A Selected Bibliography of Publications by, and about, Richard Phillips Feynman

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

31 October 2024 Version 1.190

Title word cross-reference

$14.95 [Oni15]. $15 [Ano54b]. $18.00 [Dys98]. $19.99 [Oni15]. 2 + 1 [Fey81, Fey82c]. $22.00 [Dys98]. $22.95 [Oni15]. $24.95 [Dys11a, RS12]. $26.00 [Bro06, Ryc17, Dys05]. $29.99 [Oni15, Roe12, Dys11a]. $30.00 [Kra08, Lep07, Wüt07]. $35 [Ano03b]. $50.00 [DeV00, Ano99]. $500 [Ano39]. $55.00 [Noe11]. $80.00hb/$30.00pb [Cao06]. $9.95 [Oni15]. α [GN87, Sla72]. e [BC18]. E = mc^{2} [KN19]. F (t) · r [BS96]. λ [Fey53c, Fey53a]. SU(3) [Fey65a]. U(6) ⊗ U(6) [FGMZ64]. π [BC18]. r [EFK^{+}62].

-Transition [Fey53a].

0-19-853948-7 [Tay97]. 0-226-42266-6 [Wüt07]. 0-226-42267-4 [Kra08]. 0-691-03327-7 [Bro96c]. 0-691-03685-3 [Bro96c].

1 [Rel22]. 1965 [Fey64e]. 1988 [Meh02].

2 [Rel23]. 2.0 [BCKT09]. 2002 [FRRZ04]. 2007 [JP08]. 2010 [KLR13]. 20th [Anoxx, Bre97, Gin01, Kai02]. 235 [FdHS56].

3 [Ish19, Ryc17]. 3.0 [Sem09]. 3.2 [Sem16].

40th [MKR87]. 469pp [Cao06].

8 [Roe12].

9 [BFB82]. 978 [Ish19, Roe12, Ryc17]. 978-0-06135-132-7 [Oni15]. 978-0-300-20998-3 [Ryc17]. 978-0-8090-9355-7 [Oni15]. 978-1-58834-352-9 [Oni15]. 978-1-59643-259-8 [Roe12]. 978-1-59643-827-9 [Oni15]. 978-1-59691-452-0 [Oni15]. 978-3-319-96835-3 [Jac20, Ish19]. 9th [JP08].

Abenteuer [FL14]. Abner [CHS97]. Absolute [Fey53b]. Absorber [FW41, Pri94, WF45]. Absorbing [FW46]. Absorption [Ber47]. Acquired [TF70, TF71]. Across [Bus10, II08]. Action [Fey42, WF49]. Actors [FW20]. Addison [Dys98]. Addison-Wesley [Dys98]. Advanced [Sca91]. adventure [MB11a]. Adventures [BFB82, Dre89, FL06, FL14, Sac85, Tei86, FL88, FLH85b, FLH85a]. advice [FGL^{+}12]. Again [Lee95]. Against [Wüt13a]. Age [Dys98]. Alamos [MKR87, BHB80, BFB82, Bet91b, Bet93, Fey74c, Fey80b, Fey82a, Wil75]. Albert [Hor00]. Albuquerque [Dys79b]. algorithm [XWhZ13]. ALOHA [dALM^{+}12]. Altas [Fey72a]. America [Fey63c]. American [Lit89]. among [FFF77]. Amplifier [BF46, dDD^{+}19]. amplitude [Fie06, Wan93]. amplitudes [Wes93, dALM^{+}12]. analysis [CH70, FRRZ04]. analyzer [TF00]. Ancestor [DE88]. and/or [Mil12]. Anecdotes [BFB82]. ANESTOC [FRRZ04]. Anniversary [MKR87]. anthology [War12]. Anthony [Ano99, Ano03a, DeV00]. Anthropic [Spe91]. anthropology [Fey15]. Antiparticles [Fey89b, Faĭ89]. Anton [Fis10, Fis12, Fis10, Fis12]. Any [Sta97, Kan95]. Apart [Cao06, Fre06, Kai05b, Kra08, Lep07, Wüt07]. Apostolos [Oni15]. Appendix [Fey86d]. Application [Fey46, Fey55b, Fey69a, Wüt13b]. Applications [Fey51c, Mor04, Mar06, Maz09, Sca91]. Applied [DMR57]. applying [Fey88a]. Approach [Bro05, Cal72, Fey48c, Fey49a, FBD05, ST09, BKSS10, BKSS12, Cra93, FFF78, Rel23]. approximated [BC18]. Archibald [Kla72]. Art [Roe12, Ano96, Dys11a, FF95]. articles [Ano89]. Artist [Fey89a]. aspects [Sim80]. Association [Lit89]. astounding [Haw11]. astronomy [FF91]. Atomic [Edw85, Fey53d, Fey53b, Fey53e, Fey53c, Fey54a, Oni15, Bet91b]. Atoms [Fey91a, Bro98]. atual [Fey51b]. Audio [Rat06]. auroras [Fey15]. Autobiography [For20, Dys18]. automated [BFK98]. Automatic [Nog93, Sas76, Wan93, dALM^{+}12, Sem98, Sem09, Sem16, XWhZ13]. available [Jac20]. Averages [Pop98]. Award [Ano54a, Ano54b]. Awarded [Ano35, Dys65].

B [Ano03b, Haf65, Lin66, Tre64]. Back [Ano11a, Wüt11a]. bacteriophage [EFK^{+}62]. Bang [Spe91, Spe91]. Barrow [Spe91]. Baryonic [FPT70]. Based [FMT47, FMT49b, FMT49a]. Basic [Ano05, Bro06, Dys05]. basis [Mar06, RLER07]. Be [Fey47, Fey66e, Mor95]. Beat [Cus96, SGT^{+}95, Tay97, Meh94]. Beaten [Bro06, Dys05, Fey05a, Sch06]. Beaulieu [SGT^{+}95]. beauty [Mlo03a]. Been [TSxx, Bro11]. beginners [DT16]. behavior [Fey69b, Fey81, Fey07a]. belieben [FL14]. Belief [Dys98]. Bell [Mer17, Whi16]. Below [Fey82a, Fey74c, Fey80b]. Berlin [KLR13]. Best [FR99, FR05, Pic17d, Fey00c, Fey06a]. Bethe [Goo99, Kai05a]. between [Cos50, Stö18]. bibl [Jac20]. Big [Spe91]. Biographies [FW20, Jac20, Wei88, Nob72]. Biography [Fey64e, Fey98a, For20, Oni15]. Birth [BH83a, BH83b, BH82]. birthday [Kla72]. bis [Fis10, Fis12]. Black [Lig19, Ove08, Spe91]. Black-Hole [Lig19]. Blackboard [Lig19]. Blackboards [Paz89]. Blocks [Noe11, Coo10, Coo15]. Bloomsbury [Oni15]. Blunder [BFB82]. Body [Bus10, Kai05a, II08, Mat67, Mat76, Mat92]. Bohm [For07, Jac10, Tür10, For07]. Bomb [Ral06, Oni15]. bombs [Hen10]. bongos [Hen10]. Book [And88, Ano99, Ano03a, Ano03b, Bro96b, Bro96c, Bro06, Cao06, Cus96, DeV00, Dre89, Dys91, Dys92, Dys98, Dys11a, Fey86b, For19, Fre06, Gro14, Haf65, Hug86, Hut68, Ish19, Jac20, Kai01, KYT^{+}92, Kra08, Lan11, Leb73, Lep07, Lin66, Mic67, Mul74, Noe11, Oni15, RS11, RS12, Roe12, Ryc17, Sac85, Sax94, Sch06, Spe91, Tay97, Tei86, Tre64, Tür10, Way95, Wei66, Wüt07, Dan00]. Books [Ano99, Bro06, DeV00, Dys05, Oni15, Roe12, Sta98]. Bottom [Fey59a, Fey59b, Fey61c, Fey92, Fey11e, JR06]. Boulder [Ano03b]. Boys [Ano35]. Breaking [TSxx]. Brian [Ano03b]. Brief [Spe91]. brilliant [FLS95, Fey11c]. British [DE88]. Broglie [Cos50, Fie06, Vis51]. Brooklyn [Ano35, Ano39]. Brown [Wei88]. Built [Ral06]. Bulletin [FR98]. Bust [Dys91, Lei00, Par03]. Byrne [Oni15].

C [Smi15, Stu10]. Cage [CHT15, Tre16]. Calculation [Cal72, PJ95, BFK98, FT92, Wan93]. calculations [Gro07]. cálculo [Fey51d]. Calculus [Fey33, Fey51c, Jef04, MW66, BKSS12, Fey51d, JL86]. Caltech [Ano65a]. Cambridge [Ano03a]. Can [Ano96, Fey72b]. Care [Dre89, Spe91, FL88]. Cargo [Fey74b]. Carolina [DMR57]. CD [KYT^{+}92]. CD-ROM [KYT^{+}92]. celle [Cos50]. centre [Goo01, UMF16]. Century [Anoxx, Bre97, Gin01, Kai02]. CERN [Acz10]. Challenger [DE88, Fey88c]. Cham [Ish19, Jac20]. champ [Cos50, Vis51]. Channel [Fey56b]. Chapel [DMR57]. Character [FLH85b, FLH85a, Fey94b, FL06, Hut68, Mic67, FC55, Fey65e, Fey70c, FL88, FF95, Fey12, Dre89, Sac85, Tei86]. Charles [Hor00]. chasing [Ano96]. Chemistry [Hel37]. Chicago [Cao06, Kra08, Lep07, Wüt07]. Chief [RS11]. Childs [Ano35]. Chile [FRRZ04]. Christian [Jac10, Tür10]. Christos [Oni15]. chromodynamic [FFF78]. chromodynamics [Fey82c]. Cipher [Rel22, Rel23]. ciphers [Rel23]. Citizen [Dys98, Fey98d]. City [Ano39]. Classic [FL06]. Classical [Fey48b, WF49, FK86]. Climbing [Kai05a]. Closed [Fey71a]. cloth [Bro06, Jac20, Noe11, Roe12]. CO [Ano03b]. Coined [Ove08]. Cold [CF57, For07]. Collected [Fey05a]. collection [Bus10, Kla72, Rat06]. College [Ano39]. Collider [Acz10]. Collins [Oni15]. Collisions [Fey69d, Fey69b, Fey77a]. Coloring [Roe12, Dys11a]. Colors [Ano96]. Combinatorial [Cra93]. Commentaries [Ano96]. commentary [Bro00c, Dys01]. Comments [FPT70]. Commission [Fey86d]. Common [Spe91]. Company [Lan11]. Competition [Ano39]. Complete [Rat06]. Components [FGMZ64]. Compton [BF52]. Computation [DP04, DeV00, Hey99a, Ano99, Ano03a, FHA96, HA96, Hey99b, HF02, Sei00, TF04]. computational [HM24]. computations [dALM^{+}12]. Computer [San87, Ano02, Bro00b, CH70]. Computers [Fey84, Fey86a, Hey99a, Ano02, Fey82d, Fey85b, Fey86c, Fey87b, HF02, Sei00, DeV00, Ano99, Ano03a]. computing [Fey85a, MD02]. Concept [Noe11, Coo10, Coo15, Fey51a]. Condensed [Pin89, Jis13, Jis14]. Conditions [MF56]. Conference [Fey48a, Fey74a, JP08, KLR13, DMR57]. conhecimentos [Fey51b]. Connection [Cos50, Hil89]. Connexion [Cos50]. Consequences [Fey65a]. conserved [Nog17a, Nog17b]. Conspiracy [Bro00a]. constant [BC18]. constants [Kan95]. Constructing [Ral06]. containing [Ano89]. Contest [Ano35, Ano39]. Conventional [Usd09]. Conversation [Goo99]. conversations [Bri95a]. Cooper [Ano39]. Coopersmith [Noe11]. Cornell [Bet93, Dys89, Dys93]. Correction [BBF53]. Corrections [Ano87, BF52]. Correlations [FFF77, Fey77a]. Cosmic [VF39]. Cosmological [Spe91]. Cosmology [Spe91]. cosmos [Dan00]. Could [Mer04]. Counting [Kan18]. coupling [Kan95]. Course [Fey78, DT16]. Creation [Acz10]. Creativity [Fey80a]. creators [Str11]. cryptanalysis [Rel23]. Crystal [Fey55a, FHIP62, TF70, TF71]. Cult [Fey74b]. Culture [Dyl20b, Fey66e]. Curie [Oni15, Hor00]. Curious [Dre89, FLH85b, FLH85a, FL06, Sac85, Tei86, FL88, FF95, FL14]. Current [Fey51b, FGMZ64, FKR71]. Cut [Fey48b, Fey48d]. Cut-Off [Fey48b, Fey48d].

D [Spe91]. Dabney [Ano65b]. Damping [FW41]. Darwin [Hor00, Oni15, SNA^{+}06]. data [Fey92]. David [Cao06, For07, Jac10, Tür10, Wüt07, For07, Fre06, Kra08, Lep07]. Day [Lig19, HHW99]. DC [Oni15]. Dead [Gle88, Ove08]. decomposiTion [ST09]. Definition [Her13, Cos50, Cos50]. Definitive [Fey70a]. degree [RLER07]. Derivation [Eps55]. description [Fie06]. Design [Fri14, KYT^{+}92]. Determination [Ber47]. Deuteron [LF52, LF54]. Development [Fey66c, Wüt11b, Fey66a, Fey66b, Fey98b, HM24, Wüt13b]. Deviations [Bro06, Dys05, Fey05a, Sch06]. diagonal [FC82]. Diagram [PJ95, BFK98, Jis13, Jis14, Lai98, TF00, XWhZ13, XW15, dALM^{+}12]. diagrammatic [For24]. Diagrammatica [Vel94]. Diagrams [Bro18a, Cal72, DR18, Fre06, KIH04, Kai05a, Kai05b, Kra08, Lep07, Mey08, Mey18, Stö17, Stö18, Vel94, Wüt07, Wüt11g, Wüt11h, Wüt11i, Wüt12, ABD^{+}18, Bil74, BT04, BCKT09, CH70, Cao06, DT16, Dar19, Fey71a, FT92, For24, Gro12, Gro07, HL08, HM24, JL86, Kre00, Lev90, Mar06, Mat67, Mat76, Mat92, Ohl95, Sca91, Stö22, TF04, Wan93, vO91]. DIANA [TF00, TF04]. dice [MB11a]. Dick [GM89, Ral06]. Difference [FS54]. Different [Cus96, Tay97, Wüt18, Meh94, SGT^{+}95]. differential [Mar06]. Difficulties [Fey88a]. Diffusion [FW46]. Digital [Ano96]. dignified [FL84]. dimensions [Fey81, Fey82c]. Dinge [FRL01]. Dining [Cha89]. d’interaction [Vis51]. d’Inverno [SGT^{+}95]. Dirac [And88, Bus10, DR93, FW87, II08, Wüt11c]. Direct [WF49]. Directions [MKR87, Spe91]. discourse [Edw85]. Discovery [CA15, Noe11, Coo10, Coo15]. discrete [DT08]. Discussing [Mil12, Fri11]. Discussion [Ber47, Hei62, TFJ^{+}72, Fey82c, Hei62]. Dispersion [Cao06, FdHS55, FdHS56, Fre06, Kai05b, Kra08, Lep07, Wüt07]. Dispositions [Aro97]. dissipative [FV63, FV00]. Distance [CHS97]. Distinction [Fey91b]. Disturbing [Dys79a]. Divergence [Wüt13a]. Divergence-Free [Wüt13a]. Diversions [Mer16]. Do [Ano39, Bro18a, Dre89, Spe91, FL88]. does [MB11a]. done [Str08]. Don’t [FF05]. Down [GM89]. Doxiadis [Oni15]. Dr [Ano54a]. Dramatic [Dys11a]. Drawing [Kai05b, Kra08, Ohl95, BT04, BCKT09, HL08, Lev90, Fre06, Lep07, Cao06, Wüt07]. dreams [Haw11]. Dresden [JP08]. Drum [Cus96, SGT^{+}95, Tay97, Meh94]. Dvizhenie [FLS76a]. Dynamics [Cha89]. Dyson [Bro96b, Bro96c, Sax94, SGT^{+}95, Spe91, Way95, CK74, Eps55, Hur52, JL86, Kai05c, Sch94b, Sch94a, Sta98, Wüt13b].

E-book [Jac20]. Early [SGT^{+}95, Gro12, Wüt13b]. Earth [UMF16]. easy [CD09, FLS95, Fey97, Fey11c, Fey11d]. EasyFeynDiag [XW15]. ed [Bro06, DeV00]. Edited [Ano99, Ano03a, Ano03b, Dys05]. Edition [Fey70a, Fey10a]. Editor [FD93, Lai98, Fey67b]. education [Cec11]. Effective [FK86]. einfach [FRL01]. Einführung [SGT^{+}95, Hel37]. Einstein [Hor00, Mil12, RS11, Ano54a, Ano54b, Anoxx, Coo10, Coo15, Fey97, Fri11, Kai02, KN19, Noe11, Sau08, SNA^{+}06]. Einstein’s [Fey11d]. elastic [FF77]. Electric [TF70, TF71]. Electricity [SGT^{+}95]. Electricity/Zell [SGT^{+}95]. Electrodynamics [Ano54b, Fey48b, Fey48d, Fey49a, Fey51c, Fey66c, Sch58, Sch89, Sch03, SGT^{+}95, WF49, Wüt11i, Wüt13a, Wüt18, Dar19, Fey61d, Fey61a, Fey62a, Fey66a, Fey66b, FH69, FHH^{+}62, Fey98b, For22]. Electromagnetic [Fey50, Fey65d, Fey69c]. Electromagnetism [Pic17b, FLS64, Fey10b]. Electron [TF70, TF71]. Electronic [SGT^{+}95]. Electrons [Fey55a, FHIP62]. Elektrichestvo [FLS77a]. Elektrodinamika [FL77]. Elementary [And88, BH83a, Fey62c, FW87, Her13, Lar19, BH82, Zic65]. elementary-particle [BH82]. Elements [FMT47, FMT49b, FKR71, FMT49a]. elliptic [ABD^{+}18]. Emission [FdHS55, FdHS56]. Energias [Fey72a]. energies [Fey69b]. Energy [Coo10, Coo15, FHRK51, FC56, Fey69d, Noe11]. Entanglement [CHS97, Sta97, MD02]. entdecken [FRL01]. entre [Cos50]. Entretiens [Fey00b]. Enumeration [Hur52]. Enz [SGT^{+}95]. Epic [Oni15]. Epistemology [Mey18]. Equation [FVH57, Wüt11c]. Equations [FMT47, FMT49a, FMT49b, Dys90]. Erforschung [Fis10, Fis12]. errors [Sty11]. Espionage [Usd09]. essays [Kla72]. essentials [FLS95, Fey11c]. Estado [Fey51b, Fey63b, Fey89c]. Eugene [Oni15]. EUR [Jac20]. evaluate [BS96, vO91]. Evaluation [ST09]. Evander [Ano35]. Events [Ber47]. Everything [DB88b]. Excitations [FC56, Fey58]. Excursion [Her13]. Exercises [Fey64a, FLV69, FLS14]. Exigencies [Wüt18]. experiments [EFK^{+}62]. explained [FLS95, Fey11c]. explainer [KYT^{+}92, LeV10]. Explores [Dyl20b]. Exploring [Ano99, Ano03a, DeV00, Hey99a, HF02, Sei00]. expressing [Wes93]. Extended [Fey70a, XW15]. Extension [Pop98]. Extraordinary [Jac20, Muk11, Res18, Ish19]. extreme [Fey69b].

F [Fey86d]. failings [Cre14]. Fallout [Oni15]. Fantasy [Fey80a, Fey00a]. Far [Ano35]. Faraday [CHT15, Tre16]. fast [XW15]. Father [FD93]. FBI [BLM19]. February [Meh02]. Feinmanovskie [FLS76a, FLS76b, FLS77b, FLS77a, FL77, FLS78]. Fejnman [Dai67]. Fenômenos [Fey63b, Fey89c]. Fermi [FMT47, FMT49a, FMT49b, FGM58, FG92]. Fernando [Ano03b]. Ferris [Dys05]. Ferromagnetic [MF56]. Fetter [Oni15]. Fetter-Vorm [Oni15]. FeynEdit [HL08]. Feynman [And88, Ano03b, Bet88, Bro96b, Bro96c, Bro06, Bus10, Cao06, Dai67, Dre89, Dys91, Dys98, Dys05, Dys11a, Edw85, For19, Fre06, Haf65, Hor00, Hug86, Hut68, Kai01, Kra08, Lan11, Leb73, Lep07, Lin66, Meh02, Mil12, Mul74, Noe11, Rat06, RS12, Sac85, Sax94, Sch06, SGT^{+}95, Spe91, Tei86, Tre64, Way95, Wei66, Wüt07, ABD^{+}18, AHK76, AHKM08, AEMS10, Ano35, Ano54a, Ano54b, Ano65b, Ano87, Ano88, Ano89, Ano96, Ano02, Ano05, Anoxx, Aro97, BS96, BMGW88, Bee80, Ber47, Bet91a, Bet93, Bil74, BT04, BCKT09, Bjo89, BKSS10, BKSS12, Bri95b, Bro98, BR93, Bro00a, Bro00c, Bro05, Bro06, Bro11, Bro18a, BFK98, Bry09, Cal72, CH70, Cao06, Cec11, Cha89, CD09, Coo10, Coo15, Cos50, Cra93, Cre14, CP14, CK74, DT16, Dar19]. Feynman [DB88a, Dav79, DeV00, DR18, DT08, Dur00, DR93, Dyl20a, Dys49, Dys58, Dys65, Dys89, Dys90, Dys93, Dys01, Dys03, Dys11a, Dys11b, Eps55, EFV03, Fer52, FC82, FLS65b, FLS64, FLS65a, FLS66, Fey70a, Fey72d, Fey80a, Fey86b, FL87b, FLS89, Fey89a, FD93, Fey94a, FF95, FMWH95, FHA96, Fey98a, FR99, Fey00c, Fey00b, FHM03, FBD05, FR05, Fey05a, FL06, FLS06, FGL^{+}06, Fey06a, Fey10b, Fey10a, FGL^{+}12, FL14, FLS14, Fie06, FT92, For22, For24, For07, For20, Fri11, Fri14, Gal98, GN87, GM89, Gle88, Gle92, Goo89, GFG97, GG00, Goo01, GG97, Gro12, Gro07, HL08, Hal17, HM24, Hei62, Hen10, Her13, HA96, Hey96, Hey99a, Hey99b, HF02, Hil89, Hur52, II08, Jac20, Jef04, Jis13, Jis14, JL86, Kai02, KIH04, Kai05a, Kai05b]. Feynman [Kai05c, Kan95, Kan18, KYT^{+}92, Kra11a, Kra11b, Kre00, Lai98, LeV10, Lee94, Lee95, Lei00, Lev90, Lig19, Lit89, Lub89, Mar06, Mar10, Mat67, Mat76, Mat92, Maz09, Meh94, Mer04, Mer17, Mey08, Mey18, Mic67, MD02, MW66, Mlo03a, Mlo03b, MB11b, Mor04, Muk11, NC89, Nog93, Nog17a, Nog17b, Ohl95, Oni15, Ott11, OM13, PJ95, Par03, Paz89, Pic17a, Pic17b, Pic17c, Pic17d, Pin89, Pop98, Pri94, Pup02, Res18, RLER07, Sas76, Sau08, Sca91, SNA^{+}06, Sch86, Sch94b, Sch94a, Sch94c, Sei00, Sem98, Sem09, Sem16, She05, Sim80, Sla72, ST09, Smi15, Sta98, Sta97, Stö17, Stö18, Stö22, Str08, Stu10, Sty11, Syk94, Tay97, TVOT98, Tel89, TF00, TF04, Vel94, Vis51, Wan93, Wei88, Wes93, Whe89, Whi16, Wüt11c, Wüt11d, Wüt11g]. Feynman [Wüt11h, Wüt11i, Wüt12, Wüt13a, Wüt13b, Wüt18, XWhZ13, XW15, Zeh11, dDD^{+}19, dALM^{+}12, vO91, Cus96, Dys92, Dys05, Jac10, Sch06, Tür10, Ano03b, Ben11, FGL^{+}06, Gro14, Ish19, Ano99, Ano03a]. Feynman-Diagram [PJ95]. Feynman-graph [Kan95]. Feynman. [Roe12]. Feynman/Miller [SGT^{+}95]. FeynmanParameter [Wes93]. Feynman’s [Ano89]. FeynRules [CD09]. FF [vO91]. Field [Fey67a, TF70, TF71, Wüt11d, Cos50, Dar19, Fey71c, Fey88a, Sem09, Vis51]. field-less [Dar19]. FIESTA [ST09]. Fifty [CP14]. figs [Jac20]. Files [BLM19]. Film [Hut68]. Finding [FR99, FR05, Fey00c, FRL01]. Findings [Ano65a]. fine [BC18]. Fingerübungen [Fey11b]. Finite [Fey46, TF70, TF71]. FIRE5 [Smi15]. First [Ano35, Ano39, Dys11a, Oni15, Roe12]. Fisica [Fey72a, Fey53f, Fey05b]. Fission [FdHS55, FdHS56]. Fissionable [FW46]. Five [Cec11]. Fizika [FLS70, FLS77b]. fizike [Dai67, FLS76a, FLS76b, FLS77b, FLS77a, FL77, FLS78]. flavour [Fie06]. flow [Fey55c]. Fluid [Fey54a]. fonction [Vis51]. Forces [Fey39a, Fey39b]. foreword [Dys05]. form [BS96, Wan93, FT92]. Formulae [Mor04]. formulas [FC82]. Formulation [Aro97, Fey50]. Forstner [Jac10, Tür10]. Fortgeschrittene [Fey11b]. found [Ano96]. Foundation [Aro97]. Fragments [Fey86b]. Frank [Spe91]. Free [Wüt11e, Wüt13a]. Freeman [Spe91, Sta98]. French [Cos50, Hei62, Vis51]. Fritzsch [RS11, Mil12]. Front [Ano11b, SGT^{+}95, Wüt11f]. Frontiers [H^{+}58]. frosh [Ano65b]. fulfillment [RLER07]. Full [San87]. function [Vis51]. functions [FK86, Str08]. Fundamental [Fey47, Fey59c, Fey54b, Fey61b, Fey62b, Fey95]. Further [Dre89, FL88]. Future [Ber47, Fey85a, Fey91b].

G [Ano99, Ano03a, Ano03b, DeV00]. Galaxy [VF39]. Galileo [Hor00, Cro01, RS11, SNA^{+}06]. Gauge [Fey77b]. Genealogical [Mor04]. general [FV63, FV00, Mil12]. Generalized [FMT47, FMT49b, GN87, JL86, FMT49a, Str08]. Generated [FGMZ64]. Generation [Nog93, RLER07, Wan93, Sas76, Sem98, Sem09, Sem16, XWhZ13]. generator [Kan95, XW15]. Genesis [Wüt11g, Wüt18]. genidentity [Stö22]. Genius [Ano93, Bri95a, Fey80a, For20, Gle92, SNA^{+}06, Syk94, Fey00a, Dys92]. Geometric [Sim80]. Geometrical [FVH57]. geometry [Mar06]. George [Wei88]. German [FH69, Fey86e, FRL01, FL14, Fis10, Fis12, For07, Hel37]. Germany [JP08]. Gesetze [Fey10c]. Getting [Ano39]. Gleick [Dys92]. Goals [Wüt11b]. God [Dys98, Dys98, MB11a]. Goes [FL87b]. Gold [Ano35]. Good [BR93]. graduate [Bus10]. grandeurs [Cos50]. Graph [Nog93, Kan95]. Graphic [Oni15]. graphical [BT04, BCKT09]. graphics [Lev90]. Graphs [Hur52, Kan18, Sas76]. Gravitation [Fey71b, FHM03, DMR57, Fey63a, Fey64b, Fey64c, FMWH95, Ano03b]. gravitational [Fey71c]. Great [Cro01, Hal09, Wüt11c, Bre97, LeV10]. Gribbin [Kai01, Kai01]. Group [FGMZ64]. Groveling [Ano65b]. Guide [EFV03, Fey67a, Mat67, Mat76, Mat92]. guided [HG07]. Gurr [Oni15]. Guy [GM89].

H [Oni15, Wei88]. Hadron [Acz10, Fey69b, Fey72c, Fey77a]. Hadrons [Fey69d]. Hall [Cha89, GM89]. Halpern [For19]. Hamiltonian [Wüt11j]. Hancock [SGT^{+}95]. Hancock-Beaulieu [SGT^{+}95]. Hans [Goo99, Stö22]. Harald [Mil12, RS11]. hardback [Ish19, Tay97]. Hardbound [DeV00]. hardcover [Oni15]. Harper [Oni15]. Hatfield [Ano03b]. Having [Fey51c]. Hawking [Spe91, Cro01, Dan00]. Heisenberg [Mil12, Bre97, Fri11]. Heitler [Hei62]. helicity [dALM^{+}12]. Helium [CF57, Fey53b, Fey53c, Fey53a, Fey54a, Fey55c, FC56, Fey53d, Fey53e, Fey55b, FC55, Fey58]. Hellmann [Bee80, FC82, GN87, Pop98, Pup02, RLER07, Sla72]. Heraclitus [Dan00]. Here [Ano39, Bre97]. Heuristics [Wüt13a]. Hey [Ano99, Ano03a, DeV00]. Hibbs [Wei66, Sty11]. High [Ano35, FHRK51, Fey69d, Fey77a, FF77, RLER07]. High-Energy [Fey69d]. high-transverse-momentum [FF77]. Highest [Ano35]. Hilary [Dys11a, Roe12]. Hill [DMR57, Oni15]. Hintertreppe [Fis10, Fis12]. historic [Fey57b]. Historical [Dys11b]. Histories [Oni15]. History [Dyl20b, FW20, GG00, Oni15, Spe91, KLR13]. Hole [Lig19, Ove08]. Holes [Spe91]. honor [Kla72]. Honors [Ano35]. House [Ral06]. HQ [KLR13]. HQ-3 [KLR13]. Human [Bus10, II08]. humbly [Ano65b]. hypergeometric [ABD^{+}18]. Hypotheses [SF56].

Idea [Wol16, JR06]. ideas [Bre97, Wol16]. II [Fey55c, Nog17b]. III [FLS66]. illus [Oni15]. Illustrated [Syk94]. im [For07, Jac10, Tür10]. images [FF95]. imagining [Dan00]. implementation [Fri14, Smi15]. Impossible [Wüt13a]. including [Nob72]. Independent [Eps55]. Individual [Ano35]. Industrial [Usd09]. Industrial-Scale [Usd09]. Inelastic [CF57]. Infinite [Spe91]. Infinitesimal [Fey93]. Infinity [Bro96a]. Influential [Rog10]. information [Fri14]. Inquiry [Fey88c]. insights [FGL^{+}12]. inspiration [Ano65b]. inspired [Par03]. Institutions [FW20]. Integral [ST09, Smi15, BS96, BKSS12, DT08, JL86, Sim80, Stu10]. Integrals [AHK76, AHKM08, Her13, ABD^{+}18, BKSS10, FH65, FH68, FH93, FHS10, JP08, Maz09, Str08, Wes93, Wei66]. integraly [FH68]. integration [Cra93, Sau08]. Integrity [Dyl20a, Fey90]. Interacting [Mor04, FV63, FV00]. Interaction [Fey50, FGM58, FG92, WF45, BS96, Vis51]. interactions [Fey65a, Fey65d, Fey67a, Fey69c, Fey72c]. interest [Mil12]. interface [BT04, BCKT09]. Interfaces [SGT^{+}95]. Interfaces/Meadows [SGT^{+}95]. International [DMR57, JP08, KLR13]. Interparticle [WF49]. Interpretation [Wüt11h, Zeh11]. Interpreting [Wüt12]. Interprets [Fey80a]. Interview [Ano05, Dav79]. intriguing [dDD^{+}19]. Introduction [AHKM08, Bil74, Bro18b, Lar19, Wüt11h, Dys05, Hel37]. Introductory [FLV69]. Intuitive [Lar19]. invariance [HZ01]. Invention [CA15]. Invisible [Fri91]. IP [KYT^{+}92]. ISBN [Bro96c, Ish19, Jac20, Kra08, Oni15, Roe12, Ryc17, Tay97, Wüt07]. Issue [Lub89, Ano89]. Iterated [ABD^{+}18]. Ivancevic [Bus10, Bus10]. Izluchenie [FLS76b].

J [Ano99, Ano03a, DeV00, Gro14, Spe91]. J. [PC06]. Jagdish [Cus96, Tay97]. James [Ryc17, Dys92]. January [DMR57, FRRZ04]. Japan [KIH04]. JaxoDraw [BT04, BCKT09]. Jennifer [Noe11]. jets [FFF77, Fey77c, FFF78, FF78]. Jim [Dys11a, Oni15, Roe12]. John [Dys98, For19, Kai01, Spe91, Hal17, Kla72, Ove08]. Joking [FLH85b, FLH85a, Fey86b, FL14, Sac85, Tei86]. Jonathan [Oni15]. Jörg [Ish19, Jac20]. Joseph [Spe91]. Journey [Lei00, Dys91]. Journeys [Dyl20b]. July [KLR13]. June [FRRZ04, KLR13].

Kac [Mor04]. Kaiser [Cao06, Fre06, Kra08, Lep07, Wüt07]. Kalten [For07, Jac10, Tür10]. kernel [Cos50]. Kinetika [FLS76b]. kleinsten [Fis10, Fis12]. Knots [Kre00]. knowledge [Fey51b]. Known [Ano54b]. Krauss [Dys11a, Lan11, RS12, Kra11a]. Krieg [For07, Jac10, Tür10]. kurs [Fey78]. Kvantovaía [FH68]. Kvantovaya [FLS78]. Kvanty [FLS76b].

L [Ano39, Gro14]. L. [Cos50]. Lab [Dys98]. Labyrinth [For19, Hal17]. Ladder [Kai05a]. Lagrangian [Dir33, Sem98, Sem16]. Lamb [BBF53]. LanHEP [Sem98, Sem09, Sem16]. large [FFF77, FFF78, Acz10]. large-transverse-momentum [FFF78]. Last [Dys91, Lei00, Paz89]. lasting [Dur00]. Latin [Fey63c]. Latter [Ano39]. Lattice [FW46]. Laureates’ [Nob72]. laureaty [Dai67]. Lauren [Oni15]. Law [Fey94b, Fey65e, Fey70c, Fey12, Hut68, Mic67]. Lawrence [Dys11a, Lan11, Kra11a, RS12]. Laws [And88, Fey66d, Ano02, FW87, Fey87b]. Leading [Gle88, Cro01]. Leap [Bus10, Fis10, Fis12, II08]. learned [Cec11]. Least [Fey42]. Lecture [FH69, Fey78, Fey62a, Fey62b, FHH^{+}62, Fey95, GFG97]. Lectures [CP14, Edw85, Fey59c, Fey62c, Fey64a, FLS66, Fey70a, Fey71b, FHM03, FGL^{+}06, Gro07, Haf65, Lin66, Rat06, SS98, Tre64, Ano03b, FLS65b, FLS64, FLS65a, Fey72d, FW87, FLS89, FMWH95, FHA96, Fey98c, FLS06, Fey10b, Fey10a, FLS14, Gro12, HA96, Nob72, Fey06a, And88, Leb73, Mul74]. Leeds [Pri94]. legacy [Dur00, FRRZ04]. Leibniz [Noe11, Coo10, Coo15]. Leighton [Dys91, Haf65, Lin66, Tre64, Fey00b, Par03]. lektsii [FLS76a, FLS76b, FLS77b, FLS77a, FL77, FLS78]. lektsiĭ [Fey78]. Leland [Dys11a, Roe12]. less [Dar19]. Letter [Fer52, Fey67b]. Letters [Bro06, Dys05, FD93, Fey05a, Sch06, SF56, Dys18]. Level [Ben11, Bus10, Mil12]. Levels [Wüt18]. Leyland [Oni15]. libraries [dALM^{+}12]. Lies [KYT^{+}92]. Life [Cus96, Dys92, Dys11a, Gle92, Jac20, Kai01, Kra11a, Lan11, Res18, RS12, SGT^{+}95, Tay97, Ano89, Cro01, GG97, Ish19, Kra11b, Meh94, Mlo03a, PC06]. Lifetime [Gin01]. Light [Fey13, MB11b, Pic17a, Fey88d, Fey06b, Fey10b]. Limits [DeV00, Hey99a, Ano99, Ano03a, HF02, Sei00]. linear [FV63, FV00]. Liquid [CF57, Fey53b, Fey53a, Fey54a, FC56, Fey53d, Fey53e, Fey55b, FC55, Fey58]. Lisbon [FRRZ04]. literature [HG07]. lives [Bre97, Wol16]. Logic [Bro00a]. Logicomix [Oni15]. London [Kra08, Lan11]. loop [Fey71a, FT92, Gro07, Wan93, vO91]. loop-diagrams [Wan93]. Lorentz [HZ01]. lost [GFG97]. Louis [Vis51]. Love [Oni15]. Lowbrow [Tel89]. luminaries [Bri95a].

M [Dys11a, Lan11]. M.I.T. [Ano39]. MA [Ano99, Ano03a]. Machine [Hil89]. machinery [Fey93]. machines [Bro00b, Fey85a]. Made [Bro96b, Bro96c, Sax94, Way95, CD09, Haw11, Sch94a, SGT^{+}95]. Magic [Kla72, Dar19]. Magnetic [NC89]. magnetizm [FLS77a]. magnitudes [Cos50]. Mai [FLS70]. Mainly [FLS64]. Maker [Dys18]. Makers [AG02, Wol16]. Making [For20, Bro96c]. Man [Dys05, Dys11a, Fey33, Lan11, Res18, RS12, Kra11b, Jac20, Ish19]. Manhattan [HHW99]. manipulation [BS96]. Many [Kai05a, Mat67, Mat76, Mat92]. Many-Body [Kai05a, Mat67, Mat76, Mat92]. Mapping [EFK^{+}62]. Marcolli [Ben11]. Marie [Hor00, Oni15]. Markvart [SGT^{+}95]. Mary [Kai01]. Maser [FVH57]. masers [Fey10b]. Mass [FS54]. massless [Fey71c]. Materials [FW46, SGT^{+}95]. Materials/Mönch [SGT^{+}95]. Math [Ano39, Pic17d]. Mathematical [AHK76, AHKM08, Fey50, Her13, Maz09, FRRZ04]. Mathematics [Ano35, CHT15, Stö18, FF91, Fey65b, Fey65c, Fey69a, War12]. Mathematische [SGT^{+}95]. Matilde [Ben11]. Matrix [FKR71]. Matter [Ano11a, Ano11b, Fey13, MB11b, Pin89, Wüt11a, Wüt11f, FLS64, Fey88d, Fey06b, Fie06, Jis13, Jis14, WH07]. Matthew [Haf65, Lin66, Tre64]. Max [Fis10, Fis12, Fis10, Fis12]. Maxwell [Dys90]. May [Meh02]. Meadows [SGT^{+}95]. Meaning [Dys98, Fey98d]. Means [Wüt11b, Wüt13b]. Mechanical [CHS97, Fey84, Fey86a, Ano02, Fey65f, Fey85b, Fey86c, Fey87b].

Mechanics [Aro97, DP04, Fey42, Fey48c, Fey78, FRRZ04, Leb73, Mul74, Pic17a, Pic17c, Wei66, Wüt11j, Dir33, DR93, Fey51a, Fey55b, FLS65a, FH65, FLS66, FH68, Fey72d, FH93, Fey98c, Fey10b, FHS10, Fie06, For07, Fri11, Lin66, Mil12]. Mechanism [FW41, WF45]. Mechanistic [Wüt18]. Medal [Ano35]. Medalist [Lit89]. Mehra [SGT^{+}95, Tay97, Cus96]. mekhanika [FLS78, FH68, Fey78]. mekhaniki [FLS76a]. Memorial [And88, Ano89, FW87]. Memories [BR93, Hey96]. Men [Bro96b, Bro96c, Sax94, SGT^{+}95, Way95, Eds67, Sch94a]. Merely [Ber01]. Meson [FHRK51, LF52, LF54]. mesons [Fey51b, FF77, Fey51b]. Method [Sla72, GN87]. methods [HM24]. Meÿenn [SGT^{+}95]. Michelle [Dys05, Ano05, Bro06]. Milestones [Wüt11b]. millennium [Fey10a]. Miller [SGT^{+}95]. Mills [Fey71c, Fey81]. mind [II08, Bus10]. Minds [Bro00b]. misspelling [Ano87]. Mobility [FHIP62]. Model [Fey54a, FKR71]. Modeling [Stö18]. Modelling [Gal98, Wüt18]. Models [Stö17, Wüt12]. modern [Fey66e]. Modest [Ano65a]. Molecules [Fey39a, Fey39b]. moment [Fey57b]. momenta [FFF77]. momentum [Fey77a, FFF78, FF77]. Mönch [SGT^{+}95]. monograph [Ben11, Bus10]. Monsieur [Fey00b]. Moringo [Ano03b]. Most [BR93, Rog10, FLS95, Fey11c, Haw11]. Motion [Fey65f, Fey91a, GFG97]. motives [Mar10, Ben11]. Mr [FLH85b, FLH85a, FL14, Mil12, RS11, Sac85, Tei86]. Mr. [FL87b, FL14, Fri11]. multi [Gro07]. multi-loop [Gro07]. multiverse [Bro00b]. mutants [EFK^{+}62]. My [Lig19]. Myrick [Dys11a, Oni15, Roe12, Gro14]. Mystery [Sta97, FR98, dDD^{+}19]. myths [Bro96c]. myths-in-the-making [Bro96c].

N.Y.U. [Ano35]. Nachdruck [FH69]. name [Ano87]. Nation [Ano39]. Nation-Wide [Ano39]. Nature [DR18, Gal98, Fey65f]. Naturwissenschaft [Fey86e]. nauka [FLS76a]. Near [Fey53b, Pup02]. Negative [Fey87a]. Net [Ano96]. Netze [SGT^{+}95]. Netze/Enz [SGT^{+}95]. neugierigen [FL14]. Neuronaler [SGT^{+}95]. neutrino [Fie06]. neutrinos [Fey72b]. Neutron [FW46, FS54, FdHS55, FdHS56]. Neutrons [CF57]. Newton [Mil12, Fri11, SNA^{+}06]. No [Ano93, Syk94, Fey63b, Fey89c]. Nobel [Dai67, Fey64e, Ano65a, Ano65c, Bri95a, Dys65, SS98, SAY^{+}82]. nobelevskoj [Dai67]. Non [Fey48c, Ano89]. Non-Relativistic [Fey48c]. non-technical [Ano89]. North [DMR57]. Norton [Dys11a, Lan11, RS12]. Not-so-easy [Fey97, Fey11d]. notable [Wol16]. notacao [Fey51d]. notation [Fey51d]. Note [Dys11b, Fey62a, Fey62b, FHH^{+}62, Fey95]. Notes [RS11, RS12, BCKT09, FH69]. Nothing [Mor95, Wea16, Ryc17]. nova [Fey51d]. noyaux [Cos50]. Nuclear [Fey53f, Fey05b, Usd09, Bet91b]. number [Kan18]. numbers [Nog17a, Nog17b, War12].

o [FLS76a, Fey51d]. obeying [Ano02, Fey87b]. Obituary [Bet88]. Objects [FW20]. Observations [Ber01, Fey86d]. Oersted [Lit89]. Off [Fey48b, Fey48d, FC82]. off-diagonal [FC82]. Office [GM89, Paz89]. Olum [Rel22, Rel23]. on-shell [FT92]. One [Fey88b, Kra11a, Gro07, vO91]. one- [Gro07]. one-loop [vO91]. Open [Fey56b, FRRZ04, Str08]. operacional [Fey51d]. Operational [Jef04, Fey51d, JL86]. Operator [Fey51c, MW66, Wüt11j]. Oppenheimer [Anoxx, Kai02, MS15, PC06]. optics [Fie06]. Oracle [Ano65b]. order [Kan95]. Ordinary [Ano93, Syk94]. Origin [Wüt11h, Sau08]. original [FH69]. Originalarbeiten [FH69]. oscillations [Fie06]. Other [Dre89, Fey64d, Mer16, Spe91, Bri95a, FL88]. Ottaviani [Dys11a, Gro14, Oni15, Roe12]. our [Wil75]. Outsider [Fey88c]. Owen [Ryc17]. Oxford [Noe11, Tay97].

P [And88, Ano35, Ano54a, Ano88, Ano03b, Bet91a, Bro06, Dre89, Dys98, Dys05, Fey64e, Fey72d, FF95, Fey98a, FR99, Fey00c, FRRZ04, FR05, Fey05a, For20, Goo89, Haf65, Hut68, Leb73, Lin66, Lit89, Mul74, Sac85, Sch06, Spe91, Tei86, Tre64, Vis51, Wei66, Wei88, Ryc17]. package [FT92, Sem98, Sem09, Sem16, vO91]. Page [SGT^{+}95]. pages [Ano99, Ano03a, Ano03b, Bro06, DeV00, Lep07, Noe11]. Panel [TFJ^{+}72]. Papadimitriou [Oni15]. paper [Lep07, Oni15]. Paperback [Wüt07, Bro96c, Kra08]. Papers [EFV03, Sch58, Sch03, Bro00c, Haw11]. para [Fey51d]. Parallel [TF04]. parameterization [FF78]. parameters [Wes93]. parametric [CK74]. Part [Pic17a, Pic17b, Pic17c, Pic17d]. Particle [BH83a, BH83b, Lar19, Mor04, BH82, DT16, Zic65]. Particles [And88, Fey47, FFF77, FFF78, FW87, Fis10, Fis12]. partition [FK86]. Partons [Bjo89, Fey70b, Fey74d, Fey82b, Fey72b]. Passion [CA15, CHS97]. Passion-at-a-Distance [CHS97]. past [Ber47, Fey91b]. Pasta [TSxx]. Path [AHK76, AHKM08, Her13, JP08, Sch89, Vel94, Wei66, BS96, Cra93, FH65, FH68, FH93, FHS10, Fie06, Maz09, Sau08, Str08]. Paths [Sta97, TVOT98]. patterns [Dys18]. Paul [For19]. Pauli [SGT^{+}95]. pedagogy [Gro12]. People [Dre89, Spe91, FL88, Wol16]. Perfectly [Fey05a, Sch06, Bro06, Dys05]. Perseus [Ano99, DeV00]. Personal [Ber01, Fey86d, Bet91b, Wol16]. Personalities [Dyl20b, Gin01]. perspective [Bet91b]. perspectives [JP08, KN19, Wol16]. Phenomena [FHRK51, For24, Fey63b, Fey89c]. Phillips [Bet88, EFV03, Meh02, Muk11]. Philosophy [SGT^{+}95]. Philosophy/Mehra [SGT^{+}95]. Photographic [KYT^{+}92]. Photon [Fey72c]. Photon-hadron [Fey72c]. Physical [Fey66d, Fey94b, Hal09, Hut68, MF56, Mic67, Fey65e, Fey70c, Fey12, Fie06, Mar06]. Physicist [Dyl20b, Gle88, Kra11a, Muk11, Ove08, FL14, KYT^{+}92]. Physicists [Pic17d, Bre97, Cro01]. Physics [And88, Anoxx, BH83a, BH83b, Fey62c, Fey64e, Fey64d, FLS66, FLV69, Fey86b, Fey05b, FW20, Gin01, Haf65, Her13, Jac20, Kai02, Kai05b, Kra08, Lar19, Lin66, Lit89, MKR87, Pin89, Rat06, Res18, Ryc17, SS98, Sax94, SGT^{+}95, Stö18, Ano65b, Ano65c, Ano89, Ano02, BH82, DT16, DMR57, FF91, Fey53f, Fey54b, Fey57b, FLS65b, Fey63c, FLS64, Fey64a, FLS65a, Fey82d, FW87, FLS89, FLS95, FRRZ04, FLS06, FGL^{+}06, Fey10b, Fey10a, Fey11c, FGL^{+}12, FLS14, Haw11, Ish19, Jis13, Jis14, KLR13, Mlo03a, Stö22, Str11, Wea16, WH07, Zic65, Dai67, Dys65, Fey70a, Lep07, Cao06, Fre06, Tre64, Wüt07, Edw85, FGL^{+}06]. Physics/Markvart [SGT^{+}95]. Physik [SGT^{+}95]. Physik/Sutton [SGT^{+}95]. physikalische [Fey11b]. physikalischer [Fey10c]. Physikers [FL14]. piano [Bry09]. Picture [Dys11a, Wüt13a, Fey65f]. Pictures [Gro12]. Picturing [Mey18]. Pieces [Fey97, FLS95, Fey11c, Fey11d]. Pierre [Oni15]. Pioneer [Sta87]. pioneers [DR93]. Places [Ano39]. Planck [Fis10, Fis12, Fis10, Fis12]. planets [GFG97]. play [MB11a, Par03]. Pleasure [FR99, FR05, Fey00c, FRL01]. Plenty [Fey59a, Fey59b, Fey61c, Fey92, Fey11e, JR06]. Pocono [Fey48a]. Poincaré [HZ01]. point [FH74]. Polar [Fey55a, FHIP62, TF70, TF71]. Polkinghorne [Dys98, Sta98]. popular [War12]. Pork [Mer16]. Portugal [FRRZ04]. Portuguese [Fey51b, Fey51d, Fey53f, Fey63b, Fey89c, Fey05b]. Positrons [Fey49c, Fey49b]. Postscript [Sch94b]. Postwar [Cao06, Fre06, Kai05b, Kra08, Lep07, Wüt07]. Potentiality [CHS97]. power [Bet91b]. Pp [Kra08, Tay97, Bro96c, Dys98, Dys05, Dys11a, Ish19, Jac20, Oni15, Roe12, Wüt07]. pp. [Lan11]. Practical [Fey33, Gro07]. practice [FGL^{+}12]. premii [Dai67]. Present [Acz10, Fey65d, Fey69c, Fey54b, Fey61d]. presentation [Nob72]. Press [Ano03a, Ano03b, Bro96c, Cao06, Dys98, Kra08, Lep07, Noe11, Ryc17, Tay97, Wüt07]. Price [Roe12, Lee94, Lee95]. Princeton [Bro96c]. Principle [Fey42, Fey88a, She05, Spe91]. prirode [FLS76a]. Prize [Ano39, Dai67, Dys65, Fey64e, Ano65c]. Prizes [Ano35]. probabilistic [Stö22]. Probability [Fey65f, Fey51a, Fey87a]. Probably [TSxx, Bre97]. Problem [Kai05a, Wüt18, Fey63c, FGL^{+}06, Mat67, Mat76, Mat92]. problem-solving [FGL^{+}06]. Problems [FVH57, Fey71c, Gin01, Str08]. proceedings [FRRZ04, JP08]. Processes [Fey59c, Sch94c, Fey61b, Fey62b, Fey95, Sch86]. processing [Fri14]. processor [MD02]. produced [FFF77]. production [FFF78]. professional [Bus10]. Professor [Ano54b, FL84]. Profound [Bet91b]. programs [Wes93]. Project [HHW99]. pronunciation [Mer17]. proof [Dys90]. Propagation [Wüt11e]. properties [FF78]. Prostranstvo [FLS76a]. Proton [FS54, Fey74e]. Proton-Neutron [FS54]. Pseudoscalar [LF52, LF54]. Psychology [SAY^{+}82]. Pure [DMR57]. Putnam [Ano39].

QCD [Gro07, Kan18]. QED [SGT^{+}95, Bro96c, Dar19, Fey88d, Fey06b, Fey13, Gro07, Par03, Sas76, Sch94a, Hug86, Bro96b, Sax94, Way95]. qualitative [Fey81, Fey82c]. Quantenchemie [Hel37]. Quantenelektrodynamik. [FH69]. Quantenmechanik [For07, Jac10, Tür10]. Quantensprung [Fis10, Fis12]. quantizing [Fey71c]. Quantum [Aro97, Bro00a, Bro05, CHS97, DP04, Fey42, Fey48d, Fey48c, Fey49a, Fey50, Fey51c, Fey61a, Fey62a, FH65, Fey66c, FLS66, FHH^{+}62, FFF78, Fey84, Fey85b, Fey86c, Fey86a, FH93, FRRZ04, FBD05, Fey07a, FHS10, Fie06, For19, Hal17, Hel37, HW86, II08, Kra11b, Lin66, Mil12, Pic17c, Pop98, Sch58, Sch89, Sch03, SGT^{+}95, Sta97, Wei66, Wüt11i, Wüt11j, Wüt13a, Wüt18, Ano02, Bro00b, Dir33, DR93, Fey51a, Fey55b, Fey61d, Fey63a, FV63, Fey64b, Fey64c, Fey65f, Fey66a, Fey66b, Fey82c, Fey87b, Fey88a, Fey98b, FV00, Fey10b, Fis10, Fis12, For22, Fri11, Fri14, Haw11, KLR13, MD02, MB11a, Nog17a, Nog17b, Sca91, Stö22, Str11, TVOT98, Zeh11, FLS65a, FH68, FH69, For07, RS12, Bus10, Dys11a, Lan11]. Quantum-chromodynamic [FFF78]. Quantum-Mechanical [Fey84]. Quark [FKR71, Fey77c, FF77, Mer16, Eds67, FF78, Fie06, Mer17]. Quarks [Fey67b, Fey73, Hen10]. Quest [FL87a, Bro00b]. quotable [Hor00].

R [Ano54a, FRRZ04, Wei66]. R. [Ano35, Fey72d, Mul74, Vis51, Wei66]. Radiation [Ber47, Dys01, Lig19, WF45, Dys49, Dys58, Dys03]. Radiative [BF52, FW41]. Radically [Her13]. Radioactive [Oni15]. rainbow [Mlo03a]. Ralph [Par03, Dys91, Fey00b]. Ramanujan [AEMS10]. rapports [Hei62]. Rays [VF39]. Reaction [FW41]. Reader [Spe91]. readers [Mil12]. Reading [Ano99, DeV00]. readings [WH07]. real [FR98]. Reality [Hal17, For19]. Reason [Fey89b, Faĭ89]. Reasonable [Bro06, Dys05, Fey05a, Sch06]. Receive [Ano54b]. Receives [Ano35]. Recruit [Wil75]. Redniss [Oni15]. REDUCE [Cal72]. reduction [Stu10, Smi15]. Reduze [Stu10]. reef [Ano96]. reference [dDD^{+}19]. Reflections [Gin01, FGL^{+}12]. Reflects [Kra11a]. Reichenbach [Stö22]. Reinterpreting [Pri94]. Relation [Fey64d, Fey56a]. Relativistic [BBF53, Fey48b, Fey48d, Fey48c, FKR71]. Relativitätstheorie [SGT^{+}95]. Relativitätstheorie/Wong [SGT^{+}95]. Relativity [Pic17a, DT16, Fey11d, Fey64f, Fey97, HZ01]. release [BCKT09]. Reliability [Fey86d]. Religion [SF56, Fey56a]. Remarks [Sau08]. Remembering [BMGW88]. Reminiscences [BHB80, BFB82]. renormalization [Gro07]. Reply [Lee95, Pri94]. Report [Fey86d]. Reports [Hei62]. Represent [Mey08]. Representation [DR18, FVH57, Wüt11b, BC18, For24, Wüt13b]. reprint [Fey62a, FHH^{+}62]. reproduction [FH69]. Resag [Ish19, Jac20]. research [Bus10]. researcher [Ben11]. Reshaping [Sax94]. Resonance [MF56]. Response [Sta98, BF46, dDD^{+}19]. returns [Goo01]. Revealed [Usd09]. Review [And88, Ano99, Ano03a, Ano03b, Bro96b, Bro96c, Bro06, Cao06, Cus96, DeV00, Dre89, Dys91, Dys92, Dys98, For19, Fre06, Gro14, Haf65, Hug86, Ish19, Jac20, Kai01, Kra08, Lan11, Leb73, Lep07, Lin66, Mic67, Mul74, Noe11, Rat06, Roe12, Ryc17, Sac85, Sax94, Sch06, Spe91, Tei86, Tre64, Tür10, Way95, Wei66, Wüt07, Sta98, Tay97]. Reviews [Dys11a, Hut68, Jac10, KYT^{+}92, Oni15, Ano96]. revolution [MD02]. Revolutionized [For19, Hal17]. Rhymes [Mer16]. Richard [Ano03b, Bro06, Cus96, Dys91, Dys92, Dys98, Dys05, Dys11a, For07, For19, Hor00, Jac10, Lan11, Lit89, Mic67, RS12, Sch06, SGT^{+}95, Spe91, Tay97, Tür10, And88, Ano87, Ano88, Ano02, BMGW88, Bet88, Bet91a, Bri95b, BR93, Bro00c, Bro11, Bry09, Cec11, DB88a, Dav79, Dre89, Dur00, Dys11a, Dys11b, EFV03, Fey64e, Fey80a, Fey89a, Fey94a, FF95, Fey98a, FR99, Fey00c, FR05, Fey05a, For07, For20, Gle88, Gle92, Goo89, GG00, GG97, Haf65, Hal17, Hen10, Hey96, Hey99b, Hil89, Hug86, Hut68, KYT^{+}92, Kra11a, Kra11b, LeV10, Leb73, Lei00, Lin66, Lub89, Meh94, Meh02, MB11b, Muk11, Par03, Pin89, Sac85, Sch94c, Sch06, Syk94, Tei86, Tre64, Wei88, Whi16, Kai01]. ride [Dys79b]. Ring [Usd09]. rire [Fey00b]. road [Bet91b]. Robert [Haf65, Lin66, Tre64]. Rockaway [Ano35]. Rogers [Fey86d]. role [DMR57, Fey57c, Fey66e, Gro12]. ROM [KYT^{+}92]. Room [Fey59a, Fey59b, Fey61c, Fey92, Fey11e, JR06]. Rosenberg [Usd09]. rotational [Fey55c]. roton [FC55]. routine [Lev90]. Rules [Eps55, CD09, CK74, Nog17a, Nog17b, Sem98, Sem09, Sem16]. Russian [Dai67, FH68, Fey78].

S [Bro96b, Bro96c, Way95]. S43 [Ano03a]. Said [Mer04]. SAMP [FRRZ04]. SAMP/ANESTOC [FRRZ04]. Sands [Haf65, Lin66, Tre64]. Santiago [FRRZ04]. Scale [Usd09]. Scattering [BF52, CF57, VF39, Wüt11e, FF77]. scherzen [FL14]. scholar [Fey00a]. School [Ano35]. Schrödinger [FVH57]. Schur [AEMS10]. Schweber [Bro96c, SGT^{+}95, Bro96b, Way95]. Schwinger [Bro96b, Bro96c, Dai67, Sax94, SGT^{+}95, Way95, Dys49, Dys58, Dys65, Dys01, Dys03, Sch94b, Sch94a]. Science [AG02, Ber01, CA15, Cus96, Dyl20b, Dys92, Fey56b, Fey69e, Fey74b, Gle92, Jac20, Res18, SGT^{+}95, SF56, Sta98, Tay97, Bet91b, Cec11, Fey55d, Fey56a, Fey57c, Fey86e, Fey07b, Fey11a, FR98, GG97, H^{+}58, Ish19, Kra11b, Meh94, SNA^{+}06, Dys98, RS12, Dys11a, Kai01, Lan11]. Sciences [Fey64d, Hal09, RS11]. Scientific [Dyl20a, Dyl20b, Fey90, Mer16, Fey66e, HG07, Haw11]. Scientist [Fey98d, Hor00, Dys98]. Scientists [Ber01, BLM19, Rog10, Bri95a, Hal09, Str11]. Scope [Ben11, Bus10, Mil12]. Scorer [Ano35]. Search [Oni15, Wüt13a, Eds67, KYT^{+}92, Mlo03a]. Searching [dDD^{+}19]. Sechs [Fey11b]. Second [Ano35, Ano39, Dys11a, Oni15, Roe12]. Sector [ST09]. seeks [Ano65b]. Selected [Bro00c, Sch58, Sch03]. Semiconductor [SGT^{+}95]. September [JP08]. Series [Fey59c, BC18, JL86]. Set [Leb73, Mul74, Fey72d, Fey98c]. sets [RLER07]. seven [Bro98]. shell [FT92]. SHELL2 [FT92]. Shift [BBF53]. Shimony [CHS97]. shook [Haw11]. Short [FR99, FR05, Fey00c]. should [Fey66e]. Shuttle [Fey86d]. Shvinger [Dai67]. Silk [Spe91]. Silvan [Bro96b, Way95, Bro96c]. Simon [Oni15]. Simple [Fey88b, FRL01, XWhZ13, XW15]. Simplified [Pic17a, Pic17b, Pic17c, Pic17d, Wan93]. Simulating [Fey82d]. Simulation [SGT^{+}95, Ano02]. situation [Fey54b]. Six [FLS95, Fey97, Fey11c, Fey11d]. sixtieth [Kla72]. Skeptic [She05]. slept [Bre97]. Slow [Fey55a, FHIP62]. small [Bro98]. smallest [Fis10, Fis12]. Smithsonian [Oni15]. sobre [Fey51b]. society [Fey66e, Ano35]. Solar [SGT^{+}95]. Solid [Fey63b, Fey89c]. Sólido [Fey63b, Fey89c]. Solving [FVH57, Rel22, Rel23, FGL^{+}06]. Some [Bro96c, FPT70, Mlo03b, Wol16]. Sommerfeld [BC18]. source [FF77]. Sourcebook [SGT^{+}95]. Sourcebook/Schweber [SGT^{+}95]. soustractif [Vis51]. Soviet [KIH04]. Sovremennaya [FLS76a]. Space [FW46, Fey48c, Fey49a, Fey66c, Fey97, Sch94c, CK74, Fey66a, Fey66b, Fey98b, Fey11d, Fie06, For22, Sch86, Str08]. space-time [Fie06, Sch86]. Special [Ano89, Lub89, Pic17a, DT16, Fey64f]. Spectrum [FC56]. speeches [Nob72]. Split [Kai05c]. sploshnykh [FLS77b]. sponsorship [DMR57]. Spreading [KIH04]. Springer [Ish19, Jac20]. sred [FLS77b]. stage [Goo01]. staircase [Fis10, Fis12]. Stars [VF39]. State [FMT47, FMT49b, Fey63b, Fey89c, FMT49a, Fey51b, FC55]. States [FPT70]. Statistical [Fey72d, Fey98c, FRRZ04, Leb73, Mul74, Pop98, Fey78]. Statisticheskaĭa [Fey78]. status [Fey61d, Fey65d, Fey69c]. stay [Feyxx, Sha17]. Stephen [Spe91]. Steven [And88]. Stiff [Bry09]. Stiff-string [Bry09]. Stink [Wüt18]. Stochastic [FRRZ04]. storage [Fey92]. Stories [CA15]. storm [Ano96]. Story [Acz10, Cao06, LeV10]. Strange [Fey13, MB11b, Ryc17, Fey88d, Fey06b, Wea16]. stresses [Fey39b]. string [Bry09]. strong [Fey65d, Fey67a, Fey69c]. Structure [Fey74e, SGT^{+}95, BC18]. Struggle [Wüt11c, Wüt13b]. student [Bro11, Bus10]. Studies [CHS97]. study [Fis10, Fis12]. Stuff [BR93, Haw11]. Subtle [Noe11, Coo10, Coo15]. subtractive [Vis51]. Successive [Wüt11e]. sum [TVOT98]. sum-over-paths [TVOT98]. Summary [Fey74a]. summation [BKSS10, BKSS12]. sun [GFG97, HHW99]. Superconductivity [Fey57a, GG00]. Superfluidity [Fey57a]. Superstrings [DB88b]. supplement [FGL^{+}06]. Surely [FLH85b, FLH85a, Fey86b, Sac85, Tei86, FL14]. Surfaces [SGT^{+}95]. surprise [Wüt13b]. Surprises [San87, Tre16]. Surveillance [BLM19]. survey [H^{+}58]. Sutton [SGT^{+}95]. Switzerland [Jac20]. Sycamore [Dys11a, Roe12]. Symbolic [CH70, PJ95, BS96, BKSS10, BKSS12]. Symmetries [Zic65]. Symmetry [Fey66d, Fey97, Fey65a, Fey11d]. system [FV63, FV00]. Systematization [Wüt11d]. Systems [FRRZ04, Mor04, RS11].

T [Bus10]. T4d [EFK^{+}62]. T5 [Fey49b]. Take [FH74]. Tale [Oni15]. talk [Bri95a]. Talking [Fey80a]. Tampers [Fey46]. Tannu [FL87a]. Tapes [NC89]. TCP [KYT^{+}92]. TCP/IP [KYT^{+}92]. Teacher [Goo89, Muk11, FLS95, Fey11c]. Teachers [Lit89]. Teaching [Fey86b, TVOT98, Fey63c]. Team [Ano39, Ral06]. technical [Ano89]. Technique [Hur52]. techniques [Jis13, Jis14]. Teilchen [Fis10, Fis12]. tell [Fey72b]. Téorica [Fey53f, Fey05b]. Teplota [FLS76b]. Term [Ove08]. Terms [WF49]. textbooks [Fey65b, Fey65c]. theater [FR98]. Their [HHW99, Maz09]. theology [Sta98]. Theorem [Bee80, Fey46, GN87, Pup02, RLER07, Mer17, Whi16]. Theorems [Pop98, Sla72]. Theoretical [Gle88, Wüt11d, Wüt18, DT16, Fey54b, dDD^{+}19, Fey53f, Fey05b]. théorie [Vis51]. Theories [Cao06, Dys01, FHRK51, Fre06, Kai05b, Kra08, Lep07, Wüt07, Dys49, Dys58, Dys03, Fey77b, Fey88a]. Theory [AHK76, AHKM08, Ber47, Bro05, CF57, DB88b, Eps55, FMT47, FMT49b, Fey49c, Fey49b, Fey50, Fey53b, Fey53c, Fey54a, FGM58, Fey59c, Fey64f, FG92, FBD05, Fey13, KIH04, Lee94, LF52, LF54, MB11b, Pri94, Wüt11d, Bry09, FMT49a, Fey53d, Fey53e, Fey61b, Fey62b, Fey63a, FV63, Fey64b, Fey64c, Fey67a, Fey81, Fey88d, Fey95, FV00, Fey06b, Sca91, Sem09, TVOT98, Vis51, Zeh11]. There [Fey61c, Fey92, Fey11e, JR06, Sta97]. Thesis [Bro05, FBD05]. Things [FR99, FR05, Cec11, Fey00c, FRL01, Str08]. Think [Dre89, FF05, Spe91, FL88]. Third [KLR13]. Thomas [FMT47, FMT49a, FMT49b]. Those [NC89]. Thought [Sta87]. Thoughts [Dys98, Fey98d]. Three [Fey88b]. threshold [Pup02]. thrive [Ano96]. Tie [Ano39]. Time [Eps55, FF05, For19, Hal17, Rog10, Spe91, TSxx, Wil75, Fey48c, Fey49a, Fey66a, Fey66b, Fey66c, Fey97, Fey98b, Fey11d, Fie06, For22, Mlo03b, Sch86, Sch94c, Stö22]. Time-Independent [Eps55]. times [Bre97, Cro01, KYT^{+}92]. Timothy [Dys05]. Tiny [Fey87b, Ano02]. Tipler [Spe91]. tips [FGL^{+}06, FGL^{+}12, KYT^{+}92]. today [Fey57c]. Tomonaga [Bro96c, Dai67, Dys49, Dys58, Dys65, Dys01, Dys03, Sch94b, Sch94a, SGT^{+}95, Bro96b, Way95, Sax94]. Tomorrow [SAY^{+}82]. too [Mor95]. tool [HL08]. Tools [KIH04, Fri14]. topologies [Stö22]. tour [HG07]. trace [Wes93]. Track [Bro06, Dys05, Fey05a, Sch06, Feyxx]. tracks [Sha17]. Traditions [KLR13]. traektoriíam [FH68]. trains [Feyxx, Sha17]. transformations [KLR13]. Transition [Fey53c, Fey53a]. transposition [Rel23]. transverse [FFF77, Fey77a, FFF78, FF77]. treasury [FF91]. tree [Fey71a]. trends [JP08]. Trinity [Oni15]. Trivial [FC82]. true [Mor95]. Truth [FD93, KYT^{+}92, Oni15]. Tufte [She05]. tuning [Bry09]. Turn [Bro00a, Fey47]. Tuva [Dys91, FL87a, Lei00, Par03]. Two [CHS97, Eds67, Fey54a, Fey88b, FT92, RS11]. Two-Fluid [Fey54a]. two-loop [FT92].

U [Ryc17, SGT^{+}95, Bro00a, FdHS56]. U-235 [FdHS56]. U-Turn [Bro00a]. U235 [FdHS55]. uncertainty [Fey65f, Fey07b]. undergraduates [Mil12]. Understanding [Mey18, Wüt18]. Union [DMR57, Ano39, KIH04]. Universe [Bus10, Fri91, HW86, Dan00, II08, Dys79a]. University [Bro96c, Cao06, DMR57, Dys98, Kra08, Lep07, Noe11, Tay97, Wüt07]. USA [KIH04]. Use [Wüt11h]. user [BT04, BCKT09]. Using [BS96].

V [Bus10]. value [Fey55d]. variational [Fey88a]. Velocity [TF70, TF71]. Vergnügen [FRL01]. version [XW15, BCKT09, Sem09, Sem16]. Very [Fey69d, Bro98, Fey06a]. View [Fey66c, Fey88c, SAY^{+}82, Tel89, Fey65f, Fey66a, Fey66b, FH74, Fey98b, For22]. viewpoints [Bet91b]. Virial [Pop98, Sla72]. vision [JR06]. Visual [Wüt12, KYT^{+}92]. Visualization [Sch94c, For24, Sch86]. Void [Wea16, Ryc17]. Vol [Fey10b, Haf65, Lin66, FLS64, FLS65a]. Volny [FLS76b]. Volume [Fey86d, Fey62a, Fey62b, FHH^{+}62, Fey95, CHS97, MKR87]. vom [FRL01, Fey10c]. Vorlesungsmitschrift [FH69]. Vorm [Oni15]. voulez [Fey00b]. Vous [Fey00b]. Vremya [FLS76a]. Vyp [FLS76a, FLS76b].

W [Lan11, RS12, SGT^{+}95, Spe91]. W. [Ano39]. Wagner [Ano03b]. Wallis [AEMS10]. Wang [Oni15]. War [Gal98, Wüt18, For07]. Was [Fey86e]. Washington [Oni15, FL87b]. Waves [Pic17a, Fie06]. weak [Fey65a, Fey65d, Fey69c]. wealth [War12]. Weapons [Gal98, Bet91b]. Weatherall [Ryc17]. Weinberg [And88]. Weird [Str11]. Well [Ano39]. Wesen [Fey10c]. Wesley [Dys98]. Westview [Ano03a, Ano03b]. Wheeler [For19, Ber47, Hal17, Kla72, Lee94, Lee95, Ove08, Pri94]. Who [Bro96b, Bro96c, Ove08, Sax94, Way95, Sch94a, SGT^{+}95]. Whole [TSxx]. Wide [Ano39]. Will [Ano54b, Fey47]. William [Ano03b]. Winner [Ano65a, Ano65c]. Wins [Ano35, Ano39]. Wires [Bro98]. wisdom [Hor00]. Wise [Dys05]. Without [Wüt11i, Wüt11j, Kla72]. wohl [FL14]. women [HHW99]. wonderful [Mor95]. Wong [SGT^{+}95]. words [Hor00]. Work [Ano54b, Bro18a, Kra11a, Ano89, FH69]. Worked [Lig19]. worker [Bus10]. Works [FR99, FR05, Fey00c]. World [RS11, Bro98, FF91, Fey57c, FH74, Haw11]. writing [War12]. Writings [SGT^{+}95, Par03]. Wrong [Mil12, RS11, TSxx, Fri11].

X [GN87, Sla72]. Xfey [Lai98]. xii [Jac20]. xiv [Noe11]. xix [Lep07, Wüt07]. xloops [BFK98]. xvii [RS12]. xx [Kra08]. xxiii [Bro06, DeV00]. xxvii [Bro96c]. xxx [Fey74f].

Yale [Dys98, Ryc17]. Yang [Fey71c, Fey81]. Years [CP14, HZ01, War12]. York [Bro06, Lan11, Oni15, Roe12]. Young [Whe89, UMF16]. You’re [FLH85b, FLH85a, Fey86b, FL14, Sac85, Tei86]. You’ve [TSxx].

Zakony [FLS76a]. Zeilinger [Fis10, Fis12]. Zell [SGT^{+}95]. Zero [Fey53b]. zum [Fis10, Fis12]. Zvuk [FLS76b].

References

Ablinger:2018:IEH

[ABD^{+}18] J. Ablinger, J. Blümlein, A. De Freitas, M. van Hoeij, E. Imamoglu, C. G. Raab, C.-S. Radu, and C. Schneider. Iterated elliptic and hypergeometric integrals for Feynman diagrams. Journal of Mathematical Physics, 59(6):062305, June 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aczel:2010:PCS

[Acz10] Amir D. Aczel. Present at the Creation: the Story of CERN and the Large Hadron Collider. Harmony Books, New York, NY, USA, 2010. ISBN 0-307-59167-0. xvi + 271 + 8 pp. LCCN QC787.P73 A29 2010.

Amdeberhan:2010:WRS

[AEMS10] T. Amdeberhan, O. Espinosa, V. H. Moll, and A. Straub. Wallis– Ramanujan–Schur–Feynman. American Mathematical Monthly, 117(7):618–632, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.

org/stable/pdfplus/10.4169/000298910X496741.pdf.

Allaby:2002:MS

[AG02] Michael Allaby and Derek Gjertsen. Makers of Science. Makers of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-521680-6 (set). 96 (vol. 1) pp. LCCN Q141 .A44 2002. URL http://www.gbv.de/dms/

goettingen/360423353.pdf;   http://www.loc.gov/catdir/
enhancements/fy0611/2001048396-d.html.

Albeverio:1976:MTF

[AHK76] Sergio A. Albeverio and Raphael J. Høegh-Krohn. Mathematical Theory of Feynman Path Integrals, volume 523 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LN-MAA2. ISBN 3-540-07785-5 (print), 3-540-38250-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 137 pp. LCCN

QA3 .L28 no. 523. URL http://link.springer.com/book/10.
1007/BFb0079827;   http://www.springerlink.com/content/
978-3-540-38250-8.

Albeverio:2008:MTF

[AHKM08] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. Mathematical Theory of Feynman Path Integrals: An Introduction, volume 523 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 177 pp. LCCN QA3 .L28 no. 523. URL http://link.

springer.com/book/10.1007/978-3-540-76956-9;   http://
www.springerlink.com/content/978-3-540-76956-9.

Andersen:1988:BRR

[And88] Per H. Andersen. Book review: Richard P. Feynman and Steven Weinberg, Elementary Particles and the Laws of Physics: The 1986 Dirac Memorial Lectures. Physics Today, 41(4):96, April 1988. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1935:PAM

[Ano35] Anonymous. Prizes awarded in mathematics: Boys High School, Brooklyn, wins first honors in N.Y.U. Society’s contest. Evander Childs second. R. P. Feynman of Far Rockaway, highest individual scorer, receives gold medal. New York Times, ??(??): N3, May 19, 1935. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/101522107.

Anonymous:1939:BCW

[Ano39] Anonymous. Brooklyn College wins ‘math’ prize: Places first, getting $500, in second nation-wide W. L. Putnam Contest. M.I.T. team is second. City College, Cooper Union do well — Latter and Brooklyn tie in a competition here. New York Times, ?? (??):27, April 27, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.proquest.

com/hnpnewyorktimes/docview/102795540/.

Anonymous:1954:EAD

[Ano54a] Anonymous. Einstein Award: Dr. R. P. Feynman. Nature, 173 (4403):524, March 20, 1954. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v173/n4403/pdf/173524a0.pdf.

Anonymous:1954:EAP

[Ano54b] Anonymous. Einstein Award to professor, 35: Feynman is known for his electrodynamics work — will receive $15,000. New York Times, ??(??):87, March 14, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.

proquest.com/hnpnewyorktimes/docview/112897379/.

Anonymous:1965:CNW

[Ano65a] Anonymous. Caltech Nobel winner modest on findings. Los Angeles Times, ??(??):2, October 22, 1965.

Anonymous:1965:GFH

[Ano65b] Anonymous. Groveling frosh humbly seeks physics inspiration from Oracle of Feynman at Dabney. California Tech, ??(??):1, October 28, 1965.

Anonymous:1965:NPP

[Ano65c] Anonymous. Nobel physics prize winner, October 21, 1965. 1 photographic print from a photograph taken in 1963.

Anonymous:1987:CMR

[Ano87] Anonymous. Corrections: [misspelling of Richard Feynman’s name]. IEEE Spectrum, 24(4):10, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:RPF

[Ano88] Anonymous. Richard P. Feynman 1918–1988. Scientific American, 258(6):38–42, June 1988. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v258/n6/pdf/
scientificamerican0688-38b.pdf.

Anonymous:1989:SFM

[Ano89] Anonymous. Special Feynman memorial issue containing nontechnical articles on Feynman’s life and work in physics. Physics Today, 42(2):??, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://ptonline.aip.

org/dbt/dbt.jsp?KEY=PHTOAD&Volume=42&Issue=2.

Anonymous:1993:QTT

[Ano93] Anonymous. No ordinary genius. BBC TV ’Horizon’, 1993. A two-part documentary (50 minutes each) about Feynman’s life and work, with contributions from colleagues, friends and family. Republished by PBS Nova as “The Best Mind Since Einstein”.

Anonymous:1996:RCC

[Ano96] Anonymous. Reviews and Commentaries: Can art thrive on the Net?. colors of the reef. digital storm chasing. Feynman found. Scientific American, 275(2):104–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/0896issue/0896currentissue.html.

Anonymous:1999:BRFf

[Ano99] Anonymous. Book review: Feynman and computation: Exploring the limits of computers: Edited by Anthony J. G. Hey. Perseus Books, Reading, MA. (1999). 438 pages. $50.00. Computers and Mathematics with Applications, 37(10):173, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122199903358.

Anonymous:2002:RFQ

[Ano02] Anonymous. Richard Feynman on “quantum physics and computer simulation” and “tiny computers obeying quantum mechanical laws”. Los Alamos Science, 27:viii–ix, November 27, 2002. CODEN LASCDI. ISSN 0273-7116. URL http://library.lanl.

gov/cgi-bin/getfile?27-02.pdf.

Anonymous:2003:BRFc

[Ano03a] Anonymous. Book review: Feynman and computation: Exploring the limits of computers: Edited by Anthony J. G. Hey. Westview Press, Cambridge MA. (2002). 438 pages S43. Computers and Mathematics with Applications, 45(10–11):1781–1782, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122103801529.

Anonymous:2003:BRFd

[Ano03b] Anonymous. Book review: Feynman lectures on gravitation: By Richard P. Feynman, Fernando B. Moringo, and William G. Wagner. Edited by Brian Hatfield. Westview Press, Boulder, CO. (2003). 232 pages. $35. Computers and Mathematics with Applications, 45(10–11):1782, May/June 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0898122103801530.

Anonymous:2005:BFI

[Ano05] Anonymous. Basic Feynman: An interview with Michelle Feynman. Web site, 2005. Comments about the preparation of [Fey05a].

Anonymous:2011:BMa

[Ano11a] Anonymous. Back matter. In Adrian Wüthrich, editor, The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 191–208. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.springer.

com/content/pdf/bbm:978-90-481-9228-1/1.pdf.

Anonymous:2011:FMa

[Ano11b] Anonymous. Front matter. In Adrian Wüthrich, editor, The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xvii. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.springer.

com/content/pdf/bfm:978-90-481-9228-1/1.pdf.

Anonymous:20xx:EOF

[Anoxx] Anonymous. Einstein, Oppenheimer, Feynman: Physics in the 20th Century. Technical Report MIT OpenCourseWare STS.042J / 8.225J, MIT, Cambridge, MA, USA, 20xx. Free, independent study course that explores the changing roles of physics and physicists during the 20th Century.

Aronson:1997:DFF

[Aro97] Jerrold L. Aronson. Dispositions as the foundation for Feynman’s formulation of quantum mechanics. Dialectica: International Review of Philosophy of Knowledge, 51(1):35–64, March 1997. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baranger:1953:RCL

[BBF53] M. Baranger, H. A. Bethe, and R. P. Feynman. Relativistic correction to the Lamb shift. Physical Review (2), 92(2):482–501, October 15, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Bucknum:2018:SFS

[BC18] Michael J. Bucknum and Eduardo A. Castro. Sommerfeld’s fine structure constant approximated as a series representation in e and π. Journal of Mathematical Chemistry, 56(3):651–655, March 2018. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s10910-017-0835-8.

Binosi:2009:JGU

[BCKT09] D. Binosi, J. Collins, C. Kaufhold, and L. Theussl. JaxoDraw: a graphical user interface for drawing Feynman diagrams. Version 2.0 release notes. Computer Physics Communications, 180(9):1709– 1715, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000757.

Beebe:1980:HFT

[Bee80] Nelson H. F. Beebe. On the Hellmann–Feynman theorem. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, August 27, 1980. 3 pp.

Bengtsson:2011:FMM

[Ben11] Ingemar Bengtsson. Feynman Motives, by Matilde Marcolli, Scope: monograph. Level: researcher. Contemporary Physics, 52(4):366, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berenda:1947:DPF

[Ber47] Carlton W. Berenda. The determination of past by future events. A discussion of the Wheeler–Feynman absorption–radiation theory. Philosophy of Science, 14(1):13–19, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185530.

Bernstein:2001:MPO

[Ber01] Jeremy Bernstein. The Merely Personal: Observations on Science and Scientists. Ivan R. Dee, Chicago, IL, USA, 2001. ISBN 1-56663-344-3. 245 pp. LCCN Q171 .B5375 2001.

Bethe:1988:ORP

[Bet88] Hans A. Bethe. Obituary: Richard Phillips Feynman (1918– 1988). Nature, 332(6165):588, April 14, 1988. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v332/n6165/332588a0.

Bethe:1991:RPF

[Bet91a] Hans A. Bethe. Richard P. Feynman. In The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science [Bet91b], pages 239–244. ISBN 0-88318-707-8, 0-671-74012-1. LCCN U264 .B455 1991. US$24.95.

Bethe:1991:RAP

[Bet91b] Hans Albrecht Bethe. The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science. Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1991. ISBN 0-88318-707-8, 0-671-74012-1. xvii + 286 pp. LCCN U264 .B455 1991. US$24.95.

Bethe:1993:FAC

[Bet93] Hans Albrecht Bethe. Feynman in Los Alamos and Cornell. In Brown and Rigden [BR93], pages 33–36. ISBN 0-88318-870-8. LCCN QC16.F49 A3 1993. Commentary by Joan Feynman, John Wheeler, Hans Bethe, Julian Schwinger, Murray Gell-Mann, Daniel Hillis, David Goodstein, Freeman Dyson, and Laurie Brown.

Bacher:1946:AR

[BF46] R. F. Bacher and R. P. Feynman. Amplifier response. Report, Los Alamos Scientific Laboratory, Atomic Energy Commission, Los Alamos, NM, USA, 1946. 15 pp. Edited and declassified work from the Manhattan Project. Department of State Publication 2661 USGPO.

Brown:1952:RCC

[BF52] Laurie M. Brown and Richard P. Feynman. Radiative corrections to Compton scattering. Physical Review (2), 85(2):231–244, January 15, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Baker:1982:AAB

[BFB82] C. L. Baker, Richard P. Feynman, and Bernice Brode. Anecdotes: The Adventures of a Blunder 9; Reminiscences of Los Alamos. Annals of the History of Computing, 4(1):60–64, January/March 1982. CODEN AHCOE5. ISSN 0164–1239. URL

http://dlib.computer.org/an/books/an1982/pdf/a1060.pdf;
 http://www.computer.org/annals/an1982/a1060abs.htm.

Brucher:1998:XAF

[BFK98] L. Brücher, J. Franzkowski, and D. Kreimer. xloops — automated Feynman diagram calculation. Computer Physics Communications, 115(2–3):140–160, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465598001623.

Brown:1982:BEP

[BH82] Laurie M. Brown and Lillian Hoddeson. The birth of elementaryparticle physics. Physics Today, 35(4):36–43, April 1982. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v35/

i4/p36_s1. This is a much-abridged version of [BH83a].

Brown:1983:BEP

[BH83a] Laurie M. Brown and Lillian Hoddeson. The birth of elementary particle physics. In The Birth of Particle Physics [BH83b], pages 3–36. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Brown:1983:BPP

[BH83b] Laurie M. Brown and Lillian Hoddeson, editors. The Birth of Particle Physics. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Badash:1980:RA

[BHB80] Lawrence Badash, Joseph Oakland Hirschfelder, and Herbert P. Broida, editors. Reminiscences of Los Alamos, 1943–1945, volume 5 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). xxi + 188 pp. LCCN QC791.96 .R44.

Bilenskii:1974:IFD

[Bil74] S. M. (Samoil Mikhelevich) Bilenskii, editor. Introduction to Feynman diagrams, volume 65 of International series of monographs in natural philosophy. Pergamon Press, New York, NY, USA, 1974. ISBN 0-08-017799-9. viii + 186 pp. LCCN QC794.6.F4 B5413. Translated and edited by Frances Pardee.

Bjorken:1989:FP

[Bjo89] James D. Bjorken. Feynman and partons. Physics Today, 42(2): 56–59, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Blumlein:2010:SSA

[BKSS10] Johannes Blümlein, Sebastian Klein, Carsten Schneider, and Flavia Stan. A symbolic summation approach to Feynman integrals. ACM Communications in Computer Algebra, 44(3):95–96, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Blumlein:2012:SSA

[BKSS12] Johannes Blümlein, Sebastian Klein, Carsten Schneider, and Flavia Stan. A symbolic summation approach to Feynman integral calculus. Journal of Symbolic Computation, 47(10):1267–1289, October 2012. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0747717111002471.

Brown:2019:SUS

[BLM19] J. Patrick Brown, Beryl Lipton, and Michael Morisy. Scientists under Surveillance: the FBI Files. MIT Press, Cambridge, MA, USA, 2019. ISBN 0-262-53688-9 (paperback), 0-262-35302-4 (e-book). xviii + 413 pp. LCCN Q125 .S4366 2019. URL http://mitpress. mit.edu/9780262536882. Foreword by Steven Aftergood and introduction by Walter V. Robinson.

Barker:1988:RRF

[BMGW88] Robert J. Barker, Foster Morrison, Alexander J. Glass, and Robert Sherman Wolff. Remembering Richard Feynman. Physics Today, 41(5):134–138, May 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/PHTOAD/v41/i5.

Brown:1993:MGS

[BR93] Laurie M. Brown and John S. Rigden, editors. Most of the Good Stuff: Memories of Richard Feynman. American Institute of Physics, Woodbury, NY, USA, 1993. ISBN 0-88318-870-8. 181 + 16 pp. LCCN QC16.F49 A3 1993. Commentary by Joan Feynman, John Wheeler, Hans Bethe, Julian Schwinger, Murray Gell-Mann, Daniel Hillis, David Goodstein, Freeman Dyson, and Laurie Brown.

Brennan:1997:HPS

[Bre97] Richard P. Brennan. Heisenberg probably slept here: the lives, times, and ideas of the great physicists of the 20th Century. Wiley, New York, NY, USA, 1997. ISBN 0-471-15709-0 (cloth). xi + 274 pp. LCCN QC15 .B74 1997; 97.E02533.

Brian:1995:GTC

[Bri95a] Denis Brian. Genius talk: conversations with Nobel scientists and other luminaries. Plenum Press, New York, NY, USA; London, UK, 1995. ISBN 0-306-45089-5. xii + 423 pp. LCCN fa1941.

Brian:1995:RF

[Bri95b] Denis Brian. Richard Feynman. In Genius talk: conversations with Nobel scientists and other luminaries [Bri95a], pages 35–60. ISBN 0-306-45089-5. LCCN fa1941.

Broderick:1996:I

[Bro96a] Matthew Broderick. Infinity. Movie, 1996. Directed by Matthew Broderick, and starring Matthew Broderick as Feynman, depicting Feynman’s love affair with his first wife and ending with the Trinity test.

Brown:1996:BRB

[Bro96b] Laurie M. Brown. Book review: QED and the Men Who Made It: Dyson, Feynman, Schwinger, and Tomonaga by Silvan S. Schweber. Isis, 87(1):204–205, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/235809.

Brown:1996:BRS

[Bro96c] Laurie M. Brown. Book review: Some QED myths-in-the-making?: Silvan S. Schweber, QED and the Men Who Made It: Dyson, Feynman, Schwinger and Tomonaga (Princeton University Press, 1994), xxvii + 732 pp., ISBN 0-691-03685-3, 0-691-03327-7 (paperback). Studies in History and Philosophy of Modern Physics, 27 (1):81–90, March 1996. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/1355219895000232. See [Sch94a].

Broglia:1998:WSA

[Bro98] Ricardo A. Broglia. Wires of seven atoms — Feynman’s very, very small world. Contemporary Physics, 39(5):371–376, September/ October 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brown:2000:LQC

[Bro00a] J. R. (Julian Russell) Brown. The logic of the quantum conspiracy: Feynman’s U-turn. In Minds, machines, and the multiverse: the quest for the quantum computer [Bro00b], page ?? ISBN 0-684-81481-1, 0-7432-4263-7. LCCN QA76.889 .B76 2000. URL http:

//catdir.loc.gov/catdir/bios/simon054/99056638.html;
http://catdir.loc.gov/catdir/description/simon032/99056638.
html;   http://catdir.loc.gov/catdir/enhancements/fy0705/
99056638-s.html;   http://catdir.loc.gov/catdir/enhancements/
fy0707/99056638-t.html.

Brown:2000:MMM

[Bro00b] J. R. (Julian Russell) Brown. Minds, machines, and the multiverse: the quest for the quantum computer. Simon and Schuster, New York, NY, USA, 2000. ISBN 0-684-81481-1, 0-7432-4263-7. 396 pp. LCCN QA76.889 .B76 2000. URL http:

//catdir.loc.gov/catdir/bios/simon054/99056638.html;
http://catdir.loc.gov/catdir/description/simon032/99056638.
html;   http://catdir.loc.gov/catdir/enhancements/fy0705/
99056638-s.html;   http://catdir.loc.gov/catdir/enhancements/
fy0707/99056638-t.html.

Brown:2000:SPR

[Bro00c] Laurie M. Brown, editor. Selected papers of Richard Feynman: with commentary, volume 27 of World Scientific Series in 20th Century Physics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2000. ISBN 981-02-4130-5, 981-02-4131-3 (paperback), 981-238-546-0 (e-book). xii + 999 pp.

LCCN QC3 .F4513 2000. URL http://www.worldscibooks.com/
physics/4270.html.

Brown:2005:FTN

[Bro05] Laurie M. Brown. Feynman’s Thesis — a New Approach To Quantum Theory. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2005. ISBN 981-256-366-0 (print), 981-256-763-1 (e-book). 142 pp. LCCN QC174.12 .F49 2005. URL http://public.eblib.

com/EBLPublic/PublicView.do?ptiID=244532;   http://www.
worldscibooks.com/promotion/feynman.html#5852.

Brown:2006:BRM

[Bro06] Laurie M. Brown. Book review: Michelle Feynman, ed., Perfectly Reasonable Deviations from the Beaten Track: The Letters of Richard P. Feynman. New York, Basic Books, 2005, xxiii + 486 pages. $26.00 (cloth). Physics in Perspective (PIP), 8(4):473–474, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Brown:2011:BSR

[Bro11] Laurie M. Brown. To have been a student of Richard Feynman. Resonance, 16(9):874–878, September 2011. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic).

Brown:2018:HDF

[Bro18a] James Robert Brown. How do Feynman diagrams work? Perspectives on Science, 26(4):423–442, July–August 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

http://www.mitpressjournals.org/doi/abs/10.1162/posc_a_
00281.

Brown:2018:I

[Bro18b] James Robert Brown. Introduction. Perspectives on Science, 26(4): 419, July–August 2018. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_a_00280.

Bryner:2009:SST

[Bry09] John C. Bryner. Stiff-string theory: Richard Feynman on piano tuning. Physics Today, 62(12):46–49, December 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bao:1996:USM

[BS96] Min-Qi Bao and Anthony F. Starace. Using symbolic manipulation to evaluate Feynman’s path integral for an interaction of the form F (t)·r. Computers in Physics, 10(1):89–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168563.

Binosi:2004:JGU

[BT04] D. Binosi and L. Theußl. JaxoDraw: a graphical user interface for drawing Feynman diagrams. Computer Physics Communications, 161(1–2):76–86, August 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465504002115.

Bussey:2010:QLD

[Bus10] Dr. P. J. Bussey. Quantum Leap: From Dirac and Feynman, Across the Universe to Human Body and Mind, by V. Ivancevic and T. Ivancevic, Scope: monograph collection. Level: graduate student and professional research worker. Contemporary Physics, 51(2): 185–186, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Charman-Anderson:2015:PSS

[CA15] Suw Charman-Anderson, editor. A Passion for Science: Stories of Discovery and Invention. FindingAda, London, UK, 2015. ???? pp.

Calmet:1972:RAC

[Cal72] Jacques Calmet. A REDUCE approach to the calculation of Feynman diagrams. Computer Physics Communications, 4(2):199–204, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900082.

Cao:2006:BRS

[Cao06] Tian Yu Cao. Book review: The story of Feynman diagrams: Drawing Theories Apart: The Dispersion of Feynman Diagrams in Postwar Physics, by David Kaiser, 2005. University of Chicago Press, 469pp $80.00hb/$30.00pb. Physics World, 19(1):36–37, January 2006. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058

(electronic). URL http://iopscience.iop.org/pwa/full/pwa-
pdf/19/1/phwv19i1a38.pdf.

Christensen:2009:FFR

[CD09] Neil D. Christensen and Claude Duhr. FeynRules — Feynman rules made easy. Computer Physics Communications, 180(9):1614– 1641, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000733.

Ceceri:2011:FTL

[Cec11] Kathy Ceceri. Five things I learned from Richard Feynman about science education. Resonance, 16(9):879–880, September 2011. CO-DEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic).

Cohen:1957:TIS

[CF57] Michael Cohen and Richard P. Feynman. Theory of inelastic scattering of cold neutrons from liquid helium. Physical Review (2), 107(1):13–24, July 1, 1957. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Campbell:1970:SAF

[CH70] J. A. Campbell and Anthony C. Hearn. Symbolic analysis of Feynman diagrams by computer. Journal of Computational Physics, 5(2):280–327, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900641.

Chao:1989:FDH

[Cha89] Benjamin Fong Chao. Feynman’s dining hall dynamics. Physics Today, 42(2):15, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Cohen:1997:PEP

[CHS97] Robert S. Cohen, Michael Horne, and John Stachel, editors. Potentiality, Entanglement and Passion-at-a-Distance: Quantum Mechanical Studies for Abner Shimony. Volume Two, volume 194 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346. LCCN QC173.96-174.52. URL https://link.

springer.com/book/10.1007/978-94-017-2732-7.

Chapman:2015:MFC

[CHT15] S. Jonathan Chapman, David P. Hewett, and Lloyd N. Trefethen. Mathematics of the Faraday cage. SIAM Review, 57(3):398–417, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cvitanovic:1974:FDR

[CK74] Predrag Cvitanovic and T. Kinoshita. Feynman–Dyson rules in parametric space. Physical Review D (Particles and Fields), 10 (??):3978–??, December 15, 1974. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.10.3978.

Coopersmith:2010:ESC

[Coo10] Jennifer Coopersmith. Energy, the subtle concept: the discovery of Feynman’s blocks from Leibniz to Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-954650-9 (hardback). xiv + 400 pp. LCCN QC72 .C66 2010.

Coopersmith:2015:ESC

[Coo15] Jennifer Coopersmith. Energy, the subtle concept: the discovery of Feynman’s blocks from Leibniz to Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, revised edition, 2015. ISBN 0-19-871674-5 (paperback). xviii + 422 pp. LCCN QC72 .C66 2015.

CostadeBeauregard:1950:CED

[Cos50] O. Costa de Beauregard. Connexion entre la définition des noyaux de Feynman et celle des grandeurs de champ de L. de Broglie. (French) [Connection between the kernel definition of Feynman and that of the field magnitudes of L. of Broglie]. Comptes rendus de l’Académie des sciences, Paris, 230(??):2073–2075, ???? 1950. ISSN 0001-4036 (print), 2419-6304 (electronic).

Crease:2014:FLF

[CP14] Robert P. Crease and Peter Pesic. The Feynman lectures, fifty years on. Physics in Perspective (PIP), 16(2):143–145, June 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-014-0137-5.

Crandall:1993:CAF

[Cra93] Richard E. Crandall. Combinatorial approach to Feynman path integration. Journal of Physics A (Mathematical and General), 26(14):3627–3648, July 21, 1993. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic). URL http://stacks.iop.

org/0305-4470/26/3627.

Crease:2014:FF

[Cre14] Robert P. Crease. Feynman’s failings. Physics World, 27(03):25, 2014. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058

(electronic). URL http://stacks.iop.org/2058-7058/27/i=03/
a=34.

Cropper:2001:GPL

[Cro01] William H. Cropper. Great physicists: the life and times of leading physicists from Galileo to Hawking. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-513748-5. xii + 500 pp. LCCN QC15 .C76 2001. URL http:

//www.loc.gov/catdir/enhancements/fy0637/2001021611-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2001021611-b.html.

Cushing:1996:BRBa

[Cus96] James T. Cushing. Book review: The Beat of a Different Drum: The Life and Science of Richard Feynman by Jagdish Mehra. Isis, 87(2):387–388, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/236135.

Daison:1967:TSF

[Dai67] F. Daison. Tomonaga, Shvinger i Fejnman — laureaty nobelevskoj premii po fizike. (Russian) [Tomonaga, Schwinger and Feynman — Nobel Prize in Physics]. Uspekhi Fizicheskikh Nauk, 91(1):71– 73, 1967. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652

(electronic). URL http://ufn.ru/ru/articles/1967/1/f/.

deAquino:2012:AAL

[dALM^{+}12] Priscila de Aquino, William Link, Fabio Maltoni, Olivier Mattelaer, and Tim Stelzer. ALOHA: Automatic libraries of helicity amplitudes for Feynman diagram computations. Computer Physics Communications, 183(10):2254–2263, October 2012. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465512001725.

Danielson:2000:BCI

[Dan00] Dennis Richard Danielson, editor. The book of the cosmos: imagining the universe from Heraclitus to Hawking. Helix books. Perseus Publishers, Cambridge, MA, USA, 2000. ISBN 0-7382-0247-9. xxxiii + 556 pp. LCCN QB981 .B724 2000. URL http://catdir.

loc.gov/catdir/enhancements/fy0831/2001268173-d.html.

Darrigol:2019:MFQ

[Dar19] Olivier Darrigol. The magic of Feynman’s QED: from field-less electrodynamics to the Feynman diagrams. European Physical Journal H, 44(4–5):349–369, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.

springer.com/article/10.1140/epjh/e2019-100025-2.

Davis:1979:FRI

[Dav79] Monte Davis. Interview: Richard Feynman. Omni (New York), 1 (8):96–98, 113–114, 136, May 1979. ISSN 0149-8711. URL http:

//archive.org/details/omni-magazine-1979-05;   http://
omnimagindex.wordpress.com/author-indices/authors-d/.

Davies:1988:RF

[DB88a] P. C. W. Davies and J. R. (Julian Russell) Brown. Richard Feynman. In Superstrings: a Theory of Everything? [DB88b], page ?? ISBN 0-521-35462-5 (hardcover), 0-521-35741-1 (paperback), 0-521-43775-X (paperback). LCCN QC794.6.S85 S869 1988.

Davies:1988:STE

[DB88b] P. C. W. Davies and J. R. (Julian Russell) Brown, editors. Superstrings: a Theory of Everything? Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-35462-5 (hardcover), 0-521-35741-1 (paperback), 0-521-43775-X (paperback). viii + 234 pp. LCCN QC794.6.S85 S869 1988.

dAlessandro:2019:SRI

[dDD^{+}19] Vincenzo d’Alessandro, Santolo Daliento, Marco Di Mauro, Salvatore Esposito, and Adele Naddeo. Searching for a response: the intriguing mystery of Feynman’s theoretical reference amplifier. European Physical Journal H, 44(4–5):331–347, November 2019. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/
10.1140/epjh/e2019-90071-6.

Davidson:1988:CBA

[DE88] Gerald T. Davidson and David S. Evans. Challenger’s British ancestor. Physics Today, 41(5):138–139, May 1988. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v41/

i5/p138_s1. See [Fey88c].

DeVries:2000:BRA

[DeV00] Paul L. DeVries. Book review: Anthony J. G. Hey, ed., Feynman and Computation: Exploring the Limits of Computers. Reading: Perseus Books, 1999, xxiii + 438 pages. $50.00 (Hardbound). Physics in Perspective (PIP), 2(2):220, June 2000. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Dirac:1933:LQM

[Dir33] P. A. M. Dirac. The Lagrangian in quantum mechanics. Physikalische Zeitschrift der Sowjetunion, 3(??):64–72, ???? 1933. CODEN PHZSAL. ISSN 0369-9811.

DeWitt-Morette:1957:CRG

[DMR57] Cécile DeWitt-Morette and Dean Rickles, editors. Conference on the role of gravitation in physics at the University of North Carolina, Chapel Hill [January 18–23, 1957, under the sponsorship of the International Union of Pure and Applied Physics, and others, WADC technical report 57-216. Wright Air Development Center, Air Research and Development Command, U.S. Air Force, Wright-Patterson Air Force Base, OH, USA, 1957. LCCN UG633 .A3776 no. 57-216.

DHooghe:2004:QMC

[DP04] Bart D’Hooghe and Jaroslaw Pykacz. Quantum mechanics and computation. Foundations of Science, 9(4):387–404, December 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10699-005-4827-y.

Dutt:1993:DFP

[DR93] Ranabir Dutt and Asim K. Ray, editors. Dirac and Feynman: pioneers in quantum mechanics. Wiley Eastern, New Delhi, India, 1993. ISBN 81-224-0493-6. viii + 214 pp. LCCN QC174.26.W28 D39 1993.

Dorato:2018:NRF

[DR18] Mauro Dorato and Emanuele Rossanese. The nature of representation in Feynman diagrams. Perspectives on Science, 26(4):443–458, July–August 2018. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://www.mitpressjournals.
org/doi/abs/10.1162/posc_a_00282.

Drell:1989:BRR

[Dre89] Sidney Drell. Book review: Richard P. Feynman, What Do You Care What Other People Think?: Further Adventures of a Curious Character. Physics Today, 42(2):106–107, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Dorlas:2008:DFI

[DT08] T. C. Dorlas and E. G. F. Thomas. The discrete Feynman integral. Journal of Mathematical Physics, 49(9):092101, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i9/p092101_s1.

DAuria:2016:SRF

[DT16] R. D’Auria and Mario Trigiante. From special relativity to Feynman diagrams: a course in theoretical particle physics for beginners. UNITEXT for Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2016.

ISBN 3-319-22013-6 (print), 3-319-22014-4 (e-book). ISSN 2198-7882. xvi + 601 pp. LCCN QC793.2; QC174.45-174.52. URL

http://link.springer.com/10.1007/978-3-319-22014-7.

Durrani:2000:LLR

[Dur00] Matin Durrani. The lasting legacy of Richard Feynman. Physics World, 13(12):10–11, December 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.

iop.org/2058-7058/13/i=12/a=12.

Dylla:2020:FSI

[Dyl20a] H. Frederick Dylla. Feynman’s scientific integrity. In Scientific Journeys: a Physicist Explores the Culture, History and Personalities of Science [Dyl20b], pages 25–28. ISBN 3-030-55799-5 (paperback), 3-030-55800-2 (e-book). LCCN Q125 .D955 2020.

Dylla:2020:SJP

[Dyl20b] H. Frederick Dylla. Scientific Journeys: a Physicist Explores the Culture, History and Personalities of Science. Springer, Cham, Switzerland, 2020. ISBN 3-030-55799-5 (paperback), 3-030-55800-2 (e-book). xxv + 222 pp. LCCN Q125 .D955 2020.

Dyson:1949:RTT

[Dys49] F. J. Dyson. The radiation theories of Tomonaga, Schwinger, and Feynman. Physical Review (2), 75(3):486–502, February 1, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.75.

486. Reprinted in [Dys58].

Dyson:1958:RTT

[Dys58] Freeman J. Dyson. The radiation theories of Tomonaga, Schwinger, and Feynman. In Schwinger [Sch58], page ?? ISBN 0-486-60444-6. LCCN QC680 .S35. Reprint of [Dys49].

Dyson:1965:TSF

[Dys65] Freeman J. Dyson. Tomonaga, Schwinger, and Feynman awarded Nobel Prize for Physics. Science, 150(3696):588–589, October 29, 1965. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/1717147;   http:
//www.sciencemag.org/content/150/3696/588.full.pdf.

Dyson:1979:DU

[Dys79a] Freeman J. Dyson. Disturbing the Universe. Harper & Row, New York, NY, USA, 1979. ISBN 0-06-011108-9. x + 283 pp. LCCN QC16.D95 A33 1979. US$12.95. Commissioned by the Science Book Program of the Alfred P. Sloan Foundation. Parts of the book were published in The New Yorker, August 6, 13, and 20, 1979, and in The Observer, October 28, 1979.

Dyson:1979:RA

[Dys79b] Freeman J. Dyson. A ride to Albuquerque. In Disturbing the Universe [Dys79a], pages 58–69. ISBN 0-06-011108-9. LCCN QC16.D95 A33 1979. US$12.95. Commissioned by the Science Book Program of the Alfred P. Sloan Foundation. Parts of the book were published in The New Yorker, August 6, 13, and 20, 1979, and in The Observer, October 28, 1979.

Dyson:1989:FC

[Dys89] Freeman J. Dyson. Feynman at Cornell. Physics Today, 42(2):32– 38, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/

42/32/1. Special issue on Richard Feynman.

Dyson:1990:FPM

[Dys90] Freeman J. Dyson. Feynman’s proof of the Maxwell equations. American Journal of Physics, 58(3):209–211, 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/58/209/1.

Dyson:1991:BRR

[Dys91] Freeman Dyson. Book review: Ralph Leighton, Tuva or Bust! Richard Feynman’s Last Journey. Physics Today, 44(10):114–115, October 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/?PTO/44/
114/1.

Dyson:1992:BRJ

[Dys92] Freeman J. Dyson. Book review: James Gleick, Genius: The Life and Science of Richard Feynman. Physics Today, 45(11):87, 1992. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://link.aip.org/link/?PTO/45/87/1. See re-

traction of one remark in [FD93].

Dyson:1993:FC

[Dys93] Freeman J. Dyson. Feynman at Cornell. In Brown and Rigden [BR93], pages 39–52. ISBN 0-88318-870-8. LCCN QC16.F49 A3 1993. Commentary by Joan Feynman, John Wheeler, Hans Bethe, Julian Schwinger, Murray Gell-Mann, Daniel Hillis, David Goodstein, Freeman Dyson, and Laurie Brown.

Dyson:1998:BRG

[Dys98] Freeman J. Dyson. Book review: Is God in the lab? The Meaning of It All: Thoughts of a Citizen Scientist, by Richard P. Feynman. Addison-Wesley, 133 pp., $22.00. Belief in God in an Age of Science, by John Polkinghorne. Yale University Press, 138 pp., $18.00. New York Review of Books, 45(9):8–??, May 28, 1998. ISSN 0028-7504 (print), 1944-7744 (electronic).

URL http://www.nybooks.com/articles/archives/1998/may/

28/is-god-in-the-lab/. See responses [?, Sta98].

Dyson:2001:RTT

[Dys01] Freeman J. Dyson. The radiation theories of Tomonaga, Schwinger, and Feynman (with commentary). In Hsu and Zhang [HZ01], pages 316–337. ISBN 981-02-4721-4. LCCN QC174.17.S9 H78 2001. URL

http://www.loc.gov/catdir/toc/fy031/2002280428.html.

Dyson:2003:RTT

[Dys03] Freeman J. Dyson. The radiation theories of Tomonaga, Schwinger, and Feynman. In Selected Papers on Quantum Electrodynamics [Sch03], page ?? ISBN 0-486-60444-6. LCCN QC680 .S35. Reprint of [Dys49].

Dyson:2005:WMB

[Dys05] Freeman J. Dyson. Wise man: Perfectly Reasonable Deviations from the Beaten Track: The Letters of Richard P. Feynman, edited and with an introduction by Michelle Feynman, with a foreword by Timothy Ferris. Basic Books, 486 pp., $26.00. New York Review of Books, 52(16):??, October 20, 2005. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/2005/oct/20/wise-man/.

Dyson:2011:BRD

[Dys11a] Freeman J. Dyson. Book reviews: The ‘dramatic picture’ of Richard Feynman: Quantum Man: Richard Feynman’s Life in Science, by Lawrence M. Krauss. Norton, 350 pp., $24.95. Feynman, by Jim Ottaviani, with art by Leland Myrick and coloring by Hilary Sycamore. First Second, 266 pp., $29.99. New York Review of Books, 58(12):39–40, July 14, 2011. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2011/jul/14/dramatic-picture-richard-feynman/.

Dyson:2011:HNR

[Dys11b] Freeman J. Dyson. Of historical note: Richard Feynman. The Institute Letter, Spring 2011. URL https://www.ias.edu/ideas/

2011/dyson-of-historical-note.

Dyson:2018:MPA

[Dys18] Freeman J. Dyson. Maker of patterns: an autobiography through letters. Liveright Publishing Corporation, New York, NY, USA, 2018. ISBN 0-87140-386-2 (hardcover). xvi + 400 pp. LCCN QC16.D95 A3 2018.

Edson:1967:TMS

[Eds67] Lee Edson. Two men in search of the quark. New York Times, ?? (??):54, 56, 64, 68, 72, October 8, 1967. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.

proquest.com/hnpnewyorktimes/docview/117894893/. See let-

ter rebutting Feynman’s involvement with quarks [Fey67b].

Edwards:1985:ADF

[Edw85] David A. Edwards. Atomic discourse in The Feynman Lectures on Physics. Synthese, 65(3):445–480, December 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869279.

Edgar:1962:MER

[EFK^{+}62] R. S. Edgar, Richard P. Feynman, S. Klein, I. Lielausis, and C. M. Steinberg. Mapping experiments with r mutants of bacteriophage T4d. Genetics, 47(2):179–186, February 1962. CODEN GENTAE. ISSN 0016-6731 (print), 1943-

2631 (electronic). URL http://www.pubmedcentral.nih.gov/
articlerender.fcgi?tool=pmcentrez&artid=1210321.

Erwin:2003:GPR

[EFV03] Charlotte E. Erwin, Carol Finerman, and David A. Valone. Guide to the papers of Richard Phillips Feynman, 1933–1988. Report, California Institute of Technology, Pasadena, CA,

USA, 2003. URL http://cdn.calisphere.org/data/13030/k0/
kt5n39p6k0/files/kt5n39p6k0.pdf.

Epstein:1955:DFD

[Eps55] Saul T. Epstein. Derivation of the Feynman–Dyson rules from time-independent theory. Physical Review, 98(??):196–198, April 1, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://link.aps.org/doi/10.1103/PhysRev.
98.196.

Fainman:1989:RAa

[Faĭ89] Richard Faĭnman. The reason for antiparticles. Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 31(64)(4):261–281 (1990), 1989. CODEN FMBMAC. ISSN 0015-3265. Translated from the Russian by D. Vachov.

Feynman:2005:FTN

[FBD05] Richard P. (Richard Phillips) Feynman, Laurie M. Brown, and P. A. M. (Paul Adrien Maurice) Dirac. Feynman’s Thesis: a New Approach to Quantum Theory. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2005. ISBN 981-256-366-0, 981-256-380-6 (paperback). xxii + 119 pp. LCCN QC174.12 .F48 2005.

Feynman:1955:CRS

[FC55] Richard P. Feynman and Michael Cohen. The character of the roton state in liquid helium. Progress of Theoretical Physics, 14(3):261– 262, September 1955. CODEN PTPKAV. ISSN 0033-068X (print),

1347-4081 (electronic). URL http://ptp.ipap.jp/link?PTP/14/
261/.

Feynman:1956:ESE

[FC56] R. P. Feynman and Michael Cohen. Energy spectrum of the excitations in liquid helium. Physical Review (2), 102(5):1189–1204, June 1, 1956. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Fernandez:1982:TDH

[FC82] Francisco M. Fernández and Eduardo A. Castro. Trivial offdiagonal Hellmann–Feynman formulas. International Journal of Quantum Chemistry, 21(3):671–672, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Feynman:1993:LET

[FD93] Joan Feynman and Freeman Dyson. Letters to the editor: The truth about Feynman’s father. Physics Today, 46(5):91, 1993. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/?PTO/46/91/1. Dyson retracts

a mistaken in a remark about the father of Joan and Richard Feynman that he made in his review [Dys92].

Feynman:1955:DNE

[FdHS55] Richard P. Feynman, F. de Hoffmann, and R. Serber. Dispersion of the neutron emission in U235 fission. Report OSTI 4354998, Los Alamos Scientific Laboratory, Atomic Energy Commission, Los Alamos, NM, USA, 1955.

Feynman:1956:DNE

[FdHS56] Richard P. Feynman, F. de Hoffmann, and R. Serber. Dispersion of the neutron emission in U-235 fission. Journal of Nuclear Energy (1954), 3(1–2):64–66, IN9, 68–69, August 1956. CODEN ???? ISSN 0891-3919 (print), 1878-2051 (electronic).

Fermi:1952:LF

[Fer52] Enrico Fermi. Letter to Feynman, January 18, 1952. Reprinted in [?, pages 844–846].

Feynman:1933:CPM

[Fey33] Richard P. Feynman. The calculus for the practical man. Highschool notebook., 1933. Copy in CalTech archive. Original in Niels Bohr Library, American Institute of Physics.

Feynman:1939:FM

[Fey39a] R. P. Feynman. Forces in molecules. Physical Review (2), 56(4): 340–343, August 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). This article, written by the author when he was a 21-year-old undergraduate physics student at MIT working under Chairman and Professor John Clarke Slater, contains an independent rediscovery of a famous theorem in quantum chemistry. The connection to the first discovery [Hel37] was found only later. Slater’s books refer to this as the Feynman Theorem, or the Feynman–Hellman Theorem, but most later authors call it the Hellmann–Feynman Theorem, crediting the discoverers in order of scientific priority.

Feynman:1939:FSM

[Fey39b] Richard Phillips Feynman. Forces and stresses in molecules. Thesis (b.s.), MIT, Cambridge, MA, USA, 1939. 30 pp.

Feynman:1942:PLA

[Fey42] Richard Feynman. The Principle of Least Action in Quantum Mechanics. Thesis (Ph.D.), Department of Physics, Princeton University, Princeton, NJ, USA, 1942.

Feynman:1946:TAF

[Fey46] Richard P. Feynman. A theorem and its application to finite tampers. Report OSTI 4341197, Los Alamos Scientific Laboratory, Atomic Energy Commission, Los Alamos, NM, USA, 1946.

Feynman:1947:WWF

[Fey47] Richard P. Feynman. What will the fundamental particles turn out to be? In G. T. Reynolds and Donald R. Hamilton, editors, The Future of Nuclear Science, page ?? Princeton University Press, Princeton, NJ, USA, 1947.

Feynman:1948:PT

[Fey48a] R. P. Feynman. Pocono conference. Physics Today, 1(2):8–10, June 1948. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Feynman:1948:RCC

[Fey48b] R. P. Feynman. A relativistic cut-off for classical electrodynamics. Physical Review (2), 74(8):939–946, October 15, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1948:STA

[Fey48c] R. P. Feynman. Space–time approach to non-relativistic quantum mechanics. Reviews of Modern Physics, 20(2):367–387, April 1, 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Feynman:1948:RCQ

[Fey48d] Richard P. Feynman. Relativistic cut-off for quantum electrodynamics. Physical Review (2), 74(10):1430–1438, November 15, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1949:STA

[Fey49a] R. P. Feynman. Space–time approach to quantum electrodynamics. Physical Review (2), 76(6):769–789, September 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1949:TTP

[Fey49b] R. P. Feynman. T5. The theory of positrons. Physical Review (2), 75(8):1321–1322, April 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL https://journals.

aps.org/pr/pdf/10.1103/PhysRev.75.1279. Brief contribution

to the Minutes of the 1948 Annual Meeting at New York, January 26–29, 1949.

Feynman:1949:TP

[Fey49c] R. P. Feynman. The theory of positrons. Physical Review (2), 76 (6):749–759, September 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1950:MFQ

[Fey50] R. P. Feynman. Mathematical formulation of the quantum theory of electromagnetic interaction. Physical Review (2), 80(3):440–457, November 1, 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1951:CPQ

[Fey51a] Richard P. Feynman. The concept of probability in quantum mechanics. In Proceedings of the Second Berkeley Symposium on Mathematical and Statistical Probability, California July 31–August 12, 1950, pages 533–541. University of California Press, Berke-

ley, CA, USA, 1951. URL http://projecteuclid.org/euclid.
bsmsp/1200500252.

Feynman:1951:EAC

[Fey51b] Richard P. Feynman. Estado atual dos conhecimentos sobre os mésons. (Portuguese) [current state of knowledge about mesons]. In Ciéncia e Cultura, III Annual Meeting of the Brazilian Society for the Advancement of Science, Belo Horizonte (MG), November, 1951, page ?? ????, ????, 1951. LCCN ???? Abstract only.

Feynman:1951:OCH

[Fey51c] Richard P. Feynman. An operator calculus having applications in quantum electrodynamics. Physical Review (2), 84(1):108–128, October 1, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1951:NNP

[Fey51d] Richard P. Feynman. Uma nova notaçao para o cálculo operacional. (Portuguese) [A new notation for the operational calculus]. In Ciéncia e Cultura, III Annual Meeting of the Brazilian Society for the Advancement of Science, Belo Horizonte (MG), November, 1951, pages 301–?? ????, ????, 1951. LCCN ???? Abstract only.

Feynman:1953:TLH

[Fey53a] R. P. Feynman. The λ-transition in liquid helium. Physical Review (2), 90(6):1116–1117, June 15, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1953:ATLb

[Fey53b] R. P. Feynman. Atomic theory of liquid helium near absolute zero. Physical Review (2), 91(6):1301–1308, September 15, 1953. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1953:ATT

[Fey53c] R. P. Feynman. Atomic theory of the λ transition in helium. Physical Review (2), 91(6):1291–1301, September 15, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1953:ATLa

[Fey53d] Richard P. Feynman. Atomic theory of liquid helium. In I. Imai, editor, Proceedings of the International Conference on Theoretical Physics, Kyoto and Tokyo, Japan, September 1953, pages 895– 901. Wiley, New York, NY, USA, 1953. LCCN ???? Published as ZAMM — Journal of Applied Mathematics and Mechanics / Zeitschrift für Angewandte Mathematik und Mechanik Volume 36, Issue 1–2, page 79, 1956.

Feynman:1953:ATLc

[Fey53e] Richard P. Feynman. Atomic theory of liquid helium. Notas de Física 12, Centro Brasileiro de Pesquisas Físicas, Av. Wenceslau Braz, 71, Rio de Janeiro, Brazil, 1953.

Feynman:1953:FSN

[Fey53f] Richard P. Feynman. Física Nuclear Téorica. (Portuguese) [Theoretical Nuclear physics]. Centro Brasileiro de Pesquisas Físicas, Av. Wenceslau Braz, 71, Rio de Janeiro, Brazil, 1953. ???? pp. LCCN ???? Edited by G. Rawitscher (CBPF). Republished in [?].

Feynman:1954:ATT

[Fey54a] R. P. Feynman. Atomic theory of the two-fluid model of liquid helium. Physical Review (2), 94(2):262–277, April 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1954:PSF

[Fey54b] R. P. Feynman. The present situation in fundamental theoretical physics. Anais Acad. Brasil. Ci., 26:51–59, 1954. CODEN ???? ISSN 0001-3765 (print), 1678-2690 (electronic).

Feynman:1955:SEP

[Fey55a] R. P. Feynman. Slow electrons in a polar crystal. Physical Review (2), 97(3):660–665, February 2, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1955:AQM

[Fey55b] Richard P. Feynman. Application of quantum mechanics to liquid helium. Progress in Low Temperature Physics, 1:17–53, 1955. CODEN PLTPAA. ISBN 0-444-53307-9. ISSN 0079-6417.

Feynman:1955:HIR

[Fey55c] Richard P. Feynman. Helium II in rotational flow. Science, 121 (3171):622–??, October 7, 1955. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Feynman:1955:VS

[Fey55d] Richard P. Feynman. The value of science. Engineering and Science (Caltech), 19(3):13–15, December 1955. CODEN ???? ISSN

0013-7812. URL http://resolver.caltech.edu/CaltechES:19.
3.Science.

Feynman:1956:RSR

[Fey56a] Richard P. Feynman. The relation of science and religion. Engineering and Science (Caltech), 19(9):20–23, June 1956. CODEN

???? ISSN 0013-7812. URL http://resolver.caltech.edu/

CaltechES:19.9.Religion. See letter and response [SF56].

Feynman:1956:SOC

[Fey56b] Richard P. Feynman. Science and the open channel. Science, 123 (3191):307–??, February 24, 1956. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Feynman:1957:SS

[Fey57a] R. P. Feynman. Superfluidity and superconductivity. Reviews of Modern Physics, 29(2):205–212, April 1, 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Feynman:1957:HMP

[Fey57b] Richard P. Feynman. An historic moment in physics. Engineering and Science (Caltech), 20(5):17–18, ???? 1957. CODEN ???? ISSN

0013-7812. URL http://resolver.caltech.edu/CaltechES:20.
5.feynman.

Feynman:1957:RSW

[Fey57c] Richard P. Feynman. The role of science in the world today. Proceedings of the Institute of World Affairs, 33(??):17–31, 1957. CO-DEN ???? ISSN ????

Feynman:1958:ELH

[Fey58] Richard P. Feynman. Excitations in liquid helium. Physica, 24(1): S18–S26, September 1958. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic). Kammerling Onnes Conference, Leiden, The Netherlands.

Feynman:1959:PRBa

[Fey59a] Richard P. Feynman. Plenty of room at the bottom. Presentation to American Physical Society, 1959. URL http://www.its.caltech.

edu/~feynman/plenty.html.

Feynman:1959:PRBb

[Fey59b] Richard P. Feynman. Plenty of room at the bottom. Engineering and Science (Caltech), 23(5):22–36, February 1959. CODEN

???? ISSN 0013-7812. URL http://resolver.caltech.edu/
CaltechES:23.5.1960Bottom.

Feynman:1959:SLT

[Fey59c] Richard P. Feynman. Series of lectures on the theory of fundamental processes. California Institute of Technology, Pasadena, CA, USA, 1959.

Feynman:1961:QE

[Fey61a] R. P. Feynman. Quantum electrodynamics. Frontiers in Physics: Lecture Note and Reprint Volume. W. A. Benjamin, Inc., New York, NY, USA, 1961. x + 198 pp. Notes corrected by E. R. Huggins and H. T. Yura.

Feynman:1961:TFP

[Fey61b] R. P. Feynman. The theory of fundamental processes. Frontiers in Physics: A Lecture Note and Reprint Volume. W. A. Benjamin, Inc., New York, NY, USA, 1961. x + 172 pp. Notes Corrected by H. T. Yura.

Feynman:1961:TPR

[Fey61c] R. P. Feynman. There’s plenty of room at the bottom. In Horace D. Gilbert, editor, Miniaturization, pages 282–296. Reinhold Publishing Corporation, New York, NY, USA, 1961. LCCN TK7870 .G52.

Feynman:1961:PSQ

[Fey61d] Richard P. Feynman. The present status of quantum electrodynamics. In ????, editor, La théorie quantique des champs. Rapports et Discussion du Douzième Conseil de Physique tenu à l’Université libre de Bruxelles, 9–14 Octobre 1961, page ?? Interscience Publishers and R. Stoops, New York, NY, USA; Brussels, Belgium, 1961. LCCN ????

Feynman:1962:QELb

[Fey62a] R. P. Feynman. Quantum electrodynamics. A lecture note and reprint volume. Frontiers in Physics. W. A. Benjamin, Inc., New York, NY, USA, 1962. xi + 198 pp. Notes taken by A. R. Hibbs. Notes corrected by E. R. Huggins and H. T. Yura. Second printing corrected by P. Cziffra.

Feynman:1962:TFP

[Fey62b] R. P. Feynman. The theory of fundamental processes. A lecture note volume. Frontiers in Physics. W. A. Benjamin, Inc., New York, NY, USA, 1962. x + 172 pp. Notes taken by P. A. Curruthers and M. Nauenberg. Notes corrected by H. T. Yura. Second printing corrected by P. Cziffra.

Feynman:1962:QELa

[Fey62c] Richard P. Feynman. Lectures in elementary physics. California Institute of Technology, Pasadena, CA, USA, 1962.

Feynman:1963:QTG

[Fey63a] R. P. Feynman. The quantum theory of gravitation. Acta Physica Polonica, 24:697–722, 1963. CODEN APPOAK. ISSN 0001-673X.

Feynman:1963:FNE

[Fey63b] Richard P. Feynman. Fenômenos no Estado Sólido. (Portuguese) [Phenomena in Solid State]. Monografias de Física. Centro Brasileiro de Pesquisas Físicas, Rio de Janeiro, Brazil, 1963. ???? pp. LCCN ????

Feynman:1963:PTP

[Fey63c] Richard P. Feynman. The problem of teaching physics in Latin America. Engineering and Science (Caltech), 27(2):21–30, ???? 1963. CODEN ???? ISSN 0013-7812. URL http://resolver.

caltech.edu/CaltechES:27.2.LatinAmerica.

Feynman:1964:LPE

[Fey64a] R. P. Feynman. Lectures on physics. Exercises. Addison-Wesley, Reading, MA, USA, 1964. 41 pp. Chapter 50.

Feynman:1964:QTGa

[Fey64b] Richard P. Feynman. The quantum theory of gravitation. In Leopold Infeld, editor, Proceedings on Theory of Gravitation: conference in Warszawa and Jabłonna, 25–31 July, 1962 = Conférence internationale sur les théories relativistes de la gravitation, page ?? Gauthier-Villars, Paris, France, 1964. LCCN ????

Feynman:1964:QTGb

[Fey64c] Richard P. Feynman. The quantum theory of gravitation. In ????, editor, Proceedings on Theory of Gravitation, pages 207–208. PWN-Polish Scientific Publisher, ????, Poland, 1964. LCCN ????

Feynman:1964:RPO

[Fey64d] Richard P. Feynman. The relation of physics to other sciences. The Physics Teacher, 2(3):111–117, March 1964. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Feynman:1964:RPF

[Fey64e] Richard P. Feynman. Richard P. Feynman, The Nobel Prize in Physics 1965: Biography. In Nobel Foundation [Nob72], page ?? ISBN 0-444-40993-9. LCCN QC71.N735. URL http://

nobelprize.org/nobel_prizes/physics/laureates/1965/feynman-
bio.html;   http://nobelprize.org/nobel_prizes/physics/
laureates/1965/feynman-lecture.html.

Feynman:1964:STR

[Fey64f] Richard P. Feynman. The Special Theory of Relativity. The Physics Teacher, 2(4):151–157, April 1964. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Feynman:1965:CSW

[Fey65a] Richard P. Feynman. Consequences of SU(3) symmetry in weak interactions. In Zichichi [Zic65], pages 111–174. LCCN QC721 .I578 1964.

Feynman:1965:NTNa

[Fey65b] Richard P. Feynman. New textbooks for the “new” mathematics. Engineering and Science (Caltech), 28(6):9–15, March 1965. CODEN ???? ISSN 0013-7812. URL http://resolver.caltech.

edu/CaltechES:28.6.feynman.

Feynman:1965:NTNb

[Fey65c] Richard P. Feynman. New textbooks for the “new” mathematics. The California Institute of Technology Quarterly, 6(??):2–9, Spring 1965. CODEN ???? ISSN ????

Feynman:1965:PSS

[Fey65d] Richard P. Feynman. Present status in strong, electromagnetic and weak interactions. In Zichichi [Zic65], pages 366–418. LCCN QC721 .I578 1964.

Feynman:1965:CPL

[Fey65e] Richard P. (Richard Phillips) Feynman. The character of physical law, volume 66 of The Messenger lectures, 1964; The MIT Press paperback series. MIT Press, Cambridge, MA, USA, 1965. 173 pp. LCCN QC71 .F44 1965b.

Feynman:1965:PUQ

[Fey65f] Richard P. (Richard Phillips) Feynman, editor. Probability and uncertainty: the quantum mechanical view of nature (Motion picture). Education Development Center, British Broadcasting Corp., London, UK, 1965. LCCN ???? 56 minute B&W film.

Feynman:1966:DSTa

[Fey66a] Richard P. Feynman. The development of the space–time view of quantum electrodynamics. In ????, editor, Les Prix Nobel 1965, pages 172–191. Imprimerie Royale P. A. Norstedt & Söner, Stockholm, Sweden, 1966. LCCN ???? URL http://www.

feynmanlectures.info/other/Feynmans_Nobel_Lecture.pdf.

Feynman:1966:DSTb

[Fey66b] Richard P. Feynman. The development of the space–time view of quantum electrodynamics. Physics Today, 19(8):31–44, August 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Feynman:1966:DSTc

[Fey66c] Richard P. Feynman. The development of the space–time view of quantum electrodynamics. Science, 153(3737):699–708, August 12, 1966. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Feynman:1966:SPL

[Fey66d] Richard P. Feynman. Symmetry in physical laws. The Physics Teacher, 4(4):161–174, April 1966. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Feynman:1966:WWS

[Fey66e] Richard P. Feynman. What is and what should be the role of scientific culture in modern society? Supp. al Nuovo Cimento, 4 (??):492–524, 1966. CODEN ???? ISSN 0029-6341 (print), 1827-6121 (electronic).

Feynman:1967:FTG

[Fey67a] Richard P. Feynman. Field theory as a guide to the strong interactions. In C. R. Hagen, G. Guralnik, and V. S. Mathur, editors, Proceedings of the 1967 International Conference on Particles and Fields, page ?? Interscience Publishers, New York, NY, USA, 1967. LCCN ????

Feynman:1967:LEQ

[Fey67b] Richard P. Feynman. Letter to the Editor: Quarks. New York Times, ??(??):??, November 5, 1967. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://

search.proquest.com/hnpnewyorktimes/docview/118036235/.

See [Eds67].

Feynman:1969:AMM

[Fey69a] Richard P. Feynman. The application of mathematics to mathematics. American Mathematical Monthly, 76(10):1178–1179, December 1969. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Feynman:1969:BHC

[Fey69b] Richard P. Feynman. The behavior of hadron collisions at extreme energies. In Chen Ning Yang et al., editors, High Energy Collisions, pages 237–256. Gordon and Breach, New York, NY, USA, 1969. LCCN ????

Feynman:1969:PSS

[Fey69c] Richard P. Feynman. Present status of strong, electromagnetic, and weak interactions. Československý Časopis pro Fysiku, A19 (??):47–59, 1969. CODEN ???? ISSN ????

Feynman:1969:VHE

[Fey69d] Richard P. Feynman. Very high-energy collisions of hadrons. Physical Review Letters, 23(24):1415–1417, December 15, 1969. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Feynman:1969:WS

[Fey69e] Richard P. Feynman. What is science? The Physics Teacher, 7(6): 313–320, September 1969. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Feynman:1970:FLP

[Fey70a] Richard P. Feynman. The Feynman Lectures on Physics: The Definitive and Extended Edition. Addison-Wesley, Reading, MA, USA, 1970. ISBN 0-8053-9045-6. ???? pp. LCCN ???? Includes Feynman’s Tips on Physics (with Michael Gottlieb and Ralph Leighton), which includes four previously unreleased lectures on problem solving, exercises by Robert Leighton and Rochus Vogt, and a historical essay by Matthew Sands.

Feynman:1970:P

[Fey70b] Richard P. Feynman. Partons. In E. C. G. Sudarshan and Y. Ne’eman, editors, The Past Decade in Particle Theory, pages 773–813. Gordon and Breach, New York, NY, USA, 1970. LCCN ????

Feynman:1970:CPL

[Fey70c] Richard Phillips Feynman, editor. The character of physical law. The Messenger Lectures. MIT Press, Cambridge, MA, USA, 1970.

ISBN 0-262-06016-7 (hardcover), 0-262-56003-8 (paperback). 173 pp. LCCN ????

Feynman:1971:CLT

[Fey71a] Richard P. Feynman. Closed loop and tree diagrams. In Klauder [Kla72], page ?? ISBN 0-7167-0337-8. LCCN QC173 .M313.

Feynman:1971:LG

[Fey71b] Richard P. Feynman. Lectures on gravitation. California Institute of Technology, Pasadena, CA, USA, 1971.

Feynman:1971:PQG

[Fey71c] Richard P. Feynman. Problems in quantizing the gravitational field and the massless Yang–Mills field. In Klauder [Kla72], page ?? ISBN 0-7167-0337-8. LCCN QC173 .M313.

Feynman:1972:FAE

[Fey72a] Richard P. Feynman. Fisica de altas energias. Cursos de Verano 1972. Mexico: COPAA-SEDICT, 1972.

Feynman:1972:WNC

[Fey72b] Richard P. Feynman. What neutrinos can tell us about partons. In Andor Frenkel and George Marx, editors, Proceedings of the Neutrino ’72 Europhysics Conference, Balatonfüred, Hungary, 11– 17 June 1972, volume II, pages 75–96. OMKD Technoinform, Budapest, Hungary, 1972. ISBN ???? LCCN ????

Feynman:1972:PHI

[Fey72c] Richard P. (Richard Phillips) Feynman. Photon-hadron interactions. Frontiers in physics. W. A. Benjamin, Inc., New York, NY, USA, 1972. ISBN 0-8053-2510-7, 0-8053-2511-5 (paperback). xvi + 282 pp. LCCN QC794 .F45 1972.

Feynman:1972:SMS

[Fey72d] Richard P. (Richard Phillips) Feynman. Statistical mechanics: a set of lectures by R. P. Feynman. Frontiers in physics. W. A. Benjamin, Inc., New York, NY, USA, 1972. ISBN 0-8053-2508-5, 0-8053-2509-3 (paperback). xii + 354 pp. LCCN QC174.8 .F48. Notes taken by R. Kikuchi and H. A. Feiveson. Edited by Jacob Shaham.

Feynman:1973:Q

[Fey73] Richard P. Feynman. Quarks. Fizikai Szemle (Budapest), 23(??): 1–7, 1973. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Feynman:1974:CS

[Fey74a] R. P. Feynman. Conference summary. AIP Conference Proceedings, 22(1):299–327, July 1, 1974. CODEN APCPCS. ISSN ???? URL

http://link.aip.org/link/?APCPCS/22/299/1.

Feynman:1974:CCS

[Fey74b] Richard P. Feynman. Cargo cult science. Engineering and Science (Caltech), 37(7):10–13, June 1974. CODEN ???? ISSN 0013-

7812. URL http://calteches.library.caltech.edu/51/02/
CargoCult.pdf;   http://resolver.caltech.edu/CaltechES:
28.6.feynman.

Feynman:1974:AB

[Fey74c] Richard P. Feynman. Los Alamos from below. Engineering and Science (Caltech), 39(2):11–30, January/February 1974. CODEN

???? ISSN 0013-7812. URL http://resolver.caltech.edu/
CaltechES:39.2.FeynmanLosAlamos.

Feynman:1974:P

[Fey74d] Richard P. Feynman. Partons. In P. N. Dobson, V. Z. Petersen, and S. F. Tuan, editors, Proceedings of the 5th Hawaii Topical Conference in Particle Physics, University of Hawaii, 8–21 August, 1973, pages 1–97. University Press of Hawaii, Honolulu, Hawaii, USA, 1974. ISBN ???? LCCN ????

Feynman:1974:SP

[Fey74e] Richard P. Feynman. Structure of the proton. Science, 183(4125): 601–610, February 15, 1974. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Feynman:1974:X

[Fey74f] Richard P. Feynman. xxx. In Peter Norman Dobson, V. Z. (Vincent Zetterberg) Peterson, and San Fu Tuan, editors, Proceedings. Hawaii Topical Conference in Particle Physics (5th: 1973: University of Hawaii), pages ix + 710. University Press of Hawaii, Honolulu, HI, USA, 1974. ISBN 0-8248-0327-2. LCCN QC794.8.S8 H38 1973.

Feynman:1977:CHC

[Fey77a] Richard P. Feynman. Correlations in hadron collisions at high transverse momentum. In Behram Kurşunoğlu, Arnold Perlmutter, and Linda F. Scott, editors, Deeper Pathways in High Energy Physics: Proceedings of a Conference Held by the Center for Theoretical Studies, University of Miami, Coral Gables, Florida, January 17–21, 1977, Studies in the natural sciences, pages 461–488. Plenum Press, New York, NY, USA; London, UK, 1977. ISBN 0-306-36912-5. LCCN QC793 .O7 1977.

Feynman:1977:GT

[Fey77b] Richard P. Feynman. Gauge theories. In Weak and electromagnetic interactions at high energy (École d’Été Phys. Théor., XXIX, Les Houches, 1976), pages 121–204. North-Holland Publishing Co., Amsterdam, The Netherlands, 1977.

Feynman:1977:QJ

[Fey77c] Richard P. Feynman. Quark jets. In ????, editor, Proceedings of the 8th International Symposium on Multiparticle Dynamics, Kayserberg, 12–17 June 1977, page ?? Université Louis Pasteur de Strasbourg. Centre de recherches nucléaires. High Energy Division, Strasbourg, France, 1977. ISBN ???? LCCN ????

Feynman:1978:SMK

[Fey78] Richard P. Feynman. Statisticheskaĭa mekhanika: kurs lektsiĭ. (Russian) [Statistical Mechanics Lecture Course]. Izdatelstvo Mir, Moscow, USSR, second edition, 1978. 407 pp. LCCN QC174.82.F43517 1978. Translation to Russian from English by Yu. G. Rudogo.

Feynman:1980:TFR

[Fey80a] R. Feynman. Talking to Richard Feynman — a genius interprets fantasy and creativity. Recherche, 11(117):1424–1428, ???? 1980. ISSN 0029-5671 (print), 1625-9955 (electronic).

Feynman:1980:AB

[Fey80b] Richard P. Feynman. Los Alamos from below. In Badash et al. [BHB80], pages 111–119, 129. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

Feynman:1981:QBY

[Fey81] Richard P. Feynman. The qualitative behavior of Yang–Mills theory in 2 + 1 dimensions. Nuclear Physics B, 188(3):479–512, October 5, 1981. CODEN NUPBBO. ISSN 0550-3213 (print),

1873-1562 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0550321381900055.

Feynman:1982:AB

[Fey82a] R. P. Feynman. Los Alamos from below. Annals of the History of Computing, 4(1):61–63, ???? 1982. CODEN AHCOE5. ISSN 0164-1239.

Feynman:1982:P

[Fey82b] Richard P. Feynman. Partons. In S. Pakvasa and San Fu Tuan, editors, Hawaii Topical Conference in Particle Physics, volume I (Selected Lectures), pages 229–424. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1982. ISBN 9971-950-36-7, 9971-950-16-2. LCCN ????

Feynman:1982:QDQ

[Fey82c] Richard P. Feynman. A qualitative discussion of quantum chromodynamics in 2 + 1 dimensions. In J. Dias de Deus and Jacques Soffer, editors, Proceedings of the International Conference on High Energy Physics, Lisbon, July 9–15, 1981, pages 660–683. European Physical Society, Geneva, Switzerland, 1982. ISBN ???? LCCN ????

Feynman:1982:SPC

[Fey82d] Richard P. Feynman. Simulating physics with computers. International Journal of Theoretical Physics, 21(6–7):467–488, 1982. CO-DEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic). Physics of computation, Part II (Dedham, Mass., 1981).

Feynman:1984:QMC

[Fey84] R. P. Feynman. Quantum-mechanical computers. Journal of the Optical Society of America. B, Optical physics, 1(3):464, ???? 1984. CODEN JOBPDE. ISSN 0740-3224 (print), 1520-8540 (electronic).

Feynman:1985:CMF

[Fey85a] Richard P. Feynman. The computing machines in the future. In Nishina Memorial Foundation, editor, Nishina Memorial Lectures: Creators of Modern Physics, volume 746 of Lecture Notes in Physics, pages 99–114. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 4-431-77056-9. LCCN ?QC71 .N57 2008.

Feynman:1985:QMC

[Fey85b] Richard P. Feynman. Quantum mechanical computers. Optics News, 11(2):11–20, February 1985. CODEN ONEWDU.

ISSN 0098-907X. URL http://www.opticsinfobase.org/on/
abstract.cfm?URI=on-11-2-11.

Feynman:1986:QMCb

[Fey86a] R. P. Feynman. Quantum mechanical computers. Uspekhi Fizicheskikh Nauk, 149(4):671–688, ???? 1986. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Feynman:1986:SYJ

[Fey86b] R. P. Feynman. Surely you’re joking, Feynman (fragments of the book on physics teaching). Uspekhi Fizicheskikh Nauk, 148(3):509– 526, ???? 1986. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Feynman:1986:QMCa

[Fey86c] Richard P. Feynman. Quantum mechanical computers. Foundations of Physics, 16(6):507–531, 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Feynman:1986:RCR

[Fey86d] Richard P. Feynman. Rogers Commission report, volume 2 appendix F – personal observations on reliability of shuttle. Report, NASA, Washington, DC, USA, 1986. URL http://history.nasa.

gov/rogersrep/v2appf.htm.

Feynman:1986:WNG

[Fey86e] Richard P. Feynman. Was ist Naturwissenschaft?. (German) [What is science?]. Physik und Didaktik, 15(2):105–116, 1986. CODEN

???? ISSN 0340-8515. URL http://www.fachportalpaedagogik.
de/fis_bildung/suche/fis_set.html?FId=2299.

Feynman:1987:NP

[Fey87a] Richard P. Feynman. Negative probability. In Quantum implications: Essays in Honor of David Bohm, pages 235–248. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1987.

Feynman:1987:TCO

[Fey87b] Richard P. Feynman. Tiny computers obeying quantum mechanical laws. In Metropolis et al. [MKR87], pages 7–25. ISBN 0-12-492155-8. LCCN QC44 .N49 1987. URL http://www.osti.

gov/energycitations/product.biblio.jsp?osti_id=6120718&
query_id=0.

Feynman:1988:DAV

[Fey88a] Richard P. Feynman. Difficulties in applying the variational principle to quantum field theories. In L. Polley and D. E. L. Pottinger, editors, Proceedings of the International Workshop on Variational Calculus in Quantum Field Theory, Wangerooge, West Germany, September 1–4, 1987, page ?? World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1988. ISBN ???? LCCN ????

Feynman:1988:SOT

[Fey88b] Richard P. Feynman. It’s as simple as one, two, three, . . .. Engineering and Science (Caltech), 52(1):11–17, ???? 1988. CODEN

???? ISSN 0013-7812. URL http://resolver.caltech.edu/
CaltechES:52.1.Feynman.

Feynman:1988:OIV

[Fey88c] Richard P. Feynman. An outsider’s inside view of the Challenger Inquiry. Physics Today, 41(2):26–37, February 1988. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Feynman:1988:QST

[Fey88d] Richard Phillips Feynman, editor. QED, the strange theory of light and matter. Alix G. Mautner memorial lectures. Princeton University Press, Princeton, NJ, USA, 1988. ISBN 0-691-08388-6, 0-691-02417-0 (paperback). x + 158 pp. LCCN QC793.5.P422 F48 1988.

Feynman:1989:RFA

[Fey89a] Michelle Feynman. Richard Feynman, artist. Physics Today, 42 (2):86–87, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Feynman:1989:RAb

[Fey89b] R. P. Feynman. The reason for antiparticles. Uspekhi Fizicheskikh Nauk, 157(1):163–183, ???? 1989. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Feynman:1989:FNE

[Fey89c] Richard P. Feynman. Fenômenos no Estado Sólido. (Portuguese) [Phenomena in Solid State]. Coleção Galileo — Textos de física, I, Documentos históricos. Centro Brasileiro de Pesquisas Físicas, Rio de Janeiro, Brazil, 1989. 215 pp. LCCN ???? Report CBPF-DH-001/89. Reprint of [Fey63b].

Feynman:1990:SI

[Fey90] R. P. Feynman. Scientific integrity. Journal of the Geological Society of India, 35(3):322, ???? 1990. ISSN 0016-7622.

Feynman:1991:AM

[Fey91a] Richard P. Feynman. Atoms in motion. In Ferris and Fadiman [FF91], pages 3–17. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Feynman:1991:DPF

[Fey91b] Richard P. Feynman. Distinction of past and future. In Ferris and Fadiman [FF91], pages 147–162. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Feynman:1992:TPR

[Fey92] R. P. Feynman. There’s plenty of room at the bottom [data storage]. Journal of Microelectromechanical Systems, 1(1):60–66, March 1992. CODEN JMIYET. ISSN 1057-7157 (print), 1941-0158 (electronic).

Feynman:1993:IM

[Fey93] R. Feynman. Infinitesimal machinery. Journal of Microelectromechanical Systems, 2(1):4–14, March 1993. CODEN JMIYET. ISSN 1057-7157 (print), 1941-0158 (electronic).

Feynman:1994:FRa

[Fey94a] R. P. Feynman. From Richard Feynman. American Mathematical Monthly, 101(6):543, ???? 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Feynman:1994:CPL

[Fey94b] Richard Phillips Feynman. The Character of Physical Law. Modern Library, New York, NY, USA, modern library edition, 1994. ISBN 0-679-60127-9. xx + 167 pp. LCCN QC71 .F44 1994.

Feynman:1995:TFP

[Fey95] Richard Phillips Feynman. The theory of fundamental processes: a lecture note volume. Frontiers in physics; a lecture note and reprint series. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-8053-2507-7. x + 171 pp. LCCN QC174.1 F4 1995.

Feynman:1997:SEP

[Fey97] Richard Phillips Feynman. Six Not-so-easy Pieces: Einstein’s Relativity, Symmetry, and Space–time. Helix books. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-32841-0 (set), 0-201-15025-5 (hardcover), 0-201-32842-9 (paperback), 0-201-31151-8 (disc 1), 0-201-31152-6 (disc 2), 0-201-31153-4 (disc 3), 0-201-31154-2 (disc 4), 0-201-31155-0 (disc 5), 0-201-31156-9 (disc 6). xxvii + 152 pp. LCCN QC793.3.S9 F49 199. Originally prepared for publication by Robert B. Leighton and Matthew L. (Matthew Linzee) Sands. New introduction by Roger Penrose. See also [FLS95].

Feynman:1998:BRP

[Fey98a] Richard P. Feynman. Biography: Richard P. Feynman. In Samuelsson and Sohlman [SS98], pages 179–180. ISBN 981-02-3404-X. LCCN QC71 .P455 1998.

Feynman:1998:DST

[Fey98b] Richard P. Feynman. The development of the space–time view of quantum electrodynamics. In Samuelsson and Sohlman [SS98], pages 155–178. ISBN 981-02-3404-X. LCCN QC71

.P455 1998. URL http://www.feynmanlectures.info/other/
Feynmans_Nobel_Lecture.pdf.

Feynman:1998:SMS

[Fey98c] Richard P. Feynman. Statistical mechanics. A set of lectures. Advanced Book Classics. Perseus Publishers, Cambridge, MA, USA, 1998. ISBN 0-201-36076-4. xiv + 354 pp. Reprint of the 1972 original.

Feynman:1998:MIA

[Fey98d] Richard Phillips Feynman. The Meaning of It All: Thoughts of a Citizen Scientist. Perseus Publishers, Cambridge, MA, USA, 1998. ISBN 0-7382-0166-9 (paperback). 133 pp. LCCN Q175.55 .F49 1998.

Feynman:19xx:HTS

[Feyxx] Richard P. Feynman. How trains stay on track. Online video

recording, 19xx. URL http://www.wimp.com/trainsstay/. See

more detailed analysis in [Sha17].

Feynman:2000:SGF

[Fey00a] R. Feynman. The “scholar,” genius and fantasy. Recherche, ?? (331):72–74, ???? 2000. ISSN 0029-5671 (print), 1625-9955 (electronic). Reprint of ????.

Feynman:2000:VVR

[Fey00b] Richard Feynman. Vous voulez rire, Monsieur Feynman ! Entretiens avec Ralph Leighton. Odile Jacob, 2000.

Feynman:2000:PFT

[Fey00c] Richard P. Feynman. The pleasure of finding things out: the best short works of Richard P. Feynman. Perseus Publishers, Cambridge, MA, USA, 2000. ISBN 0-7382-0349-1 (paperback). xvi + 270 pp. LCCN Q171 .F385 1999. Edited by Jeffrey Robbins, and foreword by Freeman Dyson.

Feynman:2005:PRD

[Fey05a] Michelle Feynman, editor. Perfectly Reasonable Deviations (from the Beaten Track): the Collected Letters of Richard P. Feynman. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0636-9. xxiii + 486 pp. LCCN QC16 .F49 A4 2005. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http://
www.loc.gov/catdir/toc/ecip055/2005000049.html. Foreword

by Timothy Ferris. See also interview with the editor [Ano05].

Feynman:2005:FST

[Fey05b] Richard P. Feynman. Física Nuclear Téorica. (Portuguese) [Theoretical Nuclear Physics]. Livraria da Física, São Paulo, Brazil, 2005. ISBN 85-88325-50-0. 133 pp. LCCN ???? Reprint of [Fey53f].

Feynman:2006:VBF

[Fey06a] Richard Feynman. The very best of the Feynman Lectures. Basic Books, New York, NY, USA, 2006.

Feynman:2006:QST

[Fey06b] Richard P. (Richard Phillips) Feynman. QED: the strange theory of light and matter. Princeton science library, Alix G. Mautner memorial lectures. Princeton University Press, Princeton, NJ, USA, expanded Princeton science library edition, 2006. ISBN 0-691-12717-4, 0-691-12575-9 (paperback). xxiv + 158 pp. LCCN QC793.5.P422 F48 2006.

Feynman:2007:QB

[Fey07a] Richard Feynman. Quantum behavior. In Whitfield and Hicks [WH07], page ?? ISBN 1-880323-91-5. LCCN QC7.5 .W53 2007.

URL http://store.greatbooks.org/adu-wtm.html. Foreword

by Alan Lightman.

Feynman:2007:US

[Fey07b] Richard Feynman. The uncertainty of science. In Whitfield and Hicks [WH07], page ?? ISBN 1-880323-91-5. LCCN QC7.5 .W53

2007. URL http://store.greatbooks.org/adu-wtm.html. Fore-

word by Alan Lightman.

Feynman:2010:FLPb

[Fey10a] Richard P. Feynman. The Feynman lectures on physics: new millennium edition. Basic Books, New York, NY, USA, 2010. ISBN 0-465-02414-9. ???? pp. LCCN ????

Feynman:2010:FLPa

[Fey10b] Richard P. (Richard Phillips) Feynman. The Feynman lectures on physics. Vol. 19, Feynman on masers and light. Vol. 20, Feynman on quantum mechanics and electromagnetism. Basic Books, New York, NY, USA, 2010. ISBN 0-7382-0933-3 (vol.19-20). 9 sound discs (ca. 9 hrs.) pp. LCCN SDB 98630.

Feynman:2010:WPG

[Fey10c] Richard Phillips Feynman. Vom Wesen physikalischer Gesetze, volume 1748 of Serie Piper. Piper, München, Germany, 2010. ISBN 3-492-21748-6. 215 pp. LCCN ????

Feynman:2011:WS

[Fey11a] Richard Feynman. What is science? Resonance, 16(9):860–873, September 2011. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic).

Feynman:2011:SPF

[Fey11b] Richard P. Feynman. Sechs physikalische Fingerübungen ; Physikalische Fingerübungen für Fortgeschrittene, volume 4999 of Serie Piper. Piper, Zürich, Switzerland, fourth edition, 2011. ISBN 3-492-24999-X. LCCN QC23 .F47214 2007.

Feynman:2011:SEPa

[Fey11c] Richard P. Feynman. Six easy pieces: essentials of physics explained by its most brilliant teacher. Basic Books, New York, NY, USA, 2011. ISBN 0-465-02527-7. ???? pp. LCCN ????

Feynman:2011:SEPb

[Fey11d] Richard P. Feynman. Six not-so-easy pieces: Einstein’s relativity, symmetry, and space–time. Basic Books, New York, NY, USA, 2011. ISBN 0-465-02526-9 (print), 0-465-02528-5 (e-book). xxvii + 154 pp. LCCN QC793.3.S9 F49 2011.

Feynman:2011:TPR

[Fey11e] Richard P. Feynman. There’s plenty of room at the bottom. Resonance, 16(9):890–905, September 2011. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL https://link.

springer.com/article/10.1007/s12045-011-0109-x.

Feynman:2012:CPL

[Fey12] Richard P. Feynman. The character of physical law (1965). In Wardhaugh [War12], pages 318–321. ISBN 0-691-14775-2 (hardcover). LCCN QA7 .W43 2012.

Feynman:2013:QST

[Fey13] Richard P. Feynman. QED: the Strange Theory of Light and Matter. Princeton Science Library. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-12575-9, 0-691-12717-4, 1-4008-4746-X. xxiv + 158 pp. LCCN QC793.5.P422 F48 2006. URL

http://press.princeton.edu/titles/8169.html.

Feynman:2015:AA

[Fey15] Joan Feynman. From auroras to anthropology. In Charman-

Anderson [CA15], page ?? URL https://findingada.com/shop/
a-passion-for-science-stories-of-discovery-and-invention/
joan-feynman-from-auroras-to-anthropology/.

Field:1977:QES

[FF77] R. D. Field and R. P. Feynman. Quark elastic scattering as a source of high-transverse-momentum mesons. Physical Review D (Particles and Fields), 15(9):2590–2616, May 1, 1977. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

Field:1978:PPQ

[FF78] R. D. Field and R. P. Feynman. A parameterization of the properties of quark jets. Nuclear Physics B, 136(1):1–76, April 17, 1978. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0550321378900159.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Feynman:1995:ARP

[FF95] Richard P. (Richard Phillips) Feynman and Michelle Feynman. The art of Richard P. Feynman: images by a curious character. GB Science Publishers SA, Basel, Switzerland, 1995. ISBN 2-88449-047-7. 173 pp. LCCN NC1075.F44 A4 1995.

Feynman:2005:DYT

[FF05] Richard P. (Richard Phillips) Feynman and Michelle Feynman. Don’t You Have Time to Think? Allen Lane, London, UK, 2005. ISBN 0-7139-9847-4 (hardcover). xxi + 486 pp. LCCN 06.E02646. US$20.00.

Feynman:1977:CAP

[FFF77] R. P. Feynman, R. D. Field, and G. C. Fox. Correlations among particles and jets produced with large transverse momenta. Nuclear Physics B, 128(1):1–65, September 19, 1977. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0550321377902991.

Feynman:1978:QCA

[FFF78] R. P. Feynman, R. D. Field, and G. C. Fox. Quantumchromodynamic approach for the large-transverse-momentum production of particles and jets. Physical Review D (Particles and Fields), 18(9):3320–3343, November 1, 1978. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500.

Feynman:1992:TFI

[FG92] R. P. Feynman and M. Gellmann. Theory of the Fermi interaction. Current Science, 63(2):76–81, July 25, 1992. CODEN CUSCAM. ISSN 0011-3891. Reprinted from Physical Review, 1958.

Feynman:2006:FTP

[FGL^{+}06] Richard P. (Richard Phillips) Feynman, Michael A. Gottlieb, Ralph Leighton, Matthew L. (Matthew Linzee) Sands, Robert B. Leighton, and Rochus Vogt. Feynman’s tips on physics: a problemsolving supplement to the Feynman Lectures on Physics. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-8053-9063-4 (paperback). xiii + 162 pp. LCCN QC23 .F47 1989 Suppl. URL

http://www.loc.gov/catdir/toc/ecip0512/2005013077.html.

Feynman:2012:FTP

[FGL^{+}12] Richard P. (Richard Phillips) Feynman, Michael A. Gottlieb, Ralph Leighton, Matthew L. (Matthew Linzee) Sands, Robert B. Leighton, and Rochus Vogt, editors. Feynman’s tips on physics: reflections, advice, insights, practice. Basic Books, New York, NY, USA, second edition, 2012. ISBN 0-465-02797-0 (hardcover), 0-465-02921-3 (e-book). xiv + 182 pp. LCCN QC23 .F47 2013.

Feynman:1958:TFI

[FGM58] R. P. Feynman and M. Gell-Mann. Theory of the Fermi interaction. Physical Review (2), 109(1):193–198, January 1, 1958. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1964:GGC

[FGMZ64] R. P. Feynman, M. Gell-Mann, and G. Zweig. Group U(6) ⊗ U(6) generated by current components. Physical Review Letters, 13(2): 678–680, November 30, 1964. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Feynman:1965:QMP

[FH65] Richard P. (Richard Phillips) Feynman and Albert R. Hibbs. Quantum mechanics and path integrals. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1965. xiv + 365 pp. LCCN QC174.1 .F39.

Feynman:1968:KMI

[FH68] Richard P. (Richard Phillips) Feynman and A. R. Hibbs. Kvantovaía mekhanika i integraly po traektoriíam. (Russian) [Quantum mechanics and path integrals]. Izdatelstvo Mir, Moscow, USSR, 1968. 382 pp. LCCN QC174.1 .F3917. Translation from English to Russian by E. M. Barlita and Yu. L. Obukhova. Edited by V. S. Barashenkova.

Feynman:1969:QVN

[FH69] R. P. Feynman and A. R. Hibbs. Quantenelektrodynamik. Eine Vorlesungsmitschrift und Nachdruck von Originalarbeiten. (German) [Quantum electrodynamics. Lecture notes and reproduction of original work], volume 401/401a of BI-Hochschultaschenbücher. Bibliographisches Institut, Mannheim, Germany and Zürich, Switzerland, 1969. 249 pp. German translation by Jochen Benecke and Veronika Wagner.

Feynman:1974:TWA

[FH74] Richard P. Feynman and Sir Fred Hoyle. Take the world from another point of view. Engineering and Science (Caltech), 37(4):11–13, February 1974. CODEN ???? ISSN 0013-

7812. URL http://calteches.library.caltech.edu/archive/
00000035/02/PointofView.pdf.

Feynman:1993:QMP

[FH93] Richard Phillips Feynman and Albert R. Hibbs. Quantum mechanics and path integrals. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-020650-3. xiv + 365 pp. LCCN ????

Feynman:1996:FLC

[FHA96] Richard P. (Richard Phillips) Feynman, Anthony J. G. Hey, and Robin W. Allen. Feynman lectures on computation. Addison-Wes-ley, Reading, MA, USA, 1996. ISBN 0-201-48991-0. xiv + 303

pp. LCCN QA76 .F45 1996. URL http://www.loc.gov/catdir/
enhancements/fy0831/96025127-d.html.

Feynman:1976:QEL

[FHH^{+}62] Richard Phillips Feynman, Albert R. Hibbs, Elisha R. Huggins, Harold T. Yura, and Peter Cziffra, editors. Quantum electrodynamics: a lecture note and reprint volume, volume 3 of Frontiers in physics. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 1962. ISBN 0-8053-2501-8 (paperback). ISSN 0429-7725. xi + 198 pp. LCCN QC680 1962.

Feynman:1962:MSE

[FHIP62] R. P. Feynman, R. W. Hellwarth, C. K. Iddings, and P. M. Platzman. Mobility of slow electrons in a polar crystal. Physical Review, 127(4):1004–1017, August 15, 1962. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:2003:FLG

[FHM03] Richard P. (Richard Phillips) Feynman, Brian Hatfield, and Fernando B. Morinigo. Feynman Lectures on Gravitation. Westview Press, Boulder, CO, 2003. ISBN 0-8133-4038-1. xl + 232 pp. LCCN ???? See correction [UMF16].

Feynman:1951:HEP

[FHRK51] Richard P. Feynman, Carl W. Helstrom, Malvin A. Ruderman, and William Karzas. High energy phenomena and meson theories. Notes on course at CIT, January–March 1951, 1951.

Feynman:2010:QMP

[FHS10] Richard P. (Richard Phillips) Feynman, Albert R. Hibbs, and Daniel F. Styer. Quantum mechanics and path integrals. Dover, New York, NY, USA, 2010. ISBN 0-486-47722-3. xii + 371 pp.

LCCN QC174.12 .F484 2010. URL http://www.loc.gov/catdir/
enhancements/fy1006/2010004550-d.html.

Field:2006:QMS

[Fie06] J. H. Field. Quantum mechanics in space-time: the Feynman path amplitude description of physical optics, de Broglie matter waves and quark and neutrino flavour oscillations. Annals of Physics, 321(3):627–707, 2006. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Fischer:2010:HQE

[Fis10] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2012:HQE

[Fis12] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Feynman:1986:ECP

[FK86] R. P. Feynman and H. Kleinert. Effective classical partition functions. Physical Review. A. Third Series, 34(6):5080–5084, December 1, 1986. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Feynman:1971:CME

[FKR71] R. P. Feynman, M. Kislinger, and F. Ravndal. Current matrix elements from a relativistic quark model. Physical Review D (Particles and Fields), 3(11):2706–2732, June 1, 1971. CODEN PRV-DAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500.

Feinman:1977:FLPc

[FL77] R. Feĭnman and R. Leĭton. Feinmanovskie lektsii po fizike. 6. Elektrodinamika. Izdatelstvo Mir, Moscow, USSR, second edition, 1977. 349 pp. Translated from the English by A. V. Efremov, G. I. Kopylov and Ju. A Simonov, Edited by Ja. A. Smorodinskiĭ.

Feynman:1984:DP

[FL84] Richard P. Feynman and Ralph Leighton. The dignified professor. Engineering and Science (Caltech), 48(2):4–10, November 1984. CODEN ???? ISSN 0013-7812. URL http://resolver.caltech.

edu/CaltechES:48.2.dignified.

Feynman:1987:QTT

[FL87a] Richard Feynman and Ralph Leighton. The quest for Tannu Tuva. BBC TV ’Horizon’, 1987. 50-minute film. Republished by PBS Nova as “Last Journey of a Genius”.

Feynman:1987:MFG

[FL87b] Richard P. Feynman and Ralph Leighton. Mr. Feynman goes to Washington. Engineering and Science (Caltech), 51(1):6–22, Fall 1987. CODEN ???? ISSN 0013-7812. URL http://calteches.

library.caltech.edu/48/2/1987Feynman.pdf.

Feynman:1988:WDY

[FL88] Richard Phillips Feynman and Ralph Leighton. What do YOU care what other people think?: Further adventures of a curious character. W. W. Norton & Co., New York, NY, USA, 1988. ISBN 0-393-02659-0. 255 pp. LCCN QC16.F49 A3 1988. US$17.95.

Feynman:2006:CFA

[FL06] Richard Phillips Feynman and Ralph Leighton. Classic Feynman: All the Adventures of a Curious Character. W. W. Norton & Co., New York, NY, USA, 2006. ISBN 0-393-06132-9. x + 511 pp. LCCN QC16.F49 A3 2006. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http://
www.loc.gov/catdir/toc/ecip0515/2005018928.html.

Feynman:2014:BWS

[FL14] Richard Phillips Feynman and Ralph Leighton. “Sie belieben wohl zu scherzen, Mr. Feynman!”: Abenteuer eines neugierigen Physikers. (German) [“Surely You’re Joking, Mr. Feynman”: Adventures of a curious physicist], volume 5155 of Serie Piper. Piper, München, Germany, eighth edition, 2014. ISBN 3-492-25155-2. 462 pp. LCCN ????

Feynman:1985:SYJc

[FLH85a] Richard P. Feynman, Ralph Leighton, and Edward Hutchings. “Surely You’re Joking, Mr. Feynman!”: Adventures of a Curious Character. Bantam Books, New York, NY, USA, 1985. ISBN 0-553-25649-1. xi + 322 pp. LCCN QC16.F49A37 1986.

Feynman:1985:SYJa

[FLH85b] Richard P. (Richard Phillips) Feynman, Ralph Leighton, and Edward Hutchings. “Surely You’re Joking, Mr. Feynman!”: Adventures of a Curious Character. W. W. Norton & Co., New York, NY, USA, 1985. ISBN 0-393-01921-7. 350 pp. LCCN QC16.F49 A37 1985.

Feynman:1964:FLP

[FLS64] Richard P. Feynman, Robert B. Leighton, and Matthew Sands. The Feynman lectures on physics. Vol. 2: Mainly electromagnetism and matter. Addison-Wesley, Reading, MA, USA, 1964. xii + 569 pp.

Feynman:1965:FLP

[FLS65a] Richard P. Feynman, Robert B. Leighton, and Matthew Sands. The Feynman lectures on physics. Vol. 3: Quantum mechanics. Addison-Wesley, Reading, MA, USA, 1965. x + 365 pp.

Feynman:1963:FLP

[FLS65b] Richard P. (Richard Phillips) Feynman, Robert B. Leighton, and Matthew L. (Matthew Linzee) Sands. The Feynman lectures on physics. Addison-Wesley, Reading, MA, USA, 1963–1965. xii + 513 pp. LCCN QC21.2 .F49 1963; QC23 .F47. Three volumes.

Feynman:1966:FVP

[FLS66] R. Feynman, R. Leighton, and M. Sands. Feynman Lectures on Physics. III. Quantum mechanics, 1, 2. Izdatelstvo Mir, Moscow, USSR, 1966. 272, 260 pp. Translated by G. I. Kopykov to Russian from the English original. Edited by Ja. Asmorodinskii.

Feynman:1969:MF

[FLS70] R. P. Feynman, R. B. Leighton, and M. Sands. Mai Fizika. Müszaki könvykiadó, Budapest, Hungary, 1969–1970. ???? pp. LCCN C23.F4351 1969. Hungarian translation of the Feynman Lectures on Physics.

Feinman:1976:FLPa

[FLS76a] R. Feĭnman, R. Leĭton, and M. Sènds. Feinmanovskie lektsii po fizike. Vyp. 1. Sovremennaya nauka o prirode. Zakony mekhaniki. Vyp. 2. Prostranstvo. Vremya. Dvizhenie. Izdatelstvo Mir, Moscow, USSR, third edition, 1976. 439 pp. Translated to Russian from the English by A. V. Efremov, G. I. Kopylov and O. A. Hrustalev, Edited by Ja. A. Smorodinskiĭ.

Feinman:1976:FLPb

[FLS76b] R. Feĭnman, R. Leĭton, and M. Sènds. Feinmanovskie lektsii po fizike. Vyp. 3. Izluchenie. Volny. Kvanty. Vyp. 4. Kinetika. Teplota. Zvuk. Izdatelstvo Mir, Moscow, USSR, third edition, 1976. 439 pp. Translated to Russian from the English by A. V. Efremov, G. I. Kopylov and O. A. Hrustalev, Edited by Ja. A. Smorodinskiĭ.

Feinman:1977:FLPb

[FLS77a] R. Feĭnman, R. Leĭton, and M. Sènds. Feinmanovskie lektsii po fizike. 5. Elektrichestvo i magnetizm. Izdatelstvo Mir, Moscow, USSR, second edition, 1977. 302 pp. Translated from the English by G. I. Kopylov and Ju. A. Simonov, Edited by Ja. A. Smorodinskiĭ.

Feinman:1977:FLPa

[FLS77b] R. Feĭnman, R. Leĭton, and M. Sènds. Feinmanovskie lektsii po fizike. 7. Fizika sploshnykh sred. Izdatelstvo Mir, Moscow, USSR, second edition, 1977. 288 pp. With a supplement “A dynamical model of crystal structure” by Lawrence Bragg and J. F. Nye in English and Russian, Translated from the English by A. V. Effemov [A. V. Efremov] and Ju. A. Simonov, Edited by Ja. A. Smorodinskiĭ.

Feinman:1978:FLP

[FLS78] R. Feĭnman, R. Leĭton, and M. Sènds. Feinmanovskie lektsii po fizike. 8, 9. Kvantovaya mekhanika. Izdatelstvo Mir, Moscow, USSR, second edition, 1978. 526 pp. Translated from the English by G. I. Kopylov, Edited by Ja. A. Smorodinskiĭ.

Feynman:1989:FLP

[FLS89] Richard P. (Richard Phillips) Feynman, Robert B. Leighton, and Matthew L. (Matthew Linzee) Sands. The Feynman lectures on physics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-51003-0 (vol. 1). ???? pp. LCCN QC21.2 .F49 1989. Three volumes.

Feynman:1995:SEP

[FLS95] Richard Phillips Feynman, Robert B. Leighton, and Matthew L. (Matthew Linzee) Sands. Six easy pieces: essentials of physics, explained by its most brilliant teacher. Helix books. Addison-Wes-ley, Reading, MA, USA, 1995. ISBN 0-201-40955-0, 0-201-40956-9 (set), 0-201-48308-4 (cassettes), 0-201-40825-2 (paperback). xxix + 145 pp. LCCN QC21.2 .F52 1995b; QC21.2 .F52 1995. US$22.00. Originally prepared for publication by Robert B. Leighton and Matthew L. (Matthew Linzee) Sands. New introduction by Paul Davies. See also [Fey97].

Feynman:2006:FLP

[FLS06] Richard P. (Richard Phillips) Feynman, Robert B. Leighton, and Matthew L. (Matthew Linzee) Sands. The Feynman lectures on physics. Pearson/Addison-Wesley, San Francisco, CA, USA, 2006. ISBN 0-8053-9046-4 (vol. 1), 0-8053-9047-2 (vol. 2), 0-8053-9049-9 (vol. 3). ???? pp. LCCN QC21.2 .F49 2006.

Feynman:2014:EFL

[FLS14] Richard P. (Richard Phillips) Feynman, Robert B. Leighton, and Matthew L. (Matthew Linzee) Sands, editors. Exercises for the Feynman lectures on physics. Basic Books, New York, NY, USA, new millennium edition, 2014. ISBN 0-465-06071-4. ???? pp. LCCN QC32 .F39 2014.

Feynman:1969:EIP

[FLV69] Richard P. Feynman, R. B. Leighton, and R. E. Vogt. Exercises in Introductory Physics. Addison-Wesley, Reading, MA, USA, 1969. ???? pp. LCCN ????

Feynman:1947:ESE

[FMT47] Richard P. Feynman, N. Metropolis, and E. Teller. Equations of state of elements based on the generalized Fermi–Thomas theory. Report OSTI 4417654, Los Alamos Scientific Laboratory, Atomic Energy Commission, Los Alamos, NM, USA, 1947.

Feynman:1949:ESEa

[FMT49a] R. P. Feynman, N. Metropolis, and E. Teller. Equations of state of elements based on the generalized Fermi–Thomas theory. Technical Report AECD-2448, Technical Information Branch, Oak Ridge Operations, AEC, Oak Ridge, TN, USA, January 20,

1949. 41 pp. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0107.pdf;   http://www.osti.gov/
servlets/purl/4417654-BCgOtj/native/.

Feynman:1949:ESEb

[FMT49b] R. P. Feynman, N. Metropolis, and E. Teller. Equations of state of elements based on the generalized Fermi–Thomas theory. Physical Review (2), 75(10):1561–1573, May 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1995:FLG

[FMWH95] Richard P. (Richard Phillips) Feynman, Fernando B. (Fernando Bernardino) Morinigo, William G. (William Gerard) Wagner, and Brian Hatfield. Feynman lectures on gravitation. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-62734-5. xl + 232

pp. LCCN QC178 .F49 1995. URL http://www.loc.gov/catdir/
enhancements/fy0831/95011076-b.html;   http://www.loc.
gov/catdir/enhancements/fy0831/95011076-d.html.

Forstner:2007:QIK

[For07] Christian Forstner. Quantenmechanik im Kalten Krieg: David Bohm und Richard Feynman. (German) [Quantum mechanics in the Cold War: David Bohm and Richard Feynman]. Verl. für Geschichte der Naturwiss. und der Technik, Diepholz, Germany, 2007. ISBN 3-928186-81-7 (paperback). 244 pp. LCCN ???? EUR 30.00EUR 30.00. URL http://paperc.de/4165-

quantenmechanik-im-kalten-krieg-9783928186810.

Forstner:2019:BRP

[For19] Christian Forstner. Book review: Paul Halpern. The Quantum Labyrinth: How Richard Feynman and John Wheeler Revolutionized Time and Reality. Isis, 110(1):197–198, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Forstner:2020:BAM

[For20] Christian Forstner. Biography and autobiography in the making of a genius: Richard P. Feynman. In Forstner and Walker [FW20], chapter 9, pages 145–159. ISBN 3-030-48508-0 (print), 3-030-48509-9 (e-book), 3-030-48510-2, 3-030-48511-0. LCCN QC16.E5

.B564 2020. URL http://www.springerlink.com/content/978-
3-030-48509-2.

Forgione:2022:FST

[For22] Marco Forgione. Feynman’s space–time view in quantum electrodynamics. Studies in History and Philosophy of Science Part A, 93(??):136–148, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368122000577.

Forgione:2024:FDV

[For24] Marco Forgione. Feynman diagrams: visualization of phenomena and diagrammatic representation. European Journal for Philosophy of Science, 14(4):??, December 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-024-00609-1.

Feynman:1970:SCB

[FPT70] R. P. Feynman, S. Pakvasa, and S. F. Tuan. Some comments on baryonic states. Physical Review D (Particles and Fields), 2(7): 1267–1270, October 1, 1970. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500.

Flynn:1998:BRM

[FR98] Michael Flynn and Linda Rothstein. Bulletin: The real mystery science theater. Bulletin of the Atomic Scientists, 54(6):9–10, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feynman:1999:PFT

[FR99] Richard P. (Richard Phillips) Feynman and Jeffrey Robbins. The Pleasure of Finding Things Out: the Best Short Works of Richard P. Feynman. Helix books. Perseus Publishers, Cambridge, MA, USA, 1999. ISBN 0-7382-0108-1. xvi + 270 pp. LCCN Q171

.F385 1999. URL http://www.loc.gov/catdir/enhancements/

fy0833/99064775-d.html. Foreword by Freeman Dyson.

Feynman:2005:PFT

[FR05] Richard P. (Richard Phillips) Feynman and Jeffrey W. Robbins, editors. The Pleasure of Finding Things Out: the Best Short Works of Richard P. Feynman. Basic Books, New York, NY, USA, 2005. ISBN 0-465-02395-9 (paperback). xvi + 270 pp. LCCN Q171 .F385 2005.

French:2006:BRD

[Fre06] Steven French. Book review: David Kaiser: Drawing Theories Apart: The Dispersion of Feynman Diagrams in Postwar Physics. Isis, 97(1):185–186, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/10.1086/504564.

Friedman:1991:IU

[Fri91] Herbert Friedman. The invisible universe. Proceedings of the American Philosophical Society, 135(3):339–345, September 1991. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/986772.

Fritzsch:2011:YWM

[Fri11] Harald Fritzsch. You are wrong, Mr. Einstein!: Newton, Einstein, Heisenberg, and Feynman discussing quantum mechanics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 981-4324-99-X. xxi + 178 pp. LCCN QC174.12 .F755 2011.

Fritzsche:2014:FTQ

[Fri14] S. Fritzsche. The Feynman tools for quantum information processing: Design and implementation. Computer Physics Communications, 185(6):1697–1718, June 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465514000356.

Feynman:2001:EVD

[FRL01] Richard P. Feynman, Jeffrey Robbins, and Inge Leipold. Es ist so einfach: vom Vergnügen, Dinge zu entdecken. (German) [It is so simple: the pleasure of finding things out].

Piper, München, Germany, 2001. ISBN 3-492-04251-1. 279 pp. LCCN ???? DM 44.00, EUR 22.50. URL http://

www.gbv.de/dms/faz-rez/FD1N20010705852023.pdf;   http:/
/www.gbv.de/dms/ilmenau/toc/327975911.PDF.

Feynman:2004:SAM

[FRRZ04] Richard P. (Richard Phillips) Feynman, Rolando Rebolledo, Jorge Rezende, and Jean-Claude Zambrini, editors. Stochastic analysis and mathematical physics (SAMP/ANESTOC 2002): proceedings of the Mathematical legacy of R. P. Feynman, Lisbon, Portugal, 3–7 June 2002: proceedings of the Open Systems and Quantum Statistical Mechanics, Santiago, Chile, 7–11 January 2002. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2004. ISBN 981-256-064-5. LCCN QC20.7.S8 S734 2002.

Feynman:1954:PNM

[FS54] R. P. Feynman and G. Speisman. Proton-neutron mass difference. Physical Review (2), 94(2):500, April 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Fleischer:1992:SPC

[FT92] J. Fleischer and O. V. Tarasov. SHELL2: a package for the calculation of two-loop on-shell Feynman diagrams in FORM. Computer Physics Communications, 71(1–2):193–205, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290083B.

Feynman:1963:TGQ

[FV63] R. P. Feynman and F. L. Vernon, Jr. The theory of a general quantum system interacting with a linear dissipative system. Annals of Physics, 24(1):118–173, October 1963. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Feynman:2000:TGQ

[FV00] R. P. Feynman and F. L. Vernon, Jr. The theory of a general quantum system interacting with a linear dissipative system. Annals of Physics, 281(1–2):546–607, April 5, 2000. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Feynman:1957:GRS

[FVH57] Richard P. Feynman, Frank L. Vernon, Jr., and Robert W. Hellwarth. Geometrical representation of the Schrödinger equation for solving maser problems. Journal of Applied Physics, 28(1):49–52, 1957. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850.

Feynman:1941:RAM

[FW41] R. P. Feynman and J. A. Wheeler. Reaction of the absorber as the mechanism of radiative damping. Physical Review (2), 59(8):683, April 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Feynman:1946:NDS

[FW46] Richard P. Feynman and T. A. Welton. Neutron diffusion in a space lattice of fissionable and absorbing materials. Report OSTI 4381097, Los Alamos Scientific Laboratory, Atomic Energy Commission, Los Alamos, NM, USA, 1946.

Feynman:1987:EPL

[FW87] Richard P. (Richard Phillips) Feynman and Steven Weinberg. Elementary particles and the laws of physics: the 1986 Dirac memorial lectures. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34000-4 (hardcover), 0-521-65862-4 (paperback). x + 110 pp.

LCCN QC793.28 .F49 1987. URL http://www.loc.gov/catdir/
description/cam023/87026362.html;   http://www.loc.gov/
catdir/toc/cam023/87026362.html.

Forstner:2020:BHP

[FW20] Christian Forstner and Mark Walker, editors. Biographies in the History of Physics: Actors, Objects, Institutions. Springer International Publishing, Cham, Switzerland, 2020. ISBN 3-030-48508-0 (print), 3-030-48509-9 (e-book), 3-030-48510-2, 3-030-48511-0. vi + 324 + 38 pp. LCCN QC16.E5 .B564 2020. URL http://www.

springerlink.com/content/978-3-030-48509-2.

Galison:1998:FWM

[Gal98] Peter Galison. Feynman’s war: Modelling weapons, modelling nature. Studies in History and Philosophy of Modern Physics, 29(3): 391–434, September 1998. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219898000136.

Goodstein:1997:FLL

[GFG97] David L. Goodstein, Richard P. (Richard Phillips) Feynman, and Judith R. Goodstein. Feynman’s lost lecture: the motion of planets around the sun. Vintage Books, New York, NY, USA, 1997. ISBN 0-09-973621-7. 191 pp. LCCN QB353 .G66 1997.

Gribbin:1997:RFL

[GG97] John R. Gribbin and Mary Gribbin. Richard Feynman: a life in science. Dutton, New York, NY, USA, 1997. ISBN 0-525-94124-X. xvii + 301 pp. LCCN QC16.F49 G75 1997.

Goodstein:2000:RFH

[GG00] David Goodstein and Judith Goodstein. Richard Feynman and the history of superconductivity. Physics in Perspective (PIP), 2(1):30– 47, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s000160050035.

Ginzburg:2001:PLR

[Gin01] V. L. (Vitaliĭ Lazarevich) Ginzburg. The Physics of a Lifetime: Reflections on the Problems and Personalities of 20th Century Physics. Physics and astronomy online library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-67534-5, 3-642-08699-3. xiii + 513 pp. LCCN QC7 .G59 2001. URL http://catdir.loc.

gov/catdir/enhancements/fy0812/00066112-d.html;   http:
//catdir.loc.gov/catdir/enhancements/fy0812/00066112-t.
html.

Gleick:1988:RFD

[Gle88] James Gleick. Richard Feynman dead at 69; leading theoretical physicist. New York Times, ??(??):A1, D27, February 17, 1988. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/110578382/.

Gleick:1992:GLS

[Gle92] James Gleick. Genius: The Life and Science of Richard Feynman. Pantheon Books, New York, NY, USA, 1992. ISBN 0-679-40836-3. x + 531 pp. LCCN QC16.F49G54 1992. US$27.50.

Gell-Mann:1989:DFG

[GM89] Murray Gell-Mann. Dick Feynman — the guy in the office down the hall. Physics Today, 42(2):50–54, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Gaspar:1987:GHF

[GN87] R. Gáspár and Á. Nagy. Generalized Hellmann–Feynman theorem in the Xα method. International Journal of Quantum Chemistry, 31(4):639–647, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goodstein:1989:RPF

[Goo89] David L. Goodstein. Richard P. Feynman, teacher. Physics Today, 42(2):70–75, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goodstein:1999:CHB

[Goo99] Judith Goodstein. A conversation with Hans Bethe. Physics in Perspective (PIP), 1(3):253–281, October 1999. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/s000160050022;
 http://www.springerlink.com/content/k4f0pllveakla47q/.

Goodstein:2001:FRC

[Goo01] David Goodstein. Feynman returns to centre stage. Physics World, 14(5):45, May 2001. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.iop.

org/2058-7058/14/i=5/a=33.

Grozin:2007:LQQ

[Gro07] Andrey G. Grozin. Lectures on QED and QCD: practical calculations and renormalization of one- and multi-loop Feynman diagrams. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2007. ISBN 981-256-914-6. xi + 224 pp. LCCN QC680 .G76 2007.

Gross:2012:PPR

[Gro12] Ari Gross. Pictures and pedagogy: the role of diagrams in Feynman’s early lectures. Studies in History and Philosophy of Modern Physics, 43(3):184–194, August 2012. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219812000226.

Gross:2014:BRB

[Gro14] Ari Gross. Book review: Feynman, by J. Ottaviani and L. Myrick. Annals of Science, 71(2):292–294, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hutchings:1958:FSS

[H^{+}58] Edward Hutchings et al., editors. Frontiers in science. a survey. Basic Books, New York, NY, USA, 1958. 362 pp.

LCCN Q162 .H83. URL http://catalog.hathitrust.org/api/
volumes/oclc/1158177.html.

Hey:1996:FLC

[HA96] Anthony J. G. Hey and Robin W. Allen, editors. Feynman lectures on computation. The advanced book program. Perseus Publishers, Cambridge, MA, USA, 1996. ISBN 0-7382-0296-7. xiv + 303 pp. LCCN QA76 .F45 1999.

Hafner:1965:BRR

[Haf65] E. M. Hafner. Book review: Richard P. Feynman and Robert B. Leighton and Matthew Sands, The Feynman Lectures on Physics, Vol. I. American Journal of Physics, 33(9):750–752, September 1965. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v33/
i9/p750_s1.

Hall:2009:GSP

[Hal09] Derek Hall, editor. Great scientists. Physical sciences. Facts at your fingertips. Brown Bear Books, Redding, CT, USA, 2009. ISBN 1-933834-46-3. 64 pp. LCCN QC15 .P468 2009.

Halpern:2017:QLH

[Hal17] Paul Halpern. The Quantum Labyrinth: How Richard Feynman and John Wheeler Revolutionized Time and Reality. Basic Books, New York, NY, USA, 2017. ISBN 0-465-09758-8 (hardcover), 0-465-09759-6 (e-book). ix + 311 pp. LCCN QC174.12 .H347 2017.

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Heisenberg:1962:DRH

[Hei62] Werner Heisenberg. Discussion des rapports de Heitler et Feynman. (French) [Discussion of the reports of Heitler and Feynman]. In R. Stoops, editor, La Théorie Quantique des Champs. Rapports et Discussions du Douzième Conseil de Physique à Bruxelles, 9-14 Octobre, 1961, pages 92–97, 169–170, 173–176, 201–203, 243–247. Interscience Publishers, New York, NY, USA, 1962. LCCN ????

Hellmann:1937:EQG

[Hel37] Hans Hellmann. Einführung in die Quantenchemie. (German) [Introduction to Quantum Chemistry]. Franz Deuticke, Leipzig, Germany and Vienna, Austria, 1937. ix + 350 pp. URL http://www.

db-thueringen.de/servlets/DocumentServlet?id=11268. Pub-

lication number 3754. See note in [Fey39a].

Henderson:2010:RFQ

[Hen10] Harry Henderson. Richard Feynman: quarks, bombs, and bongos. Makers of modern science. Chelsea House, New York, NY, USA, 2010. ISBN 0-8160-6176-9. ???? pp. LCCN QC16.F49 H46 2010.

Herzberg:2013:EMP

[Her13] Frederik S. Herzberg. Excursion to mathematical physics: a radically elementary definition of Feynman path integrals. Lecture Notes in Mathematics, 2067:71–75, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33149-7_8/.

Hey:1996:MRF

[Hey96] Anthony J. G. Hey. Memories of Richard Feynman. Physics Today, 49(9):44–49, September 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v49/i9/p44_s1.

Hey:1999:FCE

[Hey99a] Anthony J. G. Hey, editor. Feynman and Computation: Exploring the Limits of Computers. Perseus Publishers, Cambridge, MA, USA, 1999. ISBN 0-7382-0057-3. xxiii + 438 pp. LCCN QA76 .F46 1999, QC52 .F49 199. US$50.00.

Hey:1999:RFC

[Hey99b] Tony Hey. Richard Feynman and computation. Contemporary Physics, 40(4):257–265, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hey:2002:FCE

[HF02] Anthony J. G. Hey and Richard P. (Richard Phillips) Feynman, editors. Feynman and computation: exploring the limits of computers. Westview Press/Perseus Books, Cambridge, MA, USA, 2002. ISBN 0-8133-4039-X. xxiii + 438 pp. LCCN QC52 .F49 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0830/2007310126-b.
html;   http://www.loc.gov/catdir/enhancements/fy0830/
2007310126-d.html;   http://www.loc.gov/catdir/toc/fy0804/
2007310126.html.

Harmon:2007:SLG

[HG07] Joseph E. Harmon and Alan G. Gross, editors. The scientific literature: a guided tour. University of Chicago Press, Chicago, IL, USA and London, UK, 2007. ISBN 0-226-31655-6 (hardcover), 0-226-31656-4 (paperback). xxiv + 327 pp.

LCCN Q225.5 .S35 2007. URL http://catdir.loc.gov/catdir/
enhancements/fy0701/2006016547-t.html;   http://catdir.
loc.gov/catdir/enhancements/fy0707/2006016547-b.html;
http://catdir.loc.gov/catdir/enhancements/fy0707/2006016547-
d.html.

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Hillis:1989:RFC

[Hil89] W. Daniel Hillis. Richard Feynman and the Connection Machine. Physics Today, 42(2):78–83, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

kurzweilai.net/articles/art0504.html?printable=1.

Hahn:2008:FTD

[HL08] T. Hahn and P. Lang. FeynEdit — a tool for drawing Feynman diagrams. Computer Physics Communications, 179(12):931–935, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002749.

Harlander:2024:DCM

[HM24] Robert V. Harlander and Jean-Philippe Martinez. The development of computational methods for Feynman diagrams. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL

https://link.springer.com/article/10.1140/epjh/s13129-
024-00067-6.

Horvitz:2000:QSW

[Hor00] Leslie Alan Horvitz. The quotable scientist: words of wisdom from Charles Darwin, Albert Einstein, Richard Feynman, Galileo, Marie Curie, and more. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-136063-8. x + 169 pp. LCCN Q173 .H739 2000. URL http:

//www.loc.gov/catdir/description/mh021/00709492.html;
http://www.loc.gov/catdir/enhancements/fy0707/00709492-
b.html;   http://www.loc.gov/catdir/toc/mh021/00709492.
html.

Hughes:1986:BRR

[Hug86] Will Hughes. Book review: Richard Feynman, QED. American Journal of Physics, 54(8):766–767, August 1986. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.

aapt.org/resource/1/ajpias/v54/i8/p766_s2.

Hurst:1952:EGF

[Hur52] C. A. Hurst. The enumeration of graphs in the Feynman–Dyson technique. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 214(1116):44–61, August 7, 1952. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/99158.

Hutchisson:1968:BFR

[Hut68] Elmer Hutchisson. Book and film reviews: Richard P. Feynman, The Character of Physical Law. The Physics Teacher, 6(2):89– 90, February 1968. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Hey:1986:QU

[HW86] Anthony J. G. Hey and Patrick Walters. The Quantum Universe. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-26744-7 (hardcover), 0-521-31845-9 (paperback). vii + 180 pp. LCCN QC174.12 .H48 1987.

Hsu:2001:LPI

[HZ01] J. P. (Jong-Ping) Hsu and Yuanzhong Zhang, editors. Lorentz and Poincaré invariance: 100 years of Relativity, volume 8 of Advanced series on theoretical physical science. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2001. ISBN 981-02-4721-4. xxxi + 583 pp. LCCN

QC174.17.S9 H78 2001. URL http://www.loc.gov/catdir/toc/
fy031/2002280428.html.

Ivancevic:2008:QLD

[II08] Vladimir G. Ivancevic and Tijana T. Ivancevic. Quantum leap: from Dirac and Feynman, across the universe, to human body and mind. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2008. ISBN 981-281-927-4, 981-281-928-2 (e-book). xvi + 839 pp. LCCN QC174.12 .I927 2008.

Ishak:2019:BRFb

[Ish19] B. Ishak. Book review: Feynman and his physics: the life and science of an extraordinary man, by Jörg Resag, Cham, Springer, 2018, 319 pp., £24.99 (hardback), ISBN: 978-3-319-96835-3. Contemporary Physics, 60(4):340–341, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jacobsen:2010:RBI

[Jac10] Anja Skaar Jacobsen. Reviews: Quantenmechanik im Kalten Krieg: David Bohm und Richard Feynman — by Christian Forstner. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):171–172, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jacobsen:2020:BRJ

[Jac20] Anja Skaar Jacobsen. Book review: Jörg Resag. Feynman and His Physics: The Life and Science of an Extraordinary Man. (Springer Biographies). xii + 319 pp., figs., bibl. Cham, Switzerland: Springer, 2018. EUR 32.69 (cloth); ISBN 978-3-319-96835-3. E-book available. Isis, 111(4):899–900, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jefferies:2004:FOC

[Jef04] Brian Jefferies. 7. Feynman’s operational calculus. Lecture Notes in Mathematics, 1843:157–171, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-70746-2_7/.

Jishi:2013:FDT

[Jis13] Radi A. Jishi. Feynman diagram techniques in condensed matter physics. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-02517-6 (hardcover). xiv + 400 pp. LCCN QC794.6.F4 J57

2013. URL http://assets.cambridge.org/97811070/25172/
cover/9781107025172.jpg.

Jishi:2014:FDT

[Jis14] Radi A. Jishi. Feynman diagram techniques in condensed matter physics. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-02517-6 (hardcover), 1-107-65533-1 (paperback). xiv + 400 pp. LCCN QC794.6.F4 J57 2013. URL http://catalogue.bnf.

fr/ark:/12148/cb43672681b.

Johnson:1986:GDS

[JL86] Gerald W. Johnson and Michel L. (Michel Laurent) Lapidus. Generalized Dyson series, generalized Feynman diagrams, the Feynman integral, and Feynman’s operational calculus, volume 351 of Memoirs of the American Mathematical Society. American Mathematical Society, Providence, RI, USA, 1986. ISBN 0-8218-2413-9 (paperback). ISSN 0065-9266. vi + 78 pp. LCCN QA3 .A57 no. 351 QC174.12.

Janke:2008:PIN

[JP08] W. (Wolfhard) Janke and Axel Pelster, editors. Path integrals: new trends and perspectives: proceedings of the 9th International Conference: Dresden, Germany, September 23–28, 2007. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2008. ISBN 981-283-726-4, 981-283-727-2 (e-book). LCCN QC174.17.P27 I57 2007eb.

Junk:2006:IVT

[JR06] Andreas Junk and Falk Riess. From an idea to a vision: There’s plenty of room at the bottom. American Journal of Physics, 74(9):825–830, September 2006. CODEN AJPIAS.

ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

ajp.aapt.org/resource/1/ajpias/v74/i9/p825_s1;   http:/
/scitation.aip.org/journals/doc/AJPIAS-ft/vol_74/iss_9/
825_1.html.

Kaiser:2001:BRBa

[Kai01] David Kaiser. Book review: Richard Feynman: A Life in Science by John Gribbin; Mary Gribbin. Isis, 92(1):207–208, March 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237421.

Kaiser:2002:EOF

[Kai02] David Kaiser. Einstein, Oppenheimer, Feynman: Physics in the 20th Century. Technical Report MIT OpenCourseWare STS.042J / 8.225J, MIT, Cambridge, MA, USA, 2002. Free, independent study course that explores the changing roles of physics and physicists during the 20th Century.

Kaiser:2005:CBL

[Kai05a] David Kaiser. Climbing Bethe’s ladder: Feynman diagrams and the many-body problem. In DTA [Kai05b], pages 230–243. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). LCCN QC794.6.F4 K35 2005.

URL http://www.loc.gov/catdir/bios/uchi051/2004023335.
html;   http://www.loc.gov/catdir/enhancements/fy0617/
2004023335-d.html;   http://www.loc.gov/catdir/toc/ecip051/
2004023335.html.

Kaiser:2005:DTA

[Kai05b] David Kaiser. Drawing Theories Apart: the Dispersion of Feynman Diagrams in Postwar Physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2005. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). xix + 469 pp. LCCN QC794.6.F4 K35 2005. URL http://www.loc.gov/

catdir/bios/uchi051/2004023335.html;   http://www.loc.
gov/catdir/enhancements/fy0617/2004023335-d.html;   http:
//www.loc.gov/catdir/toc/ecip051/2004023335.html.

Kaiser:2005:FDS

[Kai05c] David Kaiser. The Feynman–Dyson split. In DTA [Kai05b], pages 175–194. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). LCCN QC794.6.F4 K35 2005. URL http://www.loc.

gov/catdir/bios/uchi051/2004023335.html;   http://www.
loc.gov/catdir/enhancements/fy0617/2004023335-d.html;
http://www.loc.gov/catdir/toc/ecip051/2004023335.html.

Kaneko:1995:FGG

[Kan95] Toshiaki Kaneko. A Feynman-graph generator for any order of coupling constants. Computer Physics Communications, 92(2–3):127– 152, December 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595001226.

Kaneko:2018:CNF

[Kan18] T. Kaneko. Counting the number of Feynman graphs in QCD. Computer Physics Communications, 226(??):104–113, May 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
S0010465518300018.

Kaiser:2004:STT

[KIH04] David Kaiser, Kenji Ito, and Karl Hall. Spreading the tools of theory: Feynman diagrams in the USA, Japan, and the Soviet Union. Social Studies of Science, 34(6):879–922, December 1, 2004. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL http://www.jstor.org/

stable/4144350;   https://journals.sagepub.com/doi/pdf/
10.1177/0306312704046628.

Klauder:1972:MMJ

[Kla72] John R. Klauder, editor. Magic without magic: John Archibald Wheeler; a collection of essays in honor of his sixtieth birthday. W. H. Freeman and Company, New York, NY, USA, 1972. ISBN 0-7167-0337-8. xii + 491 pp. LCCN QC173 .M313.

Katzir:2013:TTH

[KLR13] Shaul Katzir, Christoph Lehner, and Jürgen Renn, editors. Traditions and transformations in the history of quantum physics: HQ-3, Third International Conference on the History of Quantum Physics, Berlin, June 28–July 2, 2010, volume 5 of Max Planck research library for the history and development of knowledge. Proceedings. Edition Open Access, Berlin, Germany, 2013. ISBN 3-8442-5134-0. LCCN QC173.98. URL http://www.edition-open-

access.de/proceedings/5/.

Kim:2019:NPE

[KN19] Y. S. Kim and Marilyn E. Noz. New perspectives on Einstein’s E = mc^{2}. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2019. ISBN 981-323-770-8 (hardcover), 981-323-771-6 (e-book). xi + 192 pp. LCCN QC173.59.S65 K56 2018.

Kragh:2008:BRD

[Kra08] Helge Kragh. Book review: David Kaiser, Drawing Theories Apart: The Dispersion of Feynman Diagrams in Postwar Physics. Chicago and London: The University of Chicago Press, 2005. Pp. xx + 469. ISBN 0-226-42267-4. £21.00, $30.00 (paperback). British Journal for the History of Science, 41(1):151–152, March 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Krauss:2011:OPA

[Kra11a] Lawrence M. Krauss. From one physicist to another: Lawrence Krauss reflects on the life and work of Richard Feynman. Scientific American, ??(??):??, March 18, 2011. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL https://www.

scientificamerican.com/article/lawrence-krauss-reflects-
life-work-richard-feynman/.

Krauss:2011:QMR

[Kra11b] Lawrence Maxwell Krauss. Quantum man: Richard Feynman’s life in science. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2011. ISBN 0-393-06471-9. xvii + 350 pp. LCCN QC16.F49

K73 2011. URL http://www.scientificcomputing.com/news-
DS-Illuminating-the-Life-and-Legacy-of-Richard-Feynman-
032211.aspx.

Kreimer:2000:KFD

[Kre00] Dirk Kreimer. Knots and Feynman diagrams, volume 13 of Cambridge lecture notes in physics. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-58761-1 (paperback). xii + 259 pp. LCCN QC174.52.K56 K74 2000. URL http://catdir.loc.

gov/catdir/description/cam0210/99035144.html;   http://
catdir.loc.gov/catdir/toc/cam022/99035144.html.

Kenner:1992:BCRc

[KYT^{+}92] Hugh Kenner, Tom Yager, Tom Thompson, Jon Udell, and Stanford Diehl. Book and CD-ROM reviews: Photographic lies: The search for visual truth, design tips, TCP/IP explainer, the times of physicist Richard Feynman, and more. Byte Magazine, 17(14):234– ??, December 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Laina:1998:XFD

[Lai98] A. Laina. Xfey, a Feynman diagram editor. Computer Physics Communications, 111(1–3):217–242, June 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465598000277.

Lanouette:2011:BRQ

[Lan11] William Lanouette. Book review: Quantum Man: Richard Feynman’s Life in Science, by Lawrence M. Krauss. New York and London: W. W. Norton & Company, 2011, 350 pp. Issues in Science and Technology, 27(4):??, Summer 2011. ISSN 0748-5492 (print), 1938-1557 (electronic). URL http://issues.org/27-4/

br_lanouette-4/.

Larkoski:2019:EPP

[Lar19] Andrew J. Larkoski. Elementary Particle Physics: an Intuitive Introduction. Cambridge University Press, Cambridge, UK, 2019. ISBN 1-108-49698-9 (hardcover). 540 (est.) pp. LCCN QC793.2 .L37 2019.

Lebowitz:1973:BRR

[Leb73] Joel L. Lebowitz. Book review: Richard P. Feynman, Statistical Mechanics: A Set of Lectures. Physics Today, 26(10):51–53, October 1973. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://ptonline.aip.org/dbt/dbt.jsp?KEY=
PHTOAD&Volume=26&Issue=10.

Leeds:1994:PWF

[Lee94] Stephen Leeds. Price on the Wheeler–Feynman theory. British Journal for the Philosophy of Science, 45(1):288–294, March 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537

(electronic). URL http://bjps.oxfordjournals.org/content/
45/1/288.full.pdf+html;   http://www.jstor.org/stable/
687973.

Leeds:1995:WFA

[Lee95] Stephen Leeds. Wheeler–Feynman again: A reply to Price. British Journal for the Philosophy of Science, 46(3):381–383, September 1995. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.org/
content/46/3/381.full.pdf+html;   http://www.jstor.org/
stable/687664.

Leighton:2000:TBR

[Lei00] Ralph Leighton. Tuva Or Bust!: Richard Feynman’s Last Journey. W. W. Norton & Co., New York, NY, USA, 2000. ISBN 0-393-32069-3 (paperback), 0-393-02953-0. 260 pp. LCCN QC16.F49 L45 2000.

Lepage:2007:BRD

[Lep07] G. Peter Lepage. Book review: David Kaiser, Drawing Theories Apart: The Dispersion of Feynman Diagrams in Postwar Physics. Chicago: University of Chicago Press, 2005, xix + 469 pages. $30.00 (paper). Physics in Perspective (PIP), 9(1):120–122, January 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Levine:1990:LGR

[Lev90] Michael J. S. Levine. A LaTEX graphics routine for drawing Feynman diagrams. Computer Physics Communications, 58(1– 2):181–198, February/March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090144P.

LeVine:2010:GES

[LeV10] Harry LeVine III. The great explainer: the story of Richard Feynman. Profiles in science. Morgan Reynolds Publishers, Greensboro, NC, USA, 2010. ISBN 1-59935-113-7. 144 pp. LCCN QC16.F49 L48 2010.

Lopes:1952:PMT

[LF52] José Leite Lopes and Richard P. Feynman. On the pseudoscalar meson theory of the deuteron. Notas de Física 2, Centro Brasileiro de Pesquisas Físicas, Av. Pasteur, 250, Rio de Janeiro, Brazil, 1952.

Lopes:1954:PMT

[LF54] José Leite Lopes and Richard P. Feynman. On the pseudoscalar meson theory of the deuteron. In New Research Techniques in Physics, 15–29 July 1951, page ?? Brazilian Academy of Sciences, Rio de Janeiro, Brazil, 1954.

Lightman:2019:DFW

[Lig19] Alan Lightman. The day Feynman worked out black-hole radiation on my blackboard. Web site., 2019. URL https://getpocket.

com/explore/item/the-day-feynman-worked-out-black-hole-

radiation-on-my-blackboard. The author reports that Richard Feynman worked out how a black hole could emit radiation, a year before Stephen Hawking made that discovery. The blackboard with Feynman’s calculations was wiped clean by a custodian, and lost.

Lindsay:1966:BRR

[Lin66] R. Bruce Lindsay. Book review: Richard P. Feynman and Robert B. Leighton and Matthew Sands, The Feynman Lectures on Physics, Vol. 3: Quantum Mechanics. Physics Today, 19(11):80–83, November 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Little:1989:AAP

[Lit89] Robert N. Little. American Association of Physics Teachers 1971 Oersted Medalist: Richard P. Feynman. American Journal of Physics, 57(6):492, June 1989. CODEN AJPIAS. ISSN 1943-2909.

Lubkin:1989:SIR

[Lub89] Gloria B. Lubkin. Special issue: Richard Feynman. Physics Today, 42(2):22–23, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Marateck:2006:DGP

[Mar06] Samuel L. Marateck. The differential geometry and physical basis for the applications of Feynman diagrams. Notices of the American Mathematical Society, 53(7):744–752, 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.

ams.org/notices/200607/fea-marateck.pdf.

Marcolli:2010:FM

[Mar10] Matilde Marcolli. Feynman motives. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2010. ISBN 981-4271-20-9 (hardcover), 981-4304-48-4 (paperback), 981-4271-21-7 (e-book). xiii + 220 pp. LCCN

QC174.17.F45 M37 2010eb. URL http://ebooks.worldscinet.
com/ISBN/9789814271219/9789814271219.html.

Mattuck:1967:GFD

[Mat67] Richard D. Mattuck. A guide to Feynman diagrams in the manybody problem. European physics series. McGraw-Hill, New York, NY, USA, 1967. xii + 294 pp. LCCN QC174.5 .M37.

Mattuck:1976:GFD

[Mat76] Richard D. Mattuck. A guide to Feynman diagrams in the manybody problem. McGraw-Hill, New York, NY, USA, second edition, 1976. ISBN 0-07-040954-4. xv + 429 pp. LCCN QC174.17.P7 M37 1976.

Mattuck:1992:GFD

[Mat92] Richard D. Mattuck. A guide to Feynman diagrams in the many-body problem. Dover books on physics and chemistry. Dover, New York, NY, USA, second edition, 1992. ISBN 0-486-67047-3 (paperback). xv + 429 pp. LCCN QC174.17.P7

M37 1992. US$11.95. URL http://www.loc.gov/catdir/
description/dover032/92004895.html.

Mazzucchi:2009:MFP

[Maz09] Sonia Mazzucchi. Mathematical Feynman path integrals and their applications. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2009. ISBN 981-283-690-X. viii + 216 pp. LCCN QC174.17.F45 M39 2009. URL http://www.

worldscibooks.com/promotion/feynman.html#7104.

Montwill:2011:QAD

[MB11a] Alex Montwill and Ann Breslin. The quantum adventure: does God play dice? World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (e-book). ix + 248 pp. LCCN QC174.12 .M66 2012.

Montwill:2011:RFS

[MB11b] Alex Montwill and Ann Breslin. Richard Feynman — the strange theory of light and matter. In The quantum adventure: does God play dice? [MB11a], chapter 14, pages 191–212. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (e-book). LCCN QC174.12 .M66 2012.

Milburn:2002:FPQ

[MD02] Gerard J. Milburn and Paul Davies. The Feynman processor: quantum entanglement and the computing revolution. Frontiers of science; Helix books. Perseus Publishers, Cambridge, MA, USA, 2002. ISBN 0-7382-0173-1. xiv + 213 pp. LCCN ????

Mehra:1994:BDD

[Meh94] Jagdish Mehra. The beat of a different drum: the life and science of Richard Feynman. Clarendon Press, Oxford, UK, 1994. ISBN 0-19-853948-7. xxxii + 630 + 16 pp. LCCN QC16.F49 M45 1994.

Mehra:2002:RPF

[Meh02] Jagdish Mehra. Richard Phillips Feynman 11 May 1918–15 February 1988. Biographical Memoirs of Fellows of the Royal Society, 48:97–128, December 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://www.jstor.

org/stable/3650252.

Mermin:2004:CFS

[Mer04] N. David Mermin. Could Feynman have said this? Physics Today, 57(5):10–11, May 2004. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Mermin:2016:WQR

[Mer16] N. David Mermin. Why Quark Rhymes with Pork, and Other Scientific Diversions. Cambridge University Press, Cambridge, UK, 2016. ISBN 1-107-02430-7 (hardcover), 1-139-16257-8 (e-book), 1-316-47617-0 (e-book). xiii + 391 pp. LCCN QC71 .M373 2016.

Mermin:2017:FBT

[Mer17] N. David Mermin. Feynman and Bell’s Theorem, and the pronunciation of “quark”. American Journal of Physics, 85(1):5, January 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Meynell:2008:WFD

[Mey08] Letitia Meynell. Why Feynman diagrams represent. International Studies in the Philosophy of Science, 22(1):39–59, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/02698590802280902.

Meynell:2018:PFD

[Mey18] Letitia Meynell. Picturing Feynman diagrams and the epistemology of understanding. Perspectives on Science, 26(4):459–481, July– August 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://www.mitpressjournals.org/doi/abs/
10.1162/posc_a_00283.

Mercereau:1956:PCF

[MF56] J. E. Mercereau and R. P. Feynman. Physical conditions for ferromagnetic resonance. Physical Review (2), 104(1):63, October 1, 1956. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Michalos:1967:BRBa

[Mic67] Alex C. Michalos. Book review: The Character of Physical Law by Richard Feynman. Philosophy of Science, 34(2):194, June 1967. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186108.

Miller:2012:YWM

[Mil12] D. J. Miller. You are Wrong Mr Einstein! Newton, Einstein, Heisenberg and Feynman Discussing Quantum Mechanics, by Harald Fritzsch, Scope: general interest. Level: undergraduates and/or general readers. Contemporary Physics, 53(5):437–439, 2012. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Metropolis:1987:NDP

[MKR87] N. (Nicholas) Metropolis, Donald M. Kerr, and Gian-Carlo Rota, editors. New Directions in Physics: The Los Alamos 40th Anniversary Volume. Academic Press, New York, NY, USA, 1987. ISBN 0-12-492155-8. xii + 292 pp. LCCN QC44 .N49

1987. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=6120718&query_id=0.

Mlodinow:2003:FRS

[Mlo03a] Leonard Mlodinow. Feynman’s rainbow: a search for beauty in physics and in life. Warner Books, New York, NY, USA, 2003. ISBN 0-446-53045-X, 0-446-69251-4 (paperback). xiv + 171 pp. LCCN QC16.M635 A3 2003.

Mlodinow:2003:STF

[Mlo03b] Leonard Mlodinow. Some time with Feynman. Allen Lane, London, UK, 2003. ISBN 0-7139-9643-9. xiv + 171 + 2 pp. LCCN QC16.M635 A3 2003b.

Morrison:1995:NTW

[Mor95] Philip Morrison. Nothing is too wonderful to be true, volume 11 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN 1-56396-363-9 (hardcover). xi + 446 pp. LCCN Q173.

Moral:2004:FKF

[Mor04] Pierre Moral. Feynman–Kac Formulae: Genealogical and Interacting Particle Systems with Applications. Probability and its Applications, A Series of the Applied Probability Trust. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 1-4419-1902-3. ISSN 1431-7028. LCCN QA273.A1-274.9; QA274-274.9.

Morton-Smith:2015:O

[MS15] Tom Morton-Smith. Oppenheimer. Oberon Books Ltd., London, UK, 2015. ISBN 1-78319-198-8. ???? pp. LCCN

???? URL http://stratford-upon-avon-theatre.blogspot.
com/2015/01/rsc-oppenheimer-review-stratford-upon.html.

Mukunda:2011:RPF

[Muk11] N. Mukunda. Richard Phillips Feynman: Physicist and teacher extraordinary. Resonance, 16(9):796–797, September 2011. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic).

Mullin:1974:RRP

[Mul74] William J. Mullin. Book review: R. P. Feynman, Statistical Mechanics, A Set of Lectures. American Journal of Physics, 42(7): 620–622, July 1974. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v42/i7/p620_s1.

Miranker:1966:FOC

[MW66] W. L. Miranker and B. Weiss. The Feynman operator calculus. SIAM Review, 8(2):224–232, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nelson:1989:MFT

[NC89] Philip Nelson and Sidney Coleman. Those magnetic Feynman tapes. Physics Today, 42(2):13–15, February 1989. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Nobel:1972:NLI

[Nob72] Nobel Foundation, editor. Nobel lectures, including presentation speeches and Laureates’ biographies, 1963–1970. Elsevier, Amsterdam, The Netherlands, 1972. ISBN 0-444-40993-9. xi + 349 pp. LCCN QC71.N735.

Noer:2011:BRJ

[Noe11] Richard Noer. Book review: Jennifer Coopersmith, Energy, the Subtle Concept: The Discovery of Feynman’s Blocks from Leibniz to Einstein. Oxford: Oxford University Press, 2010, xiv + 400 pages. $55.00 (cloth). Physics in Perspective (PIP), 13(3):379–380, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Nogueira:1993:AFG

[Nog93] P. Nogueira. Automatic Feynman graph generation. Journal of Computational Physics, 105(2):279–289, April 1993. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999183710740.

Nogueira:2017:FRCa

[Nog17a] P. Nogueira. From Feynman rules to conserved quantum numbers, I. Computer Physics Communications, 214(??):83–90, May 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465517300371.

Nogueira:2017:FRCb

[Nog17b] P. Nogueira. From Feynman rules to conserved quantum numbers, II. Computer Physics Communications, 215(??):13–19, June 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465517300395.

Ohl:1995:DFD

[Ohl95]  Thorsten Ohl. Drawing Feynman diagrams with LaTEX and  -

. Computer Physics Communications, 90(2–3):340–354, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Ottaviani:2013:F

[OM13] Jim Ottaviani and Leland Myrick. Feynman. First Second, New York, NY, USA, 2013. ISBN 1-59643-827-4 (paperback). 266 pp. LCCN ????

Onion:2015:GHB

[Oni15] Rebecca Onion. Graphic histories: Book reviews: Eugene Byrne and Simon Gurr. Darwin: A Graphic Biography. Washington, DC:

Smithsonian Books, 2013. 96 pp., illus. ISBN 978-1-58834-352-9. $9.95 (paper). Apostolos Doxiadis and Christos H. Papadimitriou. Logicomix: An Epic Search for Truth. New York: Bloomsbury, 2009. 347 pp., illus. ISBN 978-1-59691-452-0. $22.95 (paper). Jonathan Fetter-Vorm. Trinity: A Graphic History of the First Atomic Bomb. New York: Hill and Wang, 2012. 154 pp., illus. ISBN 978-0-8090-9355-7. $14.95 (paper). Jim Ottaviani and Leyland Myrick. Feynman. New York: First Second, 2011. 272 pp., illus. ISBN 978-1-59643-827-9. $19.99 (paper). Lauren Redniss. Radioactive: Marie & Pierre Curie: A Tale of Love and Fallout. New York: Harper Collins, 2010. 208 pp., illus. ISBN 978-0-06135-132-7. $29.99 (hardcover). Historical Studies in the Natural Sciences, 45(4):621–629, September 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.jstor.org/

stable/10.1525/hsns.2015.45.4.621.

Ottaviani:2011:F

[Ott11] Jim Ottaviani. Feynman. First Second, New York, NY, USA, 2011. ISBN 1-59643-259-4. 262 pp. LCCN QC16.F49 O88

2011. URL https://www.google.com/books/edition/Feynman/

QYYWAj3jCzEC. Art by Leland Myrick, and coloring by Hilary Sycamore.

Overbye:2008:JWP

[Ove08] Dennis Overbye. John A. Wheeler, physicist who coined the term ‘black hole,’ is dead at 96. New York Times, ??(??):B7, April 14, 2008. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/897768871/.

Parnell:2003:QPI

[Par03] Peter Parnell. QED: a play: inspired by the writings of Richard Feynman and Tuva or bust! by Ralph Leighton. Applause, New York, NY, USA, 2003. ISBN 0-8222-1924-7, 1-55783-592-6. 64 pp. LCCN ????

Paz:1989:FOL

[Paz89] Robert Paz. Feynman’s office: the last blackboards. Physics Today, 42(2):88, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Pais:2006:JRO

[PC06] Abraham Pais and Robert P. Crease. J. Robert Oppenheimer: a life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-516673-6. xxii + 353 + 16 pp. LCCN QC16.O62 P35 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0636/2005002173-d.
html;   http://www.loc.gov/catdir/enhancements/fy0723/
2005002173-b.html;   http://www.loc.gov/catdir/toc/ecip056/
2005002173.html.

Piccioni:2017:FSPa

[Pic17a] Robert L. Piccioni. Feynman Simplified Part 1: Mechanics, Light, Waves, & Special Relativity. Real Science Publishing, Westlake Village, CA, USA, 2017. ISBN 1-5209-1071-1. 360 (est.) pp. LCCN ????

Piccioni:2017:FSPb

[Pic17b] Robert L. Piccioni. Feynman Simplified Part 2: Electromagnetism. Real Science Publishing, Westlake Village, CA, USA, 2017. ISBN 1-5217-1706-0. 398 (est.) pp. LCCN ????

Piccioni:2017:FSPc

[Pic17c] Robert L. Piccioni. Feynman Simplified Part 3: Quantum Mechanics. Real Science Publishing, Westlake Village, CA, USA, 2017. ISBN 1-5211-9520-X. 332 (est.) pp. LCCN ????

Piccioni:2017:FSPd

[Pic17d] Robert L. Piccioni. Feynman Simplified Part 4: Math for Physicists & The Best of Feynman. Real Science Publishing, Westlake Village, CA, USA, 2017. ISBN 1-5210-6425-3. 262 (est.) pp. LCCN ????

Pines:1989:RFC

[Pin89] David Pines. Richard Feynman and condensed matter physics. Physics Today, 42(2):61–66, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Pang:1995:SFD

[PJ95] Alex Pang and Chueng-Ryong Ji. Symbolic Feynman-diagram calculation. Computers in Physics, 9(6):589–??, November 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168553.

Popov:1998:EQV

[Pop98] Dušan Popov. Extension of the quantum virial and Hellmann– Feynman theorems to the quantum statistical averages. International Journal of Quantum Chemistry, 69(2):159–165, ????

1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-

461X (electronic). URL http://www3.interscience.wiley.com/
cgi-bin/abstract?ID=29978;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=29978&PLACEBO=IE.pdf.

Price:1994:RWF

[Pri94] Huw Price. Reinterpreting the Wheeler–Feynman absorber theory: Reply to Leeds. British Journal for the Philosophy of Science, 45(4):1023–1028, December 1994. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL

http://bjps.oxfordjournals.org/content/45/4/1023.full.
pdf+html;   http://www.jstor.org/stable/687619.

Pupyshev:2002:HFT

[Pup02] Vladimir I. Pupyshev. Hellmann–Feynman theorem near the threshold. International Journal of Quantum Chemistry, 88(4): 380–391, June 15, 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rall:2006:HDB

[Ral06] Denise N. Rall. The ‘house that Dick built’: Constructing the team that built the bomb. Social Studies of Science, 36(6):943–957, December 1, 2006. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.com/
doi/pdf/10.1177/0306312706062676.

Ratliff:2006:RFL

[Rat06] Steven T. Ratliff. Review: The Feynman Lectures on Physics — The Complete Audio Collection. American Journal of Physics, 74(9):846–??, September 2006. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://scitation.

aip.org/journals/doc/AJPIAS-ft/vol_74/iss_9/846_1.html.

Relkin:2022:SOC

[Rel22] Paul W. Relkin. Solving the Olum 1 cipher. Cryptologia, 46(4):291– 301, 2022. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/01611194.2021.1974124.

Relkin:2023:SOC

[Rel23] Paul W. Relkin. Solving the Olum 2 cipher: a new approach to cryptanalysis of transposition ciphers. Cryptologia, 47(1):38–47, 2023. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
01611194.2021.1992686.

Resag:2018:FHP

[Res18] Jörg Resag. Feynman and His Physics: The Life and Science of an Extraordinary Man. Springer Biographies. Springer International Publishing, Cham, Switzerland, 2018. ISBN 3-319-96835-1 (print), 3-319-96836-X (e-book), 3-319-96837-8. ISSN 2365-0613 (print), 2365-0621 (electronic). xii + 319 pp. LCCN QC16.F49 R4713 2018.

Rico:2007:GBS

[RLER07] J. Fernández Rico, R. López, I. Ema, and G. Ramírez. Generation of basis sets with high degree of fulfillment of the Hellmann– Feynman theorem. Journal of Computational Chemistry, 28(4): 748–758, March 2007. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Roeder:2012:BRF

[Roe12] John L. Roeder. Book review: Feynman. by Jim Ottaviani. Art by Leland Myrick. Coloring by Hilary Sycamore. 266 pp., First Second Books, New York, 2011. Price: $29.99 (cloth). ISBN 978-1-59643-259-8. American Journal of Physics, 80(1):95, January 2012. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://aapt.scitation.org/doi/full/10.1119/
1.3625870.

Rogers:2010:MIS

[Rog10] Kara Rogers, editor. The 100 Most Influential Scientists of All Time. The Britannica guide to the world’s most influential people. Britannica Educational Publishers, in association with Rosen Educational Services, New York, NY, USA, 2010. ISBN 1-61530-002-3 (library binding). 360 pp. LCCN Q162 .A15 2010.

Rigden:2011:BNG

[RS11] John S. Rigden and Roger H. Stuewer. Book notes: [Galileo’s Two Chief World Systems and Two New Sciences; Harald Fritzsch, You Are Wrong, Mr Einstein!]. Physics in Perspective (PIP), 13(3): 373–375, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0067-4.

Rigden:2012:BNL

[RS12] John S. Rigden and Roger H. H. Stuewer. Book notes [Lawrence Krauss, Quantum Man: Richard Feynman’s Life in Science (W. W. Norton, 2011, xvii + 350, $24.95)]. Physics in Perspective (PIP), 14(1):113–115, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-012-0081-1.

Ryckman:2017:BRV

[Ryc17] Thomas Ryckman. Book review: Void: The Strange Physics of Nothing, James Owen Weatherall, Yale U. Press, 2016, 196 p, $26.00, ISBN 978-0-300-20998-3. Physics Today, 70(9):59–60, September 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Sackett:1985:BRR

[Sac85] Penny D. Sackett. Book review: Richard P. Feynman, “Surely You’re Joking Mr. Feynman!” Adventures of a Curious Character. American Journal of Physics, 53(12):1214–??, December 1985. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v53/
i12/p1214_s1.

Sanger:1987:CFS

[San87] David E. Sanger. A computer full of surprises. New York Times, ??(??):D1, D3, May 8, 1987. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.proquest.

com/hnpnewyorktimes/docview/110789956/.

Sasaki:1976:AGF

[Sas76] Tateaki Sasaki. Automatic generation of Feynman graphs in QED. Journal of Computational Physics, 22(2):189–214, October 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900759.

Sauer:2008:ROP

[Sau08] Tilman Sauer. Remarks on the origin of path integration: Einstein and Feynman. In Janke and Pelster [JP08], pages 3–13. ISBN 981-283-726-4, 981-283-727-2 (e-book). LCCN QC174.17.P27 I57 2007eb.

Saxon:1994:BRR

[Sax94] David S. Saxon. Book review: A reshaping in physics: QED and the Men Who Made It, Dyson, Feynman, Schwinger, and Tomonaga. Science, 266(5192):1888–1890, December 16, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/2885047.

Singer:1982:PTN

[SAY^{+}82] I. B. Singer, K. J. Arrow, R. S. Yalow, A. Penzias, R. Feynman, and D. Baltimore. Psychology tomorrow — the Nobel view. Psychology Today, 16(12):21–??, ???? 1982. ISSN 0033-3107.

Scadron:1991:AQT

[Sca91] Michael D. Scadron. Advanced quantum theory and its applications through Feynman diagrams. Texts and monographs in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1991. ISBN 0-387-53681-7 (New York), 3-540-53681-7 (Berlin). xvii + 410 pp. LCCN QC174.12 .S3

1991. URL http://link.springer.com/openurl?genre=book\
%26isbn=978-3-540-53681-9;   http://www.loc.gov/catdir/
enhancements/fy1006/92135240-d.html.

Schwinger:1958:SPQ

[Sch58] Julian Schwinger, editor. Selected Papers on Quantum Electrodynamics. Dover books on engineering and engineering physics. Dover, New York, NY, USA, 1958. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35.

Schweber:1986:FVS

[Sch86] Silvan S. Schweber. Feynman and the visualization of spacetime processes. Reviews of Modern Physics, 58(2):449–508, April 1986. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.58.449;   http://rmp.aps.org/abstract/
RMP/v58/i2/p449_1.

Schwinger:1989:PQE

[Sch89] Julian Schwinger. A path to quantum electrodynamics. Physics Today, 42(2):42–48, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Schweber:1994:QMW

[Sch94a] S. S. (Silvan S.) Schweber. QED and the men who made it: Dyson, Feynman, Schwinger, and Tomonaga. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03685-3, 0-691-03327-7 (paperback). xxviii + 732 pp. LCCN QC680 .S34 1994. US$72.50. URL http://www.loc.gov/

catdir/description/prin021/93033550.html;   http://www.
loc.gov/catdir/toc/prin031/93033550.html.

Schweber:1994:PTS

[Sch94b] Sylvan S. Schweber. A postscript: Tomonaga, Schwinger, Feynman, and Dyson. In QED and the men who made it: Dyson, Feynman, Schwinger, and Tomonaga [Sch94a], pages 572–575. ISBN 0-691-03685-3, 0-691-03327-7 (paperback). LCCN QC680 .S34 1994.

US$72.50. URL http://www.loc.gov/catdir/description/
prin021/93033550.html;   http://www.loc.gov/catdir/toc/
prin031/93033550.html.

Schweber:1994:RFV

[Sch94c] Sylvan S. Schweber. Richard Feynman and the visualization of space–time processes. In QED and the men who made it: Dyson, Feynman, Schwinger, and Tomonaga [Sch94a], chapter 8, pages 373–473. ISBN 0-691-03685-3, 0-691-03327-7 (paperback). LCCN QC680 .S34 1994. US$72.50. URL http://www.loc.gov/

catdir/description/prin021/93033550.html;   http://www.
loc.gov/catdir/toc/prin031/93033550.html.

Schwinger:2003:SPQ

[Sch03] Julian Schwinger. Selected Papers on Quantum Electrodynamics. Dover, New York, NY, USA, 2003. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35. Reprint of [Sch58] with ISBN.

Schweber:2006:BRR

[Sch06] Sam Schweber. Book review: Richard P. Feynman, Perfectly Reasonable Deviations from the Beaten Track: The Letters of Richard P. Feynman. Physics Today, 51(2):51–52, February 2006. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Seidel:2000:FCE

[Sei00] R. W. Seidel. Feynman and computation: exploring the limits of computers. IEEE Annals of the History of Computing, 22(1):90–91, January–March 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Semenov:1998:LPA

[Sem98] A. Semenov. LanHEP — a package for automatic generation of Feynman rules from the Lagrangian. Computer Physics Communications, 115(2–3):124–139, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S001046559800143X.

Semenov:2009:LPA

[Sem09] A. V. Semenov. LanHEP — a package for the automatic generation of Feynman rules in field theory. Version 3.0. Computer Physics Communications, 180(3):431–454, March 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465508003718.

Semenov:2016:LPA

[Sem16] A. Semenov. LanHEP — a package for automatic generation of Feynman rules from the Lagrangian. Version 3.2. Computer Physics Communications, 201(??):167–170, April 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465516000199.

Sohler:1956:LNH

[SF56] Stanley E. Sohler and Richard P. Feynman. Letters: A “new hypotheses” on science and religion. Engineering and Science (Caltech), 20(1):6–12, 52, ???? 1956. CODEN ???? ISSN 0013-7812.

URL http://resolver.caltech.edu/CaltechES:20.1.letters.

See [Fey56a].

Soffel:1995:DER

[SGT^{+}95] M. Soffel, B. Geyer, H. Teichler, W. Eberhardt, F. Krull, H. Jungblut, H. Kolanoski, H. Fritzsch, and G. Münster. d’Inverno: Einführung in die Relativitätstheorie/Wong: Mathematische Physik/Sutton: Electronic Structure of Materials/Mönch: Semiconductor Surfaces and Interfaces/Meadows and Hancock-Beaulieu: Front Page Physics/Markvart: Solar Electricity/Zell: Simulation Neuronaler Netze/Enz u. Meÿenn: W. Pauli: Writings on Physics and Philosophy/Mehra: The Beat of a different Drum The Life and Science of Richard Feynman/Miller: Early Quantum Electrodynamics: A Sourcebook/Schweber: QED and the Men who made it: Dyson, Feynman, Schwinger and Tomonaga. Physikalische Blätter, 51(5):428–434, May 1995. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:

//onlinelibrary.wiley.com/doi/10.1002/phbl.19950510519/
abstract.

Shayak:2017:WTS

[Sha17] B. Shayak. Why trains stay on tracks. American Journal of Physics, 85(3):178–184, March 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See [Feyxx].

Shermer:2005:SFT

[She05] Michael Shermer. Skeptic: The Feynman–Tufte Principle. Scientific American, 292(4):38, April 2005. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v292/n4/full/
scientificamerican0405-38.html;   http://www.nature.com/
scientificamerican/journal/v292/n4/pdf/scientificamerican0405-
38.pdf.

Simms:1980:GAF

[Sim80] D. J. Simms. Geometric aspects of the Feynman integral. Lecture Notes in Mathematics, 836:167–170, 1980. CODEN LN-MAA2. ISBN 3-540-10275-2 (print), 3-540-38405-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0089732/.

Slater:1972:HFV

[Sla72] John C. Slater. Hellmann–Feynman and virial theorems in the X α method. Journal of Chemical Physics, 57(6):2389–2396, September 15, 1972. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://jcp.aip.org/resource/1/
jcpsa6/v57/i6/p2389_s1.

Smirnov:2015:FCI

[Smi15] A. V. Smirnov. FIRE5: a C++ implementation of Feynman Integral REduction. Computer Physics Communications, 189(??): 182–191, April 2015. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465514004123.

Schreiber:2006:GSE

[SNA^{+}06] Liev Schreiber, Liam Neeson, F. Murray Abraham, Peter Thomas, John Lithgow, Simon Callow, Scott Handy, Aidan McArdle, David R. Axelrod, Peter Jones, David Espar, Susan Kopman Lewis, Alastair Reid, Allan Cubitt, Christopher Oxley, Christopher Sykes, Gary Johnstone, Dava. Sobel, and David Bodanis. Genius: the science of Einstein, Feynman, Newton, Darwin, and Galileo. WGBH Boston Video, S. Burlington, VT, USA, widescreen and standard formats. edition, 2006. ISBN 1-59375-575-9, 1-59375-573-2 (Galileo’s battle for the heavens), 1-57807-858-X (Darwin’s dangerous idea), 1-59375-323-3 (Newton’s dark secrets), 1-59375-576-7 (The best mind since Einstein), 1-59375-317-9 (Einstein’s big idea). LCCN Q158.5 .G46 2006. 5 videodiscs (ca. 480 min.).

Spears:1991:BRC

[Spe91] Monroe K. Spears. Book review: Cosmology and the common reader: What Do You Care What Other People Think? by Richard P. Feynman. Infinite in All Directions by Freeman J. Dyson. A Brief History of Time from the Big Bang to Black Holes by Stephen W. Hawking. The Big Bang by Joseph Silk. The Anthropic Cosmological Principle by John D. Barrow and Frank J. Tipler. The Sewanee Review, 99(1):113–121, Winter 1991. ISSN 0037-3052 (print), 1934-421X (electronic). URL http://www.jstor.org/

stable/27546317.

Samuelsson:1998:NLP

[SS98] Bengt Samuelsson and Michael Sohlman, editors. Nobel Lectures in Physics (1963–1970), volume 4 of Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1998. ISBN 981-02-3404-X. ix + 349 pp. LCCN QC71 .P455 1998.

Smirnov:2009:FIE

[ST09] A. V. Smirnov and M. N. Tentyukov. Feynman Integral Evaluation by a Sector decomposiTion Approach (FIESTA). Computer Physics Communications, 180(5):735–746, May 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465508003858.

Stanley:1987:PT

[Sta87] Dick Stanley. A pioneer of thought. Austin American-Statesman, ??(??):??, February 8, 1987.

Stachel:1997:FPQ

[Sta97] John Stachel. Feynman paths and quantum entanglement: Is there any more to the mystery. In Cohen et al. [CHS97], pages 245– 256. ISBN 90-481-4809-X, 94-017-2732-5 (e-book). ISSN 0068-0346. LCCN QC173.96-174.52. URL https://link.springer.

com/chapter/10.1007/978-94-017-2732-7_19.

Staal:1998:STR

[Sta98] F. Staal. Science & theology (response to Freeman Dyson’s review of books by Feynman and Polkinghorne). New York Review of Books, 45(18):76–77, November 19, 1998. ISSN 0028-7504 (print), 1944-7744 (electronic).

Stoltzner:2017:FDM

[Stö17] Michael Stöltzner. Feynman diagrams as models. The Mathematical Intelligencer, 39(2):46–54, May 2017. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Stoltzner:2018:FDM

[Stö18] Michael Stöltzner. Feynman diagrams: Modeling between physics and mathematics. Perspectives on Science, 26(4):482–500, July– August 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274

(electronic). URL http://www.mitpressjournals.org/doi/abs/
10.1162/posc_a_00284.

Stoltzner:2022:PTF

[Stö22] Michael Stöltzner. From probabilistic topologies to Feynman diagrams: Hans Reichenbach on time, genidentity, and quantum physics. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03726-7.

Streit:2008:FIG

[Str08] L. Streit. Feynman integrals as generalized functions on path space: things done and open problems. In Janke and Pelster [JP08], page ?? ISBN 981-283-726-4, 981-283-727-2 (e-book). LCCN QC174.17.P27 I57 2007eb.

Strickland:2011:WSC

[Str11] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Studerus:2010:RFI

[Stu10] C. Studerus. Reduze — Feynman integral reduction in C++. Computer Physics Communications, 181(7):1293–1300, July 2010. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465510000925.

Styer:2011:EFH

[Sty11] Daniel F. Styer. The errors of Feynman and Hibbs. Resonance, 16 (9):849–853, September 2011. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic).

Sykes:1994:NOG

[Syk94] Christopher Sykes, editor. No Ordinary Genius: The Illustrated Richard Feynman. W. W. Norton & Co., New York, NY, USA, 1994. ISBN 0-297-81469-9, 0-393-31393-X (paperback), 0-393-03621-9. 272 pp. LCCN QC16.F49 A3 1994. US$19.95, CDN$25.99.

Taylor:1997:BRF

[Tay97] John C. Taylor. Book review: Feynman: Jagdish Mehra, The Beat of a Different Drum. The Life and Science of Richard Feynman. Oxford University Press, 1994. Pp. 630, £25.00. ISBN 0-19-853948-7, hardback. Notes and Records of the Royal Society of London, 51(1):156–157, January 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Teich:1986:BRR

[Tei86] Malvin C. Teich. Book review: Richard P. Feynman, Surely You’re Joking, Mr. Feynman! Adventures of a Curious Character. Physics Today, 39(9):61, September 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Telegdi:1989:LVF

[Tel89] Valentine L. Telegdi. A lowbrow’s view of Feynman. Physics Today, 42(2):85–88, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Thornber:1970:VAE

[TF70] K. K. Thornber and Richard P. Feynman. Velocity acquired by an electron in a finite electric field in a polar crystal. Physical Review B (Solid State), 1(10):4099–4114, May 15, 1970. CODEN PLRBAQ. ISSN 0556-2805. See erratum, Physical Review B 4, 674.

Thornber:1971:VAE

[TF71] K. K. Thornber and Richard P. Feynman. Velocity acquired by an electron in a finite electric field in a polar crystal. Physical Review B (Solid State), 4(2):674, July 15, 1971. CODEN PLRBAQ. ISSN 0556-2805.

Tentyukov:2000:FDA

[TF00] M. Tentyukov and J. Fleischer. A Feynman diagram analyzer DIANA. Computer Physics Communications, 132(1–2):124–141, October 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001478.

Tentyukov:2004:PCF

[TF04] M. Tentyukov and J. Fleischer. Parallel computation of Feynman diagrams with DIANA. Computer Physics Communications, 160(3):167–186, July 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504002048.

Treiman:1972:PD

[TFJ^{+}72] S. B. Treiman, R. P. Feynman, J. D. Jackson, T. D. Lee, and F. E. Low. Panel discussion. AIP Conference Proceedings, 6(??):164–185, April 1, 1972. CODEN APCPCS. ISSN ???? URL http://link.

aip.org/link/?APCPCS/6/164/1.

Treiman:1964:BRR

[Tre64] S. B. Treiman. Book review: Richard P. Feynman and Robert B. Leighton and Matthew Sands, The Feynman Lectures on Physics. Physics Today, 17(8):45–46, August 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Trefethen:2016:SFC

[Tre16] Lloyd N. Trefethen. Surprises of the Faraday cage. SIAM News, 49(6):??, July/August 21, 2016. ISSN 0036-1437. URL https://

sinews.siam.org/DetailsPage/TabId/900/ArtMID/2243/ArticleID/
757/Surprises-of-the-Faraday-Cage.aspx.

Taylor-Smith:20xx:YPB

[TSxx] Jayna Taylor-Smith. You’ve probably been breaking pasta wrong this whole time. Readers’ Digest, ??(??):??, ???? 20xx. URL

https://www.rd.com/food/fun/breaking-pasta-wrong/.

Turkes:2010:BRQ

[Tür10] Florian Türkes. Book review: Quantenmechanik im Kalten Krieg. David Bohm und Richard Feynman by Christian Forstner. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 94(1):114– 115, ???? 2010. CODEN SUARAH. ISSN 0039-4564. URL http:

//www.jstor.org/stable/20778433.

Taylor:1998:TFS

[TVOT98] Edwin F. Taylor, Stamatis Vokos, John M. O’Meara, and Nora S. Thornber. Teaching Feynman’s sum-over-paths quantum theory. Computers in Physics, 12(2):190–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168652.

Uggerhoj:2016:YCE

[UMF16] U. I. Uggerhøj, R. E. Mikkelsen, and J. Faye. The young centre of the Earth. European Journal of Physics, 37(3):035602:1–035602:8, April 2016. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

Usdin:2009:RRR

[Usd09] Steven T. Usdin. The Rosenberg ring revealed: Industrial-scale conventional and nuclear espionage. Journal of Cold War Studies, 11(3):91–143, Summer 2009. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/
268392;   http://muse.jhu.edu/article/268392/pdf.

Veltman:1994:DPFb

[Vel94] Martinus Veltman. Diagrammatica: the Path to Feynman Diagrams, volume 4 of Cambridge Lecture Notes in Physics. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-45692-4 (paperback), 0-511-56407-4 (e-book). xii + 284 pp. LCCN QC794.6.F4 V45 1994.

Vallarta:1939:SCR

[VF39] Manuel Sandoval Vallarta and Richard P. Feynman. The scattering of cosmic rays by the stars of a galaxy. Physical Review (2), 55(5):506–507, March 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL https://link.aps.

org/doi/10.1103/PhysRev.55.506.

Visconti:1951:TDC

[Vis51] Antoine Visconti. Sur la théorie du champ soustractif de Louis de Broglie et la fonction d’interaction de R. P. Feynman. (French) [On the subtractive field theory of Louis de Broglie and the interaction function of R. P. Feynman]. Comptes rendus de l’Académie des sciences, Paris, 232(??):217–219, ???? 1951. ISSN 0001-4036 (print), 2419-6304 (electronic).

vanOldenborgh:1991:FPE

[vO91] G. J. van Oldenborgh. FF — a package to evaluate one-loop Feynman diagrams. Computer Physics Communications, 66(1): 1–15, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465591900023.

Wang:1993:ACF

[Wan93] Jian-Xiong Wang. Automatic calculation of Feynman loopdiagrams: I. generation of a simplified form of the amplitude. Computer Physics Communications, 77(2):263–285, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390010A.

Wardhaugh:2012:WNA

[War12] Benjamin Wardhaugh, editor. A wealth of numbers: an anthology of 500 years of popular mathematics writing. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-14775-2 (hardcover). xv + 370 pp. LCCN QA7 .W43 2012.

Wayne:1995:BRB

[Way95] Andrew Wayne. Book review: QED and the Men Who Made It: Dyson, Feynman, Schwinger, and Tomonaga by Silvan S. Schweber. British Journal for the Philosophy of Science, 46(4):624–627, December 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://www.jstor.org/stable/687905.

Weatherall:2016:VSP

[Wea16] James Owen Weatherall. Void: the strange physics of nothing. Foundational Questions in Science. Yale University Press, New Haven, CT, USA, 2016. ISBN 0-300-20998-3 (hardcover). 196 pp. LCCN QC28 .W35 2016.

Weiss:1966:BRR

[Wei66] George H. Weiss. Book review: R. P. Feynman and A. R. Hibbs, Quantum Mechanics and Path Integrals. Physics Today, 19(6):89, June 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Weiss:1988:BGH

[Wei88] Eric A. Weiss. Biographies: George H. Brown and Richard P. Feynman. Annals of the History of Computing, 10(2):140–142, April/June 1988. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1988/pdf/a2140.pdf;
 http://www.computer.org/annals/an1988/a2140abs.htm.

West:1993:FTP

[Wes93] Todd H. West. FeynmanParameter and trace — programs for expressing Feynman amplitudes as integrals over Feynman parameters. Computer Physics Communications, 77(2):286–298, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390011Z.

Wheeler:1945:IAM

[WF45] John Archibald Wheeler and Richard Phillips Feynman. Interaction with the absorber as the mechanism of radiation. Reviews of Modern Physics, 17(2–3):157–181, April 1, 1945. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Wheeler:1949:CET

[WF49] John Archibald Wheeler and Richard Phillips Feynman. Classical electrodynamics in terms of direct interparticle action. Reviews of Modern Physics, 21(3):425–433, July 1, 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Whitfield:2007:WMR

[WH07] Donald Whitfield and James L. Hicks, editors. What’s the matter?: readings in physics. Great Books Foundation, Chicago, IL, USA, 2007. ISBN 1-880323-91-5. xx + 540 pp. LCCN QC7.5

.W53 2007. URL http://store.greatbooks.org/adu-wtm.html.

Foreword by Alan Lightman.

Wheeler:1989:YF

[Whe89] John Archibald Wheeler. The young Feynman. Physics Today, 42 (2):24–28, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Whitaker:2016:RFB

[Whi16] Andrew Whitaker. Richard Feynman and Bell’s Theorem. American Journal of Physics, 84(7):493–494, July 2016. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Wilson:1975:AOTa

[Wil75] Robert R. Wilson. All in our time: a recruit for Los Alamos. Bulletin of the Atomic Scientists, 31(3):41–47, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolfram:2016:IMP

[Wol16] Stephen Wolfram. Idea makers: personal perspectives on the lives and ideas of some notable people. Wolfram Media, Inc., Champaign, IL, USA, 2016. ISBN 1-57955-003-7 (hardcover), 1-57955-005-3 (ebook), 1-57955-011-8. 250 (est.) pp. LCCN Q141 .W678562 2016.

URL http://www.wolfram-media.com/products/idea-makers.
html.

Wuthrich:2007:BRB

[Wüt07] Adrian Wüthrich. Book review: Drawing Theories Apart: The Dispersion of Feynman Diagrams in Postwar Physics. David Kaiser. University of Chicago Press, Chicago (2005) (xix + 469 pp., $30.00, Paperback, ISBN 0-226-42266-6). Studies in History and Philosophy of Modern Physics, 38(3):586–589, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000980. See [Kai05b].

Wuthrich:2011:BM

[Wüt11a] Adrian Wüthrich. Back matter. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 191–208.

Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-

0064 (electronic). URL https://link.springer.com/content/
pdf/bbm:978-90-481-9228-1/1.pdf.

Wuthrich:2011:DNM

[Wüt11b] Adrian Wüthrich. The development of a new means of representation: Goals and milestones. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 177–189. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-

0064 (electronic). URL https://link.springer.com/chapter/
10.1007/978-90-481-9228-1_7.

Wuthrich:2011:DEF

[Wüt11c] Adrian Wüthrich. The Dirac equation: Feynman’s great struggle. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 65–111. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/10.1007/978-90-481-9228-1_4.

Wuthrich:2011:FTS

[Wüt11d] Adrian Wüthrich. The field theoretical systematization of Feynman’s theory. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 141–175. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.springer.com/chapter/10.1007/
978-90-481-9228-1_6.

Wuthrich:2011:FPS

[Wüt11e] Adrian Wüthrich. Free propagation and successive scattering. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 113–139. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.1007/978-90-481-9228-1_5.

Wuthrich:2011:FM

[Wüt11f] Adrian Wüthrich. Front matter. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xvii. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.springer.com/content/pdf/bfm:
978-90-481-9228-1/1.pdf.

Wuthrich:2011:GFD

[Wüt11g] Adrian Wüthrich. The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2011. ISBN 90-481-9227-7 (hardcover), 90-481-9228-5 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 208 pp. LCCN QC794.6.F4 W88 2010.

Wuthrich:2011:IOU

[Wüt11h] Adrian Wüthrich. Introduction: Origin, use and interpretation of Feynman diagrams. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–21. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.springer.com/chapter/10.1007/
978-90-481-9228-1_1.

Wuthrich:2011:QEF

[Wüt11i] Adrian Wüthrich. Quantum electrodynamics without Feynman diagrams. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 23–50. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/10.1007/978-90-481-9228-1_2.

Wuthrich:2011:QMH

[Wüt11j] Adrian Wüthrich. Quantum mechanics without a Hamiltonian operator. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 51–64. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/10.1007/978-90-481-9228-1_3.

Wuthrich:2012:IFD

[Wüt12] Adrian Wüthrich. Interpreting Feynman diagrams as visual models. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):172–181, ???? 2012. CODEN ???? ISSN

1913-0465. URL https://spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/article/view/
16131.

Wuthrich:2013:AIP

[Wüt13a] Adrian Wüthrich. Against the impossible picture: Feynman’s heuristics in his search for a divergence-free quantum electrodynamics. Physics and Philosophy, ??(??):??, ???? 2013. ISSN 1863-

7388. URL http://hdl.handle.net/2003/29919. Article ID 019.

Wuthrich:2013:FSD

[Wüt13b] Adrian Wüthrich. Feynman’s struggle and Dyson’s surprise: the development and early application of a new means of representation. In Katzir et al. [KLR13], pages 271–289. ISBN 3-8442-5134-0.

LCCN QC173.98. URL http://www.edition-open-access.de/
proceedings/5/.

Wuthrich:2018:EWS

[Wüt18] Adrian Wüthrich. The exigencies of war and the stink of a theoretical problem: Understanding the Genesis of Feynman’s quantum electrodynamics as mechanistic modelling at different levels. Perspectives on Science, 26(4):501–520, July–August 2018. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://www.mitpressjournals.org/doi/abs/10.
1162/posc_a_00285.

Xiao:2015:EVS

[XW15] Bo Xiao and Hao Wang. An extended version of the simple and fast Feynman diagram generator EasyFeynDiag. Computer Physics Communications, 186(??):108–109, January 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465514003191.

Xiao:2013:SAA

[XWhZ13] Bo Xiao, Hao Wang, and Shou hua Zhu. A simple algorithm for automatic Feynman diagram generation. Computer Physics Communications, 184(8):1966–1972, August 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465513001124.

Zeh:2011:FIQ

[Zeh11] H. D. Zeh. Feynman’s interpretation of quantum theory. European Physical Journal H, 36(1):63–74, July 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2011-10035-2.

Zichichi:1965:SEP

[Zic65] Antonino Zichichi, editor. Symmetries in elementary particle physics. Academic Press, New York, NY, USA, 1965. LCCN QC721 .I578 1964.